### v3

# Comparative Study of Recurrent Neural Networks for Virtual Analog Audio Effects Modeling

Abstract
Analog electronic circuits are at the core of an important category of musical devices, which includes a broad range of sound synthesizers and audio effects. The nonlinear features of their passive and active electronic components give analog musical devices a distinctive timbre and sound quality, making them highly desirable. The development of software that simulates analog musical devices, known as virtual analog modeling, is a significant sub-field in audio signal processing. Artificial neural networks are a promising technique for virtual analog modeling. They have rapidly gained popularity for the emulation of analog audio effects circuits, particularly recurrent networks. While neural approaches have been successful in accurately modeling distortion circuits, they require architectural improvements that account for parameter conditioning and low-latency response. Although hybrid solutions can offer advantages, black-box approaches can still be advantageous in some contexts. In this article, we explore the application of recent machine learning advancements for virtual analog modeling. In particular, we compare State-Space models and Linear Recurrent Units against the more common Long Short-Term Memory networks. These have shown promising ability in sequence-to-sequence modeling tasks, showing a notable improvement in signal history encoding. Our comparative study uses these black-box neural modeling techniques with a variety of audio effects. We evaluate the performance and limitations of these models using multiple metrics, providing insights for future research and development. Our metrics aim to assess the models’ ability to accurately replicate energy envelopes and frequency contents, with a particular focus on transients in the audio signal. To incorporate control parameters into the models, we employ the Feature-wise Linear Modulation method. Long Short-Term Memory networks exhibit better accuracy in emulating distortions and equalizers, while the State-Space model, followed by Long Short-Term Memory networks when integrated in an encoder-decoder structure, and Linear Recurrent Unit outperforms others in emulating saturation and compression. When considering long time-variant characteristics, the State-Space model demonstrates the greatest capability to track history. The Long Short-Term Memory networks present more tendency to introduce audio artifacts.
keywords:
Nonlinear Modeling, Virtual Analog, Audio Effect, Machine Learning
1 Introduction
The emulation of analog musical devices has become incredibly pervasive in digital commercial products. The distinct sound characteristics of both contemporary and old analog audio devices have captivated producers and musicians. nalog effect exists as standalone units as well as integrated with sound synthesizers. There are different types of popular effects such as filters, distortion, overdrive, compressor/limiter, as well as delay-based effects like chorus, flanger, phaser, and reverb. Analog circuits found in audio effects depend on intricately designed circuits that combine discrete passive and active electronic components to generate and modify electrical signals. Key components, such as operational amplifiers and diodes, are utilized in a variety of analog effects. These components present nonlinear behaviors that contribute to the unique sound of these devices. Accurate modeling of such nonlinearities is essential to replicate the behavior of analog audio devices. Consequently, developing digital audio effects capable of simulating analog counterparts has become a crucial and valuable area of study within digital audio signal processing, often referred to as Virtual Analog (VA) modeling.
This article presents a comparative study that applies recent advancements in machine learning to audio effects modeling. Machine learning has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks. Although hybrid solutions (referred to as gray-box approaches) have also been proposed, which offer their advantages  [1], black-box approaches can still be appealing in some contexts. Gray-box approaches can incorporate physics knowledge  [2] or measurements from systems  [3]. While the former may not always be accurate enough, the latter may not be feasible for certain tasks. Universal black-box models have already been proposed  [4], but these usually involve large networks that are not suitable for low-latency and real-time applications. In addition, while machine learning modeling of nearly all types of analog effects has been investigated, they often do not provide a satisfactory response when continuously manipulating conditioning parameters, which is needed to offer users runtime control over the effect’s parameters.
In this study, we compare different recurrent artificial neural network architectures for modeling a selection of audio effects. In particular, we considered overdrive, saturator, low-pass filter, equalizer, and compressor. We selected audio effects with different complexities and time dependencies to investigate how these architectural characteristics influence the results when learning directly from raw audio.
For black-box modeling purposes, Recurrent Neural Networks (RNNs) are often preferred among the different types of artificial neural networks. RNNs utilize internal states, which function as memory, making them particularly suitable in scenarios where time dependencies are crucial. Especially when the inputs are limited to a small segment of digital audio samples and the network does not have a large amount of input information to infer, as can happen with feed-forward architectures. Analog audio effects are dynamic physical systems that exhibit time-dependent characteristics: their output depends on previous internal states. Therefore, dynamic physical systems possess internal states that describe their physical evolution. This similarity to stateful machine learning architectures is often appreciated, as it allows for effective modeling of these systems.
The primary objective of this comparative study is to explore the capabilities of recurrent architectures in learning from raw digital audio and effectively modeling audio effects with different complexities. We aim to determine whether these architectures can learn various types of audio processing and be conditioned by control parameters found in audio effects. Additionally, we investigate whether novel machine learning architectures, which have shown remarkable results in sequence modeling, also perform well in the field of VA.
To address these questions, we apply models based on three different recurrent networks to a variety of audio effects: the popular Long Short-Term Memory (LSTM)  [5], the State-Space models (SSMs)  [6], and the Linear Recurrent Unit (LRU)  [7]. All models consider conditioning on control parameters using the Feature-wise Linear Modulation (FiLM) method  [8] and employ the Gated Linear Unit (GLU)  [9]. For a fair comparison, we ensure that different models have similar numbers of layers and trainable parameters, as well as utilize the same amount of input audio samples.
Moreover, we design these architectures with a limited computational complexity, measured in the number of floating points operations per output audio sample, which is $1.500$ FLOPs per sample or $72$ MFLOPs per second at $48$ KHz, ensuring theoretical real-time computational capabilities on consumer-grade devices. The designed architectures also present limited input-output audio latency, which we have limited to a maximum of $1.33$ ms.
In this study, we consider Convolutional Neural Networks (CNNs) [10] only when integrated with RNNs  [11]. However, when working with small input vectors and considering long-term temporal dependencies, an RNN-based structure with the ability to encode the signal’s history using the states is a particularly effective feature. Therefore, in this study, our focus is on models based on recurrent layers due to their effectiveness in capturing temporal dependencies, and we do not include convolutional-based networks, which are typically more suited for spatial pattern recognition.
Section  2 reviews the recent advancements in modeling analog devices, with a specific focus on the benefits and drawbacks of the recent deep learning methods against other data-driven methods. Section  3 details the methodology, architectures, losses, and datasets used in our experiments. Section  4 presents the obtained results and the consequent discussion on the limit and strength of the current methodology.
2 Background
There are two possible approaches for approximating a continuous-time system with a discrete-time model: physics-based and data-driven. Physics-based approaches discretize existing continuous-time models. Data-driven approaches identify a discrete model directly from discrete measurement data.
Physics-based modeling is used when comprehensive knowledge about the system is available. In the case of analog audio effects, this requires knowing the schematics of the circuits and specific electronic components populating the system. In addition, the circuit and components need to be mathematically modeled and solvable. A numerical scheme has to be formulated, and the obtained discrete equations have to be solved iteratively to produce the sound of the digital output. In the case of an implicit ordinary differential equations system, the Newton-Raphson (NR) iterative method is generally utilized to find the solutions. NR converges quadratically, although this is not always guaranteed. As a consequence, physics-based models generally present a high computational cost.
Data-driven modeling is used when information on the inner circuitry of the audio effect is not available. The model is considered as a mapping between input-output data: input-output measurements are taken as a reference to find a suitable mapping that approximates the system response. There are three main components in the procedure leading to a data-driven model: the dataset, the set of candidate modeling techniques, and the validation procedure. The selection of input signals for which output is measured, the configuration of the effect’s control parameters, and the data collection process have a major impact on the quality of the resulting model. Data-driven approaches rely on measurements that are static representations of the system behavior, leading to higher difficulty in accurately emulating its dynamical aspects.
Common examples of data-driven approaches are the block-oriented Hammerstein model, the Wiener model, and their combinations  [12]. In the Wiener model, a linear model is followed by a static nonlinearity, while in the Hammerstein model is the opposite. These models have a convenient representation, and they are created by estimating the model coefficients from real data and with some error minimization techniques. A proposed method to classify whether the nonlinear system should rather be modeled by a Wiener or a Hammerstein model is in  [13]. These approaches, for the reason of simplification, model linear and nonlinear behavior separately. Using this strategy, two block types can be identified: linear dynamic blocks and static nonlinearities.
Another data-driven approach for modeling analog circuits involves utilizing artificial neural networks (ANNs)  [14]. ANNs are composed of interconnected units, where each connection can transmit a signal to other artificial neurons. These neurons receive and process signals and can transmit them to connected neurons. A real number represents the signal transmitted through a connection, and the output of each neuron is determined by a nonlinear function applied to the sum of its weighted inputs. Neural networks are trained by optimizing the network’s parameters to minimize the error between the predicted output and the actual target values within a given dataset. Gradient-based techniques, such as backpropagation, are commonly employed to estimate the network’s weights. Also, in this scenario, it can be beneficial to incorporate a specific block that focuses on the ”unknown” behavior while separately utilizing a different block to handle the ”known” aspects that have already been accurately modeled.
2.1 Virtual Analog Modeling
Among physics modeling for analog effects, a popular method is based on the wave-digital principle [15]. Wave digital filters (WDF) are constructed by interconnecting simple discrete-time models of individual electronic components (inductors, capacitors, and resistors). Wave digital filters are digital filters based on physical modeling principles. Wave digital (WD) techniques were extensively investigated [16, 17]. Starting with linear elements, WD was expanded to handle nonlinear elements in [18]. More recent advances allowed the realization of multidimensional simulations with multiple nonlinearities  [19]. In [20], authors showed how its quadratic convergence strongly depends on the free parameters and proposed a novel formulation using WD formalism. Recent VA works using this approach include vacuum tube based-circuits as the preamplifier circuit of an amplifier [21], cascaded amplifiers [22], envelope filter guitar effect  [23], booster guitar pedal  [24], and a ring modulator  [25].
Other approaches consider directly deriving the finite-difference time-domain (FDTD) scheme from the continuous physical differential equations. In  [26], this method is used to model an analog tape machine, while in  [27], authors simulate the sonic effect of the Hammond tonewheel organ.
Port-Hamiltonian approach [28] uses the state-space description of power exchange between energy-storing parts, dissipative parts, and external sources in a circuit. Port Hamiltonian formalism has the advantage of preserving the power balance, a beneficial aspect in nonlinear cases  [29]. A port-Hamiltonian formulation is found in  [30], in which a nonlinear passive peaking EQ mode is modeled.
Physics-based methods are popular techniques in circuit modeling. They are more interpretable and flexible (since parametric) but efficiency in this case is not straightforward. Physics modeling can be computationally expensive since they are usually complex formulas and rely on the implicit scheme. In addition, the physics knowledge of the device and in particular the mathematical description may not be challenging to derive which can lead to introducing approximations in the simulated model.
Among data-driven block-oriented models, other popular approaches are the block-oriented Hammerstein model, the Wiener model, and their combinations [12]. In the Wiener model, a linear model is followed by a static nonlinearity, while in the Hammerstein model is the opposite. These models have a convenient representation, and they are created by estimating the model coefficients from real data and with some error minimization techniques. A proposed method to classify whether the nonlinear system should rather be modeled by a Wiener or a Hammerstein model is in [13]. Recent examples are for guitar distortion pedal  [31]  [32]  [33], amplifier  [34]  [35], distortion circuit  [36].
In the last two decades, ANNs have been extensively used for modeling analog audio devices. In particular, traditional ML techniques, initially developed for different purposes, have been investigated for creating models that emulate audio effects taking and processing raw audio samples, such as traditional DSP techniques.
In an early attempt using machine learning for audio effect modeling  [37], a multilayer feedforward network was utilized to learn a high-pass filter and the Ibanez Tube-Screamer distortion effect. The author experimented with a network having one input layer, two hidden layers with sigmoid neurons, and one output layer with a linear activation function. The hidden and output layers present the same size as the input layer, and two different input sizes were considered in the experiments: $512$ and $60$ samples recorded at $44.1$-kHz sampling rate. The training was performed using the MSE as loss function and feeding the network with overlapped segments, in particular, the hop size was half of the input size. The high-pass filter training was carried out on the frequency domain and led to unsatisfactory results, while the distortion effect was trained in the time domain and demonstrated the ability of the network to learn distortion effects. A spurious frequency was found to occur due to the overlapping process since the transition among frames was not sufficiently smooth. Feedforward networks are the simplest type of network, they do not incorporate mechanisms to learn time dependencies but their ability to encode the history of the signal only depends on the amount of input samples they take as input. This aspect limits their use to tasks not involving significant time dependencies.
Other attempts used convolutional-based networks to model the preamplifier circuit from the Fender Bassman 56F-A vacuum-tube amplifier [38]. The dataset has been built using a circuit simulator and the network consists of a stack of dilated causal convolution and fully connected layers. The first fully connected layer uses the gated activation function. The second layer has a hyperbolic tangent activation. Finally, a linear layer outputs the predicted sample. The error-to-signal ratio is the measure used as loss. The selected number of convolutional layers is $10$ with a filter width of three and the dilation pattern $d_{k}$ = ${1,2,4,...,512}$. This configuration leads to a receptive field of $2046$ samples, corresponding to $46$ ms of latency between the input and corresponding output sample, as it uses a sampling rate of $44.1$-kHz. The model presents the fully connected layers with $8$ and the convolutional layers with $16$ units each. The number of samples used as input is set equal to the receptive field. The dataset was simulated with different values of the gain, which is a variable resistor, and used to condition the model. The conditioning is realized in the convolutional layers, and using the gated activation [39]. The work shows that CNNs accurately emulate the reference device better than a fully connected feedforward network and can respond to user control changes.
Using the same architecture, the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals have been modeled as well  [40]. The datasets in this case are collected directly from the devices. Here, the authors investigated more configurations: considering different dilation patterns, activation functions, such as rectified linear unit (ReLU), tanh, and gated activation, numbers of convolutional layers ($10$, $18$, $24$) and number of convolution channels ($2$, $4$, $8$, $16$, and $32$). In this work, the conditioning is not considered. A model with gated activation, $18$ layers, and $16$ convolution channels resulted in lower ESR and real-time capability.
Similarly to a fully connected feedforward network, convolutional networks’ ability to track the past information in the signal depends on the receptive field; a longer receptive field requires a greater number of layers. To overcome this problem, Temporal Convolutional networks (TCN) have been proposed for optical dynamic range compressors such as Teletronix LA-2A Leveling Amplifier [41], which present longer time-variant characteristics than distortion units. TCNs’ strength consists of dilated convolutional networks with rapidly growing dilation factors ($2^{2}$ and $02^{1}$ 0). Increasing the dilation factors increases, in turn, the receptive field and the ability of the networks when dealing with long dependencies while, in turn, not increasing the number of layers. In this case, the TCNs have been applied to $44.1$ kHz sampled audio files. A TCN block consists of a convolutional layer, batch normalization, conditioning using FiLM, and PRelu activation function. Authors compared networks with $4$, $5$, and $10$ TCN blocks, $32$ convolutional channel for each convolutional layer, and $5$, $13$, $15$ as the kernel size. The resulting models present receptive fields of $100$, $300$, and $1000$ ms, respectively. The model with a receptive field of $300$ ms provided the best performances. TCNs provide accurate and efficient implementation with respect to computational complexity. On the other hand, convolutional architectures present an intrinsic latency equal to the length of the receptive field, which is detrimental for live audio applications.
A model, that used a combination of different deep learning architectures, was investigated in  [42]. In this case, the model is divided into three parts: an adaptive front-end, latent space, and synthesis back-end. The front-end performs time-domain convolutions with the raw audio to map it into a latent space. It also generates a residual connection, which facilitates the reconstruction of the waveform by the back-end. The adaptive front-end consists of a convolutional encoder. It contains two CNN layers, one pooling layer, and one residual connection. The first layer has $128$ filters and $64$ channels and is followed by the absolute value as a nonlinear activation function. The second layer has $128$ filters of size $128$. The latent-space DNN contains two layers. The first layer is based on locally connected dense layers of $64$ units and the second layer consists of a fully connected layer of $64$ units. Both of these layers are followed by the softplus function. The DNN modifies the latent representation of the input audio into a new latent representation which is fed into the synthesis back-end. The synthesis back-end consists of an unpooling layer, a deep neural network with smooth adaptive activation functions (DNN-SAAF), and a single CNN layer. DNN-SAAF consists of four fully connected dense layers of 128, 64, 64, and 128 hidden units, respectively. All dense layers are followed by the softplus function except the last layer. The SAAF [43] is used as the nonlinearity for the last layer. The model is first trained without the latent-space DNN and DNN-SAAF, and then these modules are incorporated for a second training. The loss function to be minimized is the mean absolute error (MAE). In both training procedures, the input and target audio are sliced into frames of $1024$ samples with a hop size of $64$ samples. Distortion, overdrive, and EQ effects have been modeled using this method, with a sampling rate equal to $16$ kHz.
In  [4] the previously described architecture is further explored and compared with different variations and with the CNN architecture based on  [38]. The variations mainly consist of incorporating Bi-LSTMs or CNNs into the latent space. In this work, the target devices were Universal Audio Vacuum-Tube Preamplifier 610-B, Universal Audio Transistor-Based Limiter Amplifier 1176LN, and 145 Leslie Speaker Cabinet. The highest score in the listening test was obtained by the variation incorporating Bi-LSTM. The input frame size was $4096$ sampled with a hop size of $2048$ samples. Another variation based on Bi-LSTM and SAAF as latent-space modules is used for plate and spring reverberators  [44]. Here, the loss function to be minimized is the sum of MAE in the time domain and MSE of the log power magnitude spectra of the target and output. These architectures represent a general-purpose deep learning framework for modeling audio effects, which has been shown to be capable of accurately modeling different types of effects. On the other hand, the models are large networks with high computational complexity, leading to difficulty when considering efficient implementation. In addition, this architecture lacks conditioning possibility and focuses on modeling a static representation of the devices (i.e., fixed parameters).
The last group of machine-learning black-box modeling methods we review in this section are those based on RNNs. These have been applied to model a vacuum-tube amplifier  [45], where authors investigated different networks and input sizes in two scenarios: slightly distorted and overdriven signals. The networks consisted of $5$ or $10$ units, while the networks’ input included a variable number of audio samples. The network is autoregressive: it generates the output sample at the current time step using current and past input samples, but also output samples generated at previous time steps. The number of past samples can differ from the input and output signal, and in this case, the experiments considered $5$, $10$, or $50$ input samples and $5$ or $10$ output samples. The work presents preliminary experiments and although the models resulted in poor performance, it demonstrates the feasibility of modeling vacuum-tube amplifiers with RNNs. Larger networks generated lower prediction errors, and the overdrive case resulted in more challenges in modeling accurately.
RNNs, in particular LSTMs, have been also compared to CNNs when applied to the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals  [46]. The architecture consists of one RNN layer and one linear fully connected layer as an output layer with one unit. LSTM and GRU are compared as RNN layers varying the number of units ($32$, $64$, and $96$). The input signal is extended to include an additional value representing the user control’s parameter. The results show that RNNs achieve similar accuracy to the CNN model while requiring significantly less computational cost and latency since requiring only the current input sample to predict the current output. LSTM performed better than GRU networks, and the error decreased when increasing the number of units. In addition, listening tests showed LSTM networks with $32$ units performing similarly to TCN with $300$ ms of latency  [41].
Another RNN-based model utilized for the Klon Centaur guitar pedal circuit is presented in  [47], which is also compared with a WDF model of the same circuit. ML implementation resulted in better computational performance than the WDF model but sounded very similar although the ML implementation has slightly damped high frequencies. The architecture consists of a GRU layer of $8$ units and one linear fully connected layer with one unit as the output layer. Also, here, a single variable control parameter of the audio effect is added as an extra input to condition the inference of the networks.
A delay-based effect is investigated in  [48], where a tape-based delay is modeled using GRU networks using a $44.1$kHz sample rate. Here, the delay trajectory is analyzed and extracted using impulse train signals. This trajectory is used to demodulate the signal before training the models or used to directly guide a delay-line built in a differentiable way and based on  [1] work. The second strategy produced smaller prediction errors. The authors include the modeling of the tape’s noise and the trajectory using diffusion techniques to approximate the probabilistic distribution. The hidden size of the GRU layer was set to $64$, and attempts with larger sizes led to no improvements.
RNNs rely on their internal memory mechanism to encode the history of the signal and system and allow the design of models to take a small number of input samples. In addition, they are compatible with real-time applications. On the other hand, they can still suffer from remembering significant long-term dependencies. Indeed, LSTMs alone have not been successful for LFO delay-based effects, such as flanger and phaser. In  [49] authors found it beneficial to use as extra input the LFO signal representing the frequency response of a time-varying system over time. The LFO signal is first extracted by analyzing the recording at $44.1$kHz and then used to help the LSTM during the learning process. The extra input has shown to be beneficial to learning the LFO-based effect and can be freely controlled after model training, although in the study is considered only static parameter values and not LFO signal changing depth and width over time. The error-to-signal ratio (ESR) is used as a loss function and the model is an LSTM layer followed by a linear fully connected layer with one unit. The size of the LSTM layer is $16$ or $32$ or $48$.
A sequence-to-sequence-based model, based on LSTMs, was designed and proposed for the case of another optical compressor, such as the TubeTech CL 1B compressor  [11], to help LSTMs’ memory. The network is based on two LSTM layers: one acts as an encoder, and the other acts as a decoder. The first takes a segment of past input samples together with the controls’ parameters, and the decoder takes the current input sample (in case of single output) or samples (in case of multi-sample output). The encoder, on processing the input, computes the internal states. These states are used as initial states for the second LSTM. These states represent a summary of the encoder input, which helps the decoder to infer the output. The sequence-to-sequence-based model is compared to LSTMs and fully connected networks, resulting in significantly better accuracy. The dataset, in this case, was at $48$kHz sampling rate. The authors initially considered two control parameters, threshold and ratio, while attack and release parameters were fixed. In  [50], the latter parameters were included. The model’s encoder was replaced with a convolutional layer, and the model was compared to TCNs. Sequence-to-sequence-based models allow for more responsiveness and interactivity but at the expense of an increasing computational cost.
Other works consider hybrid approaches. In  [3], fully connected neural networks exploit measurements within the circuits. Here, the structure includes $2$ to $3$ layers embedded within a discrete-time state-space system and explored for diode-based guitar distortion circuits and the lowpass filter of the Korg MS-20 synthesizer. The method required measurements from the devices’ circuits, which are not always accessible, and is moving away from black-box models. Continuing towards gray box approaches,  [51] is based on the Differentiable Digital Signal Processing (DDSP) framework and applies this concept to RC lowpass filter. Subsequently, the idea is applied to the parametric tone section of a guitar amplifier and a nonlinear guitar overdrive circuit in [52]. The technique learns the parameters of a circuit model from measurements of real devices using a Wiener-Hammerstein model consisting of two IIR filters and a feedforward artificial neural network. The IIR filters are, in this way, trainable and derived from the analysis of the circuit. This approach is also used in [53]. In this case, a cascade of biquad filters is trained, employing a Fourier transform, to model the BOSS MT-2 pedal distortion pedal. The model includes four tunable parameters. Differentiable Digital Filters have been applied for the case of a diode clipper circuit [54], showing higher accuracy and similar computation time when compared to traditional physics-based models. In [55], the authors proposed to learn ordinary differential equations governing the first-order and the second-order diode clipper using fully connected layers and RNNs. The formulation similarly uses the discrete-time state-space as in [3]. Neural networks have also been investigated in the case of wave digital filter formulations [2]. A tube amplifier circuit is simulated using multiport nonlinearities inside a wave digital filter and fully connected layers to learn features from a Kirchhoff-domain dataset. Similarly, in  [56], authors combine recurrent neural networks (RNNs) with the Kirchhoff nodal analysis to create a grey-box guitar amplifier model. Lastly, an example of gray box modeling in the case of the LA-2A leveling amplifier is [57], where it is modeled starting from the structure of a traditional digital compressor but using fully connected and RNN to predict the compressor’s parameters.
Gray box methods have the advantage of helping the learning process with knowledge, potentially leading to more accurate and less computationally demanding models. On the other hand, internal measurement of circuit components is not always feasible because it is invasive or difficult to achieve. Infusing external knowledge could not be trivial. For example, in the case of an optical compressor, the attack and release time are governed by a lighting system whose behavior is difficult to model.
3 Methods
This study specifically investigates recurrent neural networks for virtual analog modeling. In particular, we aim to identify which type of recurrent layer or architectural configuration provides accurate modeling of different types of audio effects, if any. These architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states rather than solely relying on the input received at each iteration. In addition, we restrict the experimentation to relatively small networks, which present low computational complexity and low input-output, allowing implementation on real-time consumer-grade digital audio systems. We have developed models that exhibit minimal input-output delay and are well-suited for the typical audio processing workflow that operates on data in blocks. The designed architectures primarily consist of a recurrent layer along with a conditioning block. This architectural choice helps us incorporate control parameters. The designed models are applied to different audio effects and finally evaluated using different metrics. The primary objective of this study is to explore the capabilities of the different recurrent models in learning from raw digital audio and effectively modeling audio effects. These insights can be used to understand the goodness and weakness of these models and determine which can better suit each audio effect. The models presented here, having limited parameters, are not expected to lead to perfectly accurate emulation but a good starting point.
After selecting several types of analog audio effects to be included in this comparative study, we recorded the dataset using analog effects within our availability. We built the dataset using software plugins emulating specific analog effects. Datasets for hardware or software effects were collected using a tool that we have developed for this specific purpose   [^1] https://github.com/stefanofasciani/DGMD.
In the rest of this section, we first detail the four architectures used in this comparative study; then we present the the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets. Finally, explain the process we used to train and test the models.
3.1 Architectures
The following architectures, based on different types of recurrent layers, are compared in this study: LSTM (a), ED (b), LRU (c), and S4D (d). The four architectures are illustrated in Figure 1. ED is based on the concept of internal state sharing  [11]. These architectures follow the same design, consisting of a linear, fully connected layer, a recurrent layer, and a conditioning block.
The number of input samples is identical for all architectures and set to $64$. This choice represents a trade-off: on the one hand, there is the amount of past input information used to generate the current output, where more is typically better; on the other hand, there is the resulting input-output latency, which increases with the number of input samples. Our goal is to keep this latency below $1.33$ ms (i.e., $64$ samples at $48$ kH.). The input of all architectures is an array, including the $64$ most recent input sample, used to generate the single current output sample. Consequently, the model must perform predictions at an audio rate, one for each output audio sample. The models map a vector of $64$ sample in one scalar output sample:
$y_{n}=g(\boldsymbol{x_{n}})=g(x_{n},x_{n-1},...,x_{n-62},x_{n-63})$ (1)
where $n$ is the discrete time index, and $g$ is the overall function describing the models. Before being fed into the recurrent layer (LSTM, LRU, or S4D), the input samples array $\boldsymbol{x_{n}}$ is processed by a fully connected layer that performs a linear projection:
$\boldsymbol{u_{n}}=\boldsymbol{W_{l}}\boldsymbol{x_{n}}+\boldsymbol{b_{l}}$ (2)
where $\boldsymbol{u_{n}}$ is the linear projected vector having a size equal to the number of hidden units of the fully connected layer. $\boldsymbol{W_{l}}$ and is the weights’ matrices, and $\boldsymbol{b_{l}}$ the vector bias terms.
Figure 1: The four compared architectures: LSTM, ED, LRU, and S4D. All architectures present the same structure: a fully connected layer, a recurrent layer, and a conditioning layer. For the case of ED, the input vector is equally split into two vectors, $\boldsymbol{x_{e}}$ and $\boldsymbol{x_{d}}$. The first vector is used to feed two convolutional layers, which are used to compute the states $[h,c]$. These states are combined with the internal states of the LSTM layer included in the decoder. Specifically, a sigmoid function is applied to the LSTM’s states, which are then element-wise multiplied to $[h,c]$. The result of this operation will be the new LSTM’s internal states. The input, identical in all four cases, is an array of the last $64$ input samples $[x_{n},...,x_{n-63}]$, which produces the output sample $y_{n}$. The inputs are firstly linearly projected and then fed to the recurrent layer. Finally, the current output sample is generated through the output layer.
The first two architectures use LSTMs as the recurrent layer, designed to improve the memory capacity of the vanilla RNN by incorporating internal mechanisms known as gates, regulating the flow of information. The LSTM architecture is governed by the following equations:
$\boldsymbol{f_{n}}$ $=\gamma(\boldsymbol{W}_{f}\boldsymbol{h_{n-1}}+\boldsymbol{U}_{f}\boldsymbol{u_{n}}+\boldsymbol{b}_{f})$
$\boldsymbol{i_{n}}$ $=\gamma(\boldsymbol{W}_{i}\boldsymbol{h_{n-1}}+\boldsymbol{U}_{i}\boldsymbol{u_{n}}+\boldsymbol{b}_{i})$
$\boldsymbol{o_{n}}$ $=\gamma(\boldsymbol{W}_{o}\boldsymbol{h_{n-1}}+\boldsymbol{U}_{o}\boldsymbol{u_{n}}+\boldsymbol{b}_{o})$
$\boldsymbol{c^{\prime}_{n}}$ $=\phi(\boldsymbol{W}_{c}\boldsymbol{h_{n-1}}+\boldsymbol{U}_{c}\boldsymbol{u_{n}}+\boldsymbol{b}_{c})$
$\boldsymbol{c_{n}}$ $=\boldsymbol{f_{n}}\circ\boldsymbol{c_{n-1}}+\boldsymbol{i_{n}}\circ\boldsymbol{c^{\prime}_{n}}$
$\boldsymbol{h_{n}}$ $=\boldsymbol{o_{n}}\circ\phi(\boldsymbol{c_{n}})$ (3)
where $\boldsymbol{h_{n}}$ is the hidden state vector, $\boldsymbol{f}$ is the forget gate, $\boldsymbol{i}$ the input gate, and $\boldsymbol{o}$ the output gate, and $\boldsymbol{c}$ is the cell state vector. The forget gate determines what information from previous steps should be kept or discarded. The input gate decides which information from the current step is relevant to add. The output gate determines the value of the next hidden state. Additionally, the cell state is updated by multiplying the forget vector and the input vector with new values that the neural network deems relevant. The forget vector and input vector are obtained using the sigmoid function, denoted by $\gamma$, which determines how much information should be kept. Values of 00 in this vector indicate information to be discarded, while a value of $1$ indicates information to be kept. Furthermore, the $\phi$ represents the hyperbolic tangent function, used to keep the value of the cell state between $-1$ and $1$. Finally, $\boldsymbol{W}$ and $\boldsymbol{U}$ are the weights’ matrices, and $\boldsymbol{b}$ the vector bias terms.
In the case of ED, the input split into two equal-size parts, $\boldsymbol{x_{d}}$ containing $[x_{n},...,x_{n-31}]$ and $\boldsymbol{x_{e}}$ containing containing $[x_{n-32},...,x_{n-63}]$. The first part feds two separate convolutional layers, which compute the states, $[\boldsymbol{h},\boldsymbol{c}]$. These states are combined with the internal states of the LSTM layer computed during the previous iteration. A sigmoid function is applied to the LSTM internal states and then element-wise multiplied to $[h,c]$. This operation determines how much information should be taken from the computed $[h,c]$. The result of this operation represents the new LSTM’s internal states. This process is detailed in Figure 2.
Figure 2: The states sharing process in ED: the encoder input $\boldsymbol{x_{e}}$ is passed to two convolutional layers, which produce the states $[\boldsymbol{h},\boldsymbol{c}]$. These states are combined with $[\boldsymbol{h_{n-1}},\boldsymbol{c_{n-1}}]$, which are the LSTM states of the previous iteration. The sigmoid function is applied to $[\boldsymbol{h_{n-1}},\boldsymbol{c_{n-1}}]$ and the results are element-wise multiplied to $[\boldsymbol{h},\boldsymbol{c}]$. The outcome will be the new states for the LSTM layer $[\boldsymbol{h_{n}},\boldsymbol{c_{n}}]$.
The second part of the input vector is linearly transformed and fed into the second LSTM, following the same procedure as the other models.
The third architecture is based on LRU. Compared to a vanilla RNN, the LRU uses a linear activation function to compute the hidden state instead of a nonlinear function. This modification allows the memory capability to improve significantly. The LRU equations are the following:
$\boldsymbol{h_{n}}$ $=\boldsymbol{W_{h}}\boldsymbol{h_{n-1}}+\boldsymbol{U_{h}}\boldsymbol{u_{n}}+\boldsymbol{b_{h}}$
$\boldsymbol{o_{n}}$ $=\boldsymbol{W_{o}}\boldsymbol{h_{n}}+\boldsymbol{b_{o}}$ (4)
where $\boldsymbol{h_{n}}$ is the hidden state at time $n$ and acts as ’memory’ of the network, $\boldsymbol{o_{n}}$ the output vector. LRU keeps the recurrent layer linear but moves the nonlinearity to a subsequent fully connected layer. Additionally, to further enhance stability, the LRU incorporates a complex-valued diagonal recurrent matrix initialization and an exponential parameterization. At the same time, a normalization scheme for the hidden states is used to boost the efficiency and accuracy [7].
Lastly, the SSMs operate by considering a continuous-time representation of the state-space formulation that is discretized to obtain the following equations:
$\boldsymbol{h_{n}}$ $=\boldsymbol{\bar{W_{A}}}\boldsymbol{h_{n-1}}+\boldsymbol{\bar{W_{B}}}\boldsymbol{u_{n}}$
$\boldsymbol{o_{n}}$ $=\boldsymbol{\bar{W_{C}}}\boldsymbol{h_{n}}+\boldsymbol{\bar{W_{D}}}\boldsymbol{u_{n}}.$ (5)
The remarkable ability of SSMs to capture long-range dependencies stems from their utilization of a specific state matrix $\boldsymbol{\bar{W_{A}}}$ known as the ”HiPPO matrix” [58]. This matrix is designed to encode all the past input history in $\boldsymbol{h_{n}}$, finding a map from $\boldsymbol{u_{n}}$ to a higher dimensional space $\boldsymbol{h_{n}}$ that represents the compression of the history. The matrix allows the model to be conceptualized as a convolutional model that decomposes an input signal onto an orthogonal system of smooth basis functions. In this way, the state $\boldsymbol{h_{n}}$ encodes the history of the signal. In our study, we employ the diagonal state space model variant, S4D model  [59], which is defined by parameterizing its state matrix as a diagonal matrix, boosting the model’s efficiency. By doing so, the basis kernels have closed-form formulas represented by normalized Legendre polynomials $L_{n}(t)$, resulting in the SSM decomposing the input signal $\boldsymbol{u_{n}}$ onto an infinitely long set of basis functions that are orthogonal with respect to an exponentially decaying measure. This characteristic gives the SSM the ability to model long-range dependencies effectively.
The recurrent layer’s output size will equal the number of hidden units of the layer. In the case of LRU and S4D, a fully connected layer with a hyperbolic tangent activation function also follows the recurrent layer to compensate for a lack of nonlinear activation functions in these layers:
$\boldsymbol{\hat{o}_{n}}=tanh(\boldsymbol{W_{nl}}\boldsymbol{o_{n}}+\boldsymbol{b_{nl}})$ (6)
The conditioning block is then applied just before the output layer, which consists of a linear fully connected layer with one unit, which is described in the following subsection. Initial experiments indicate improved performance when the conditioning block is placed after the recurrent layers rather than before. This suggests that it is more beneficial for the networks to use the information given by the control parameters to project the output of the recurrent layer and determine the extent to which this information influences the final output rather than influencing the inference of the recurrent layer based on the control parameters.
Finally, the output layer, which is a fully connected layer with one unit, computes $y_{n}$:
$y_{n}=(\boldsymbol{W_{out}}\boldsymbol{o^{c}_{n}}+b_{out})$ (7)
where $\boldsymbol{o^{c}_{n}}$ is the output of the conditioning block, as detailed in Figure  3. Since the output layer has one unit, $\boldsymbol{W_{out}}$ is a vector of the same dimensionality as $\boldsymbol{o^{c}_{n}}$.
The number of units is chosen to have approximately $800$ trainable parameters among the architectures and effects since an identical number is achievable due to major differences in the layers. As a result, the selected LSTM and ED present $8$ units in the recurrent layer and $4$ in fully connected ones. LRU and S4D present $12$ units in the recurrent layer, and $6$ in fully connected ones. The number of trainable parameters also depends on the dimensional of the conditioning input $\boldsymbol{p}$, equal to the number of variable parameters in the datasets.
3.1.1 Conditioning Block
The conditioning block consists of the FiLM method and the GLU layer, as can be seen in Figure 3.
Figure 3: The Conditioning block operates as follows: the vector with conditioning values $\boldsymbol{p}$ is first projected using a linear fully connected layer. This projection results in an output vector with twice the length $\boldsymbol{p}$. This output vector is then divided into two halves, labeled $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$. These vectors, $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$, are utilized to perform an affine transformation on the vector $\boldsymbol{o_{n}}$, coming from the recurrent layer. Subsequently, the transformed vector passes through another linear fully connected layer, producing another output vector that is double the length of its input. This output is also split into two halves, denoted as $\boldsymbol{q_{1}}$ and $\boldsymbol{q_{2}}$. To complete the process, the softsign function is applied to $\boldsymbol{q_{2}}$, and the result is multiplied element-wise with $\boldsymbol{q_{1}}$ to produce $\boldsymbol{o^{c}_{n}}$
FiLM is a technique that applies an affine transformation to a vector based on conditioning values. The parameters vector $\boldsymbol{p}$, if applicable, is the input to a linear fully connected layer. The output of this layer is then split into two vectors of equal size: $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$. If $\boldsymbol{o_{n}}$ is the vector to which the transformation is to be applied, the output of the linear fully connected layer will have a length double of $\boldsymbol{o_{n}}$. Consequently, FiLM performs am affine transform on $\boldsymbol{o_{n}}$ using $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$.
Following the projection, it is common to apply a function, such as ReLU or sigmoid, to determine the amount of information that should be passed. In our approach, we replace the activation function with a GLU layer. This layer is more stable than ReLU and exhibits faster learning compared to sigmoid. Similarly to the previous step, the GLU layer consists of a linear fully connected layer that takes the FiLM output vector as input and computes a vector with twice its length. The resulting output is then split equally into two vectors: $\boldsymbol{q_{1}}$ and $\boldsymbol{q_{2}}$. A function is applied to $\boldsymbol{q_{2}}$, and the resulting output is multiplied element-wise with $\boldsymbol{q_{1}}$. Hence, the conditioning block is described by the following:
$\boldsymbol{o^{c}_{n}}=\boldsymbol{q_{1}}\odot softsign(\boldsymbol{q_{2}}),\text{ with }\boldsymbol{q_{1}}$ $=\boldsymbol{\dot{o}_{n}}\boldsymbol{K_{1}}+\boldsymbol{d_{1}}$
$\boldsymbol{q_{2}}$ $=\boldsymbol{\dot{o}_{n}}\boldsymbol{K_{2}}+\boldsymbol{d_{2}}$
$\boldsymbol{\dot{o}_{n}}$ $=(\boldsymbol{\theta}\boldsymbol{o_{n}}+\boldsymbol{\gamma}).$ (8)
The GLU layer determines the flow of information through the network, acting as a logical gate. Typically, a sigmoid activation function is used for this purpose. However, in our approach, we introduce a softsign function instead. The softsign function controls the extent to which the control parameters should positively or negatively influence the final output. A value of 00 indicates a bypass, meaning no influence on the output. This allows for more flexibility.
3.2 Losses and Metrics
The models in this study take raw audio samples as input and produce raw audio samples as output. The most common losses used in this case include the Mean Squared Error (MSE), the Mean Absolute Error (MAE), and the Error-to-Signal Ratio (ESR). Based on our knowledge and supported by empirical experiments, there is no evidence suggesting that more complex loss functions could benefit the learning process, especially when modeling the transformation applied to a signal by a generic audio effect. While using specific and different loss functions for different effects may potentially yield better results, we decided to carry out this study using a fixed loss function across architectures and audio effects. This allows for a fair comparison between the trained models, as they all attempt to minimize the same loss function. Therefore, we selected the MSE for this comparative study as it helps to capture the dynamic aspects of the output signal more effectively [50].
In addition, our quantitative evaluation is based on various metrics to assess the models from different perspectives. In the following expressions, $y_{n}$ and $\hat{y}_{n}$ represent the target and predicted signal, and the symbol $N$ represents the overall number of samples in the output signal.
The first metric is the root-mean-square energy (RMSE) error, defined as:
$M_{RMSE}=\sqrt{\frac{1}{N}\sum_{n=1}^{N}(|y_{n}|^{2}-|\hat{y}_{n}|^{2})}$ (9)
RMSE error provides insights into the energy deviation between the target and prediction. The RMSE envelope is calculated using a frame length $W$ and hop size that is both equal to the minibatch size.
The losses and metrics presented so far calculate the average difference between the target and prediction, which may hide large errors that occur only during signal transients. Transients, such as sharp attacks or decays, may represent a small fraction of the overall duration of an audio signal. In this case, we may obtain a relatively low error even with all transients poorly modeled. Therefore, we have included a metric that considers the differences between pairs of spectrums computed on consecutive overlapping windows, as used in the library  [^2] https://github.com/magenta/ddsp developed after  [1]:
$\begin{split}M_{SF}=\Big{|}\Big{|}\Big{(}\Big{|}|STFT(y_{n})|-|STFT(y_{n-1})|\Big{|}\Big{)}\\ -\Big{(}\Big{|}|STFT(\hat{y}_{n})|-|STFT(\hat{y}_{n-1})|\Big{|}\Big{)}\Big{|}\Big{|}_{1}\end{split}$ (10)
where the Short-Time Fourier Transform (STFT) is computed with a window size equal to $2048$ and hop size equal to $512$ samples. This metric emphasizes the ability to accurately predict portions of the output sound with rapid changes.
To account for the frequency response, we use the multi-resolution STFT  [1]:
$\begin{split}M_{STFT}&=\frac{1}{N}\sum_{m}|||STFT_{m}(y_{n})|-|STFT_{m}(\hat{y}_{n})|||_{1}\\ &+\frac{1}{N}\sum_{m}||\log(|STFT_{m}(y_{n})|)-\log(|STFT_{m}(\hat{y}_{n})|)||_{1}\end{split}$ (11)
which compares linear and logarithmic spectral distances with varying frequency resolutions using the L1 norm. The error is calculated as the sum of the absolute values of the vectors. This metric quantifies the error between the spectra employing different resolutions. We compute two separate multi-resolution STFT errors: $M_{STFT-1}$ with FFT window lengths of $m=[32,64,128]$, and $M_{STFT-2}$ with $m=[256,512,1024]$. This differentiation allows us to evaluate errors occurring in windows with high frequency but low time resolution separately from errors in windows with low frequency but high time resolution. For both cases, the hop size is $25\%$ of the window length.
Similarly, we compare the MFCCs, to compute a metric linked to human sound perception:
$M_{MFCC}=\frac{1}{N}|||MFCC(y_{n})|-|MFCC(\hat{y}_{n})|||_{1}$ (12)
where $1024$ is the FFT length, and we have $80$ Mel bands between $30$ Hz and Nyquist frequency.
3.3 Audio Effects and Datasets
In this section, we introduce the audio effects selected for the study and the composition of their datasets. We proceeded to record the dataset using effects that were at our disposal. In particular, we selected the following effects: overdrive, saturation, equalization, low-pass filter, and compression. Table 1 summarizes all the datasets included in the study. Delay-based effects require a different approach than a direct black box  [49],  [48]. For this reason, they are not included in this study.
The dataset we use is derived from the following hardware devices: Behringer OD300 overdrive pedal  [^3] https://www.behringer.com/product.html?modelCode=P0608, Behringer Neutron  [^4] https://www.behringer.com/product.html?modelCode=P0CM5’s overdrive module, Behringer Neutron’s filter module in low-pass mode, TC Electronic Bucket Brigade Analog Delay used as a saturator  [^5] https://www.tcelectronic.com/product.html?modelCode=P0EBV, CL 1B TubeTech  [^6] http://www.tube-tech.com/cl-1b-opto-compressor/and Teletronix LA-2A  [^7] https://www.uaudio.com/hardware/la-2a.html optical compressors. We have also collected data from two software VA plugins: the Helper Saturator  [^8] https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ  [^9] https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
Overdrive
Both the overdrive devices feature the following control parameters: volume, distortion level, and tone. The OD300 is a distortion pedal typically used with electric guitars, which features two different types of distortion, labeled overdrive and distortion modes. To record the dataset the unit was set to overdrive mode. The device was sampled at $5$ different and equally spaced values for the distortion and tone knobs, between the minimum and maximum values, resulting in a $25$ different combination. A separate output signal signal is recorded for each combination of the control parameters. Neutron’s Overdrive is a module integrated into the Neutron synthesizer. Thanks to the semi-modular nature of this device, we were able to isolate the overdrive module. We recorded the dataset’s varying distortion and tone as we did for the OD300, but sampling $10$ equally spaced values for each parameter. In this case, we recorded the output for $100$ different combinations of control parameters. In both devices, the volume level controls an attenuator, and we fixed this parameter to the maximum (i.e. no attenuation).
Filter
Similarly to the overdrive, the Behringer Neutron integrates a filter module, which features frequency cutoff and resonance knobs. The filter can be set as low-pass, high-pass, and band-pass. We isolated the module and recorded the dataset at $10$ equally spaced values for the cutoff and $5$ for the resonance. The cutoff parameter is between the minimum and maximum, while the resonance parameter is between the minimum to the mid value. having $100$ different combinations and selecting the low-pass filter type. The filter was disconnected from the ADSR envelope module.
Saturator
Helper Saturator software emulates analog saturation, in particular tube and tape saturation. The software presents four parameters: lowpass and highpass cutoffs, saturation type switch (tape or tube), and saturation level. We recorded the dataset for $10$ different values of the saturation level between minimum and maximum. The saturation switch was set to tape mode. Lowpass and highpass cutoffs were set to $20$ and $20000$ Hz, respectively.
The other device under this category is the TC Electronic Bucket Brigade Analog Delay pedal. The pedal is equipped with the well-known bucket brigade circuit, which consists of a series of capacitors that transmit the signal. By regulating the speed at which the capacitors fill and empty, a clock controls and slows down the signal, resulting in a delay as it passes through the bucket brigade device (BBD) circuit. The signal is sampled by storing an electric charge in a capacitor at each clock pulse. The sampling creates harmonic content and saturation. After traversing the BBD, the signal undergoes another low-pass filtering process, attenuating the new harmonics, and is expanded to match the dynamic range of the original input signal as closely as possible. Before entering the BBD section, the signal is typically manipulated as well, compressing and filtering using an anti-aliased filter.
This device features four knobs: time, feedback, volume, and depth, as well as a mode switch. The time knob determines the length of the delay, adjustable up to $600$ ms. The feedback knob controls the number of repetitions of the delayed signal. The volume knob adjusts the volume of the delayed signal, while the depth knob defines the level of modulation. Activating the mode switch enables the modulation effect.
For our study, we primarily want to capture the processed signal without the delay component, which can be digitally simulated  [60]. All knobs were set to their minimum values except for the volume knob, while the mode switch remained deactivated.
Equalizer
The software plug-in Universal Audio Pultec Passive EQ emulates the tube-based equalizer Pultec EQP-1A. The effect presents low-frequency ([20, 30, 60, 100] Hz) and high-frequency ([3, 4, 5, 8, 10, 12, 16]kHz) switches to select the frequencies to boost or cut using the respective boost and attenuator knobs ([0,10]). Additionally, the emulation also has a bandwidth knob ([0,10]). The equalizer boosts and cuts the selected frequency bands, giving some saturation simultaneously. In this case, we recorded the dataset varying the boost and attenuator knobs at $5$ different equally-spaced values up to half of their range, setting $60$ Hz as the low-frequency switch value and 00 (sharp) as the bandwidth, obtaining $25$ different recordings.
Compressor/Limiter
The CL 1B and the LA-2A are analog optical compressors. In optical compressors, a lighting-emitting element is fed with the audio signal that illuminates a light-sensitive resistor. The input signal’s amplitude determines the element’s brightness, which, in turn, changes the resistance in the gain attenuation circuit. We recorded the dataset directly from the compressor. The CL 1B compressor dataset presents four parameters: attack, release, ratio, and threshold. We consider the threshold values to range from 00 to $40$ dBu, ratio values from $2$:$1$ to $10$:$1$, attack time from $0.5$ to $300$ ms, and release time from $0.05$ to $10$ seconds. We consider three equally spaced values for the ratio, attack, and release time and four values for the threshold. The resulting number of parameter combinations is $108$. On the other hand, LA-2A presents the peak reduction knob indicating the amount of compression ([00,$100$]) and a mode switch (limiter, compression). We consider ten equally spaced values for the peak reduction, resulting in $20$ parameter combinations.
Table 1: Datasets with related parameters and ranges considered in the study.
Type Parameters Range Combinations
OD300111Hardware. Overdrive Level [min, max] $25$
Tone [min, max]
Neutron’s Overdrive Level [min, max] $100$
OD module [^1] Hardware. Tone [min, max]
Neutron’s Low pass Cutoff [min, max] $100$
Filter module [^1] Hardware. filter Resonance [min, max]
Helper Saturator Saturation [min, max] $10$
Saturator [^2] Software. Saturation Type Tape
Pultec Equalizer Low frequency 60 Hz $25$
Passive EQ  [^2] Software. Frequency booster [0,10]
Bandwidth 0
Frequency Attenuator [0,10]
LA-2A [^1] Hardware.  [^3] Downloaded. Optical Peak Reduction [0,100] $20$
Compressor Switch Mode [Compressor,Limiter]
CL 1B [^1] Hardware. Optical Threshold [-40, 0] dB $108$
Compressor Ratio 1:[1, 10]
Attack [5, 300] ms
Release [0.005, 10] s
[…]
3.3.1 Data Collection
The data collection process was carried out using a MOTU M4 audio interface to feed a selection of audio signals into the system and simultaneously record its output. To achieve this, the left input channel of the audio interface was connected to the left output channel of the interface itself. The right input channel of the audio interface was connected to the output of the device, and the input of the device to the right output channel of the audio interface. This allows recording effectively both the device’s input and output signals compensating for the minor sound coloring and latency of the audio interface. The audio data was recorded at a sampling rate of $48$ kHz.
The input mono signal has a duration of $45$ seconds for each parameter combination and includes a variety of sounds, such as frequency sweeps covering a range of $20$ Hz to $20$ kHz, white noises with increasing amplitudes (both linear and logarithmic), recordings of instruments such as guitar, bass, drums (both loops and single notes), vocals, piano, pad sounds, and sections from various electronic and rock songs. Finally, the control parameter of each effect was mapped to the range of [0,1].
3.4 Experimenting and Learning
The models are trained for $60$ epochs and using the Adam [61] optimizer with a gradient norm scaling of $1$  [62]. The training was stopped earlier in case of no reduction of validation loss for $10$ epochs. We design a time-based schedule for the learning rate as follows:
$lr=\frac{LR}{e^{4}},\text{ with }LR=3\cdot 10^{4}$ (13)
where LR is the initial learning rate, and $e$ is the number of the epochs. Test losses and evaluation metrics are computed using the model’s weights that minimize the validation loss throughout the training epochs. Finally, the input signal is split into segments of $2400$ samples (equivalent to $50$ ms), to be processed before updating the weights.
We adopted the same learning schedule and minibatch size for all models across all datasets. As explained in Section 3.3, we collected the datasets using the same approach and the same input file to minimize the difference among them. However, it is important to highlight that each dataset and modeling task may benefit from specific adjustments to achieve more accurate results. Different datasets and tasks may require slightly different learning rate schedules; an excessively slow or fast learning rate can lead to suboptimal solutions. The number of samples to process before updating the weights, referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
The composition of the training, validation, and test sets is identical across datasets. We used $80\%$ for the training set, $10\%$ for the validation set, and the remaining $10\%$ for the test set. For effects with variable control parameters, each $45$s recording in the dataset is associated with a specific combination of the effect parameters. The $80-10-10\%$ split was carried out at the individual recording level, ensuring that all control parameter combinations present an equal share in each subset and are equally used for training and evaluation of the model. The initial $10\%$ of all the recordings are used as the test set. The following $10\%$ portion of the recordings is used as a validation set, while the remaining parts were used as the training set. Minor manual adjustments are made to ensure that splitting points fall within segments of silence.
4 Results
The evaluation is based on the metrics described in Section 3.2 and presented in Table 3, and 4. The MSE loss and the epoch-related epoch resulting in lower validation loss during training are reported in Table 2. The losses and the metrics consider all conditioning scenarios and provide an average error across all combinations of parameters. To further verify the model’s capacity to learn the conditioned effect, we provide a series of plots that compare the target signal with the predictions generated by the four distinct architectures. For these plots, we chose the modeling scenario that we believe to be the most challenging based on the effect parameters used as conditioning factors for the networks. Except for the low-pass filter, this typically involves parameters at the maximum of the selected range since the output signal, under these conditions, is likely to deviate most significantly from the input. Regarding the low-pass filter, the parameters are configured for minimum cutoff and maximum resonance. The plots show the RMSE in Figure  4, the spectral flux in Figure  5, the spectra in Figure  7, and the spectrograms in Figure  6. For datasets including conditioning parameters, the plots showing predictions versus the target are generated for a $2.5$ seconds input signal containing bass, guitar, and drum loop sound. These input signals are part of the respective test sets. To ensure a clearer comparative visualization of the spectral flux, we use only the first $1.25$ seconds.
Figure 4: Comparison of the RMSE between the target and predicted output. The RMSE is computed using windows of $4096$ samples with $75\%$ overlap. When the model includes conditioning, the results refer to the highest conditioning values, representing the most challenging scenario.
In the OD300 overdrive dataset, the LSTM model presents the lowest MSE, MAE, and ESR, indicating better general amplitude matching between target and prediction. The S4D model is also accurate, especially when considering the transients and the RMS envelope. This can be observed from the lower error in terms of RMSE and spectral flux. On the other hand, in the case of maximum distortion, as seen in Figures 4 and 5, this superiority is less perceived. ED and LRU are left behind except when looking $M_{SF}$ where ED show better performance than LSTM. Notably, the spectrograms shown in Figure 6 illustrate that S4D model generates more harmonics in the higher frequency region than LSTM, and especially than LRU whose produces a cleaner signal. This could be perceptually advantageous compared to the other model that poorly predicts such region of the spectrum, as summarized by the $M_{MFCC}$ errors. From Figure 7 it is evident how the LSTM-based produces the best matching for energy at low frequencies.
Figure 5: Comparison of the spectral flux between the target and predicted output. The results refer to the highest conditioning values, representing the most challenging scenario.
Models trained on the Neutron OD dataset exhibit larger errors, suggesting that this dataset presents a more challenging modeling case for the architectures being investigated. Both LSTM and LRU models yield similar results in terms of average amplitude and frequency errors. The LSTM models show slightly better performance in the time domain, while the LRU excel in the frequency domain. As observed in Figure 4, both LSTM and LRU model achieves a better match with the target. However, the LSTM model tends to produce more energy, whereas the LRU model yields less energy than the target. The S4D model surpasses previous models in terms ofSF, despite exhibiting a larger mismatch at maximum distortion settings, as visible in Figure  5. The spectrograms indicate that the S4D model exhibits a slightly better match than the LSTM model, even though the latter exhibits smaller errors. Upon informal listening, the S4D model tends to add less distortion necessary, whereas the LSTM-based models tend to do the opposite. Lastly, the LRU model exhibits a later stopping epoch than the other three cases.
Figure 6: Comparison of spectrograms of the target and predicted output. The STFT is computed using windows of $2048$ samples with $25\%$ overlap. The results refer to the highest conditioning values, representing the most challenging scenario.
For models trained on the Saturator dataset, the LRU model presents lower MSE, MAE, and $M_{RMSE}$ values. Meanwhile, the S4D model outperforms the others with respect to ESR and $M_{SF}$ metrics. The ED model is slightly less accurate but performs similarly to S4D. In terms of frequency domain errors, LRU also produces better results. As evidenced in Figures  4 and  5, the S4D and ED models achieve closer matching to the target, whereas the LRU model exhibits significantly lower energy, and the LSTM model introduces artifacts, which can be clearly heard when listening to the audio output and negatively affect its performance. These artifacts can also be seen in Figure 5 and  6. Despite these issues, the S4D and ED models still provide a better match with the target spectrograms. The same trend is observed in Figure 7, suggesting that the LRU model performs better in scenarios where minimal sound alteration is required.
[…]
Figure 7: Comparison of the frequency content between the target and predicted output. The FFT is computed using windows of $512$ samples. The results refer to the highest conditioning values, representing the most challenging scenario.
The optical compressor features a time-variant response with pronounced temporal dependencies, largely influenced by its variable attack and release parameters. This is especially noticeable in the case of the CL 1B compressor, which represents a more challenging task for neural network modeling. Among the tested models, the LSTM is the least accurate, exhibiting higher loss in both the LA-2A and CL 1B cases. In contrast, the S4D and ED models demonstrate superior accuracy for these tasks. Specifically, S4D outperforms the ED in the LA-2A case, while the textbfED excels in the CL 1B case when evaluated using time domain-based metrics. However, the trend reverses in the frequency domain, although ED consistently shows lower $M_{SF}$ and $M_{MFCC}$ values for both cases.
In the LA-2A case, we observe an overall good match for RMSE in Figure 4 and for transient phases in Figure 5. The LSTM and LRU models display slightly more mismatches in both plots. Upon examining Figure  6, spectrograms related to the LRU model appear similar to the target output than those of the S4D model, despite a discernible decrease in accuracy observed in Figure 7. In the case of the compressor, the capability to encode past information of the signal is critical, and S4D exhibits overall better accuracy. The ED and LRU models are designed to improve the temporal tracking abilities of basic LSTM, as confirmed by the results for both compressor devices. This improvement is particularly evident with the CL 1B when setting long attack and release time through the conditioning control parameters. These reach up to $300$ ms and $5$ seconds, respectively. In Figure 4 and  5, the difference across models is subtler, yet the LRU model consistently shows smaller RMS errors than other models and a more accurate match in the frequency domain as visible in Figure  6 and  7.
Modeling the low-pass filter is inherently a more complex task than the equalizer, given its more pronounced alterations to the frequency content of the processed signal. Nevertheless, informal listening evaluations of the predicted outputs indicate that all models struggle with accurately emulating the filter behavior; instead of eliminating high-frequency contents, they attenuate the whole signal, as apparent in Figure 6. However, the LRU model outperforms the others when considering the quantitative metrics. The LSTM model, in particular, tends to incorrectly predict excessive energy in the lower spectrum, as shown in Figure 7.
Table 2: MSE test loss computed on the test set. The test sets include all conditioning values, if applicable. The last column reports the epoch at which the training met the early stopping condition.
Dataset Model MSE Epoch Dataset Model MSE Epoch
OD300 LSTM $\boldsymbol{2.46\cdot 10^{-4}}$ $144$ Neutron’s LSTM $\boldsymbol{2.27\cdot 10^{-3}}$ $4$
ED $3.57\cdot 10^{-4}$ $176$ OD ED $2.79\cdot 10^{-3}$ $3$
LRU $2.98\cdot 10^{-4}$ $170$ LRU $2.43\cdot 10^{-3}$ $100$
S4D $3.23\cdot 10^{-4}$ $165$ S4D $2.62\cdot 10^{-3}$ $8$
LA-2A LSTM $1.11\cdot 10^{-4}$ $59$ CL 1B LSTM $1.78\cdot 10^{-4}$ $74$
ED $9.54\cdot 10^{-5}$ $90$ ED $\boldsymbol{8.96\cdot 10^{-5}}$ $38$
LRU $1.14\cdot 10^{-4}$ $146$ LRU $6.76\cdot 10^{-4}$ $168$
S4D $\boldsymbol{8.43\cdot 10^{-5}}$ $62$ S4D $1.22\cdot 10^{-4}$ $14$
Helper LSTM $4.56\cdot 10^{-4}$ $165$ Pultec LSTM $\boldsymbol{7.41\cdot 10^{-4}}$ $19$
Saturator ED $4.41\cdot 10^{-4}$ $138$ Passive ED $1.11\cdot 10^{-3}$ $101$
LRU $\boldsymbol{4.14\cdot 10^{-4}}$ $176$ EQ LRU $1.03\cdot 10^{-3}$ $150$
S4D $4.40\cdot 10^{-4}$ $149$ S4D $1.06\cdot 10^{-3}$ $11$
Neutron’s LSTM $4.07\cdot 10^{-3}$ $46$
Filter ED $3.36\cdot 10^{-3}$ $9$
LRU $\boldsymbol{2.43\cdot 10^{-3}}$ $188$
S4D $2.75\cdot 10^{-3}$ $12$
[…]
Table 3: MAE, ESR, $M_{RMSE}$, and $M_{SF}$ metrics computed on the test set. The test sets include all conditioning values, if applicable.
Dataset Model MAE ESR $M_{RMSE}$ $M_{SF}$
OD300 LSTM $\boldsymbol{1.14\cdot 10^{2}}$ $\boldsymbol{1.06\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $1.60\cdot 10^{-2}$
ED $1.40\cdot 10^{-2}$ $1.54\cdot 10^{-1}$ $1.30\cdot 10^{-2}$ $1.19\cdot 10^{-2}$
LRU $1.73\cdot 10^{-2}$ $2.35\cdot 10^{-1}$ $1.65\cdot 10^{-2}$ $1.77\cdot 10^{-2}$
S4D $1.29\cdot 10^{-2}$ $1.40\cdot 10^{-1}$ $\boldsymbol{1.20\cdot 10^{-2}}$ $\boldsymbol{8.50\cdot 10^{-3}}$
Neutron’s LSTM $\boldsymbol{2.02\cdot 10^{-2}}$ $\boldsymbol{2.50\cdot 10^{-1}}$ $\boldsymbol{1.86\cdot 10^{-2}}$ $5.74\cdot 10^{-3}$
OD ED $2.54\cdot 10^{-2}$ $3.07\cdot 10^{-1}$ $2.36\cdot 10^{-2}$ $6.33\cdot 10^{-3}$
LRU $2.14\cdot 10^{-2}$ $2.70\cdot 10^{-1}$ $1.94\cdot 10^{-2}$ $6.64\cdot 10^{-3}$
S4D $2.40\cdot 10^{-2}$ $2.88\cdot 10^{-1}$ $2.21\cdot 10^{-2}$ $\boldsymbol{5.71\cdot 10^{-3}}$
Helper LSTM $1.12\cdot 10^{-2}$ $2.67\cdot 10^{-1}$ $1.05\cdot 10^{-2}$ $2.78\cdot 10^{-2}$
Saturator ED $1.15\cdot 10^{-2}$ $2.58\cdot 10^{-1}$ $1.06\cdot 10^{-2}$ $1.29\cdot 10^{-2}$
LRU $\boldsymbol{1.04\cdot 10^{-2}}$ $3.46\cdot 10^{-1}$ $\boldsymbol{9.72\cdot 10^{-3}}$ $1.42\cdot 10^{-2}$
S4D $1.13\cdot 10^{-2}$ $\boldsymbol{2.57\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $\boldsymbol{1.22\cdot 10^{-2}}$
Pultec LSTM $1.36\cdot 10^{-2}$ $\boldsymbol{2.15\cdot 10^{-1}}$ $1.29\cdot 10^{-2}$ $8.85\cdot 10^{-3}$
Passive ED $1.40\cdot 10^{-2}$ $3.22\cdot 10^{-1}$ $1.24\cdot 10^{-2}$ $2.49\cdot 10^{-3}$
EQ LRU $\boldsymbol{1.32\cdot 10^{-2}}$ $3.00\cdot 10^{-1}$ $\boldsymbol{1.16\cdot 10^{-2}}$ $1.99\cdot 10^{-3}$
S4D $1.36\cdot 10^{-2}$ $3.08\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $\boldsymbol{1.79\cdot 10^{-3}}$
LA-2A LSTM $4.77\cdot 10^{-3}$ $1.03\cdot 10^{-1}$ $4.51\cdot 10^{-3}$ $1.12\cdot 10^{-2}$
ED $4.32\cdot 10^{-3}$ $8.91\cdot 10^{-2}$ $4.10\cdot 10^{-3}$ $\boldsymbol{7.28\cdot 10^{-3}}$
LRU $5.10\cdot 10^{-3}$ $1.06\cdot 10^{-1}$ $4.67\cdot 10^{-3}$ $7.84\cdot 10^{-3}$
S4D $\boldsymbol{3.70\cdot 10^{-3}}$ $\boldsymbol{7.87\cdot 10^{-2}}$ $\boldsymbol{3.51\cdot 10^{-3}}$ $1.00\cdot 10^{-2}$
CL 1B LSTM $1.06\cdot 10^{-2}$ $2.92\cdot 10^{-1}$ $9.33\cdot 10^{-3}$ $5.21\cdot 10^{-3}$
ED $\boldsymbol{4.25\cdot 10^{-3}}$ $\boldsymbol{1.46\cdot 10^{-1}}$ $\boldsymbol{4.05\cdot 10^{-3}}$ $\boldsymbol{4.18\cdot 10^{-3}}$
LRU $6.59\cdot 10^{-3}$ $1.88\cdot 10^{-2}$ $6.38\cdot 10^{-3}$ $5.49\cdot 10^{-3}$
S4D $5.19\cdot 10^{-3}$ $2.00\cdot 10^{-1}$ $5.02\cdot 10^{-3}$ $6.12\cdot 10^{-3}$
Neutron’s LSTM $3.90\cdot 10^{-2}$ $1.06$ $3.15\cdot 10^{-2}$ $3.03\cdot 10^{-2}$
Filter ED $2.51\cdot 10^{-2}$ $8.77\cdot 10^{-1}$ $2.18\cdot 10^{-2}$ $2.21\cdot 10^{-2}$
LRU $\boldsymbol{2.11\cdot 10^{-2}}$ $\boldsymbol{6.58\cdot 10^{-1}}$ $\boldsymbol{1.66\cdot 10^{-2}}$ $\boldsymbol{1.95\cdot 10^{-2}}$
S4D $2.22\cdot 10^{-2}$ $7.18\cdot 10^{-1}$ $1.97\cdot 10^{-2}$ $2.00\cdot 10^{-2}$
[…]
Table 4: $M_{STFT-1}$, $M_{STFT-2}$, and $M_{MFCC}$ metrics computed on the test set. The test sets include all conditioning values, if applicable.
Dataset Model $M_{STFT-1}$ $M_{STFT-2}$ $M_{MFCC}$
OD300 LSTM $4.53\cdot 10^{-1}$ $3.56\cdot 10^{-1}$ $6.08\cdot 10^{-1}$
ED $4.47\cdot 10^{-1}$ $3.52\cdot 10^{-1}$ $5.80\cdot 10^{-1}$
LRU $4.65\cdot 10^{-1}$ $4.55\cdot 10^{-1}$ $1.05$
S4D $\boldsymbol{4.19\cdot 10^{-1}}$ $\boldsymbol{3.11\cdot 10^{-1}}$ $\boldsymbol{5.38\cdot 10^{-1}}$
Neutron’s LSTM $5.79\cdot 10^{-1}$ $4.80\cdot 10^{-1}$ $3.49\cdot 10^{-1}$
OD ED $6.68\cdot 10^{-1}$ $5.64\cdot 10^{-1}$ $3.80\cdot 10^{-1}$
LRU $\boldsymbol{3.56\cdot 10^{-1}}$ $\boldsymbol{3.40\cdot 10^{-1}}$ $\boldsymbol{3.39\cdot 10^{-1}}$
S4D $6.40\cdot 10^{-1}$ $5.20\cdot 10^{-1}$ $3.59\cdot 10^{-1}$
Helper LSTM $6.56\cdot 10^{-1}$ $5.97\cdot 10^{-1}$ $4.72\cdot 10^{-1}$
Saturator ED $6.65\cdot 10^{-1}$ $\boldsymbol{5.33\cdot 10^{-1}}$ $4.14\cdot 10^{-1}$
LRU $\boldsymbol{5.62\cdot 10^{-1}}$ $5.44\cdot 10^{-1}$ $\boldsymbol{2.46\cdot 10^{-1}}$
S4D $6.73\cdot 10^{-1}$ $5.41\cdot 10^{-1}$ $3.67\cdot 10^{-1}$
Pultec LSTM $1.10$ $9.33\cdot 10^{-1}$ $2.06\cdot 10^{-1}$
Passive ED $6.74\cdot 10^{-1}$ $4.94\cdot 10^{-1}$ $1.91\cdot 10^{-1}$
EQ LRU $\boldsymbol{1.82\cdot 10^{-1}}$ $\boldsymbol{1.73\cdot 10^{-1}}$ $9.81\cdot 10^{-2}$
S4D $6.04\cdot 10^{-1}$ $4.09\cdot 10^{-1}$ $\boldsymbol{9.44\cdot 10^{-2}}$
LA-2A LSTM $4.68\cdot 10^{-1}$ $4.37\cdot 10^{-1}$ $3.10\cdot 10^{-1}$
ED $\boldsymbol{3.12\cdot 10^{-1}}$ $\boldsymbol{2.30\cdot 10^{-1}}$ $\boldsymbol{2.30\cdot 10^{-1}}$
LRU $2.75\cdot 10^{-1}$ $2.70\cdot 10^{-1}$ $2.93\cdot 10^{-1}$
S4D $3.19\cdot 10^{-1}$ $2.52\cdot 10^{-1}$ $2.53\cdot 10^{-1}$
CL 1B LSTM $4.70\cdot 10^{-1}$ $3.56\cdot 10^{-1}$ $2.26\cdot 10^{-1}$
ED $4.89\cdot 10^{-1}$ $4.17\cdot 10^{-1}$ $\boldsymbol{2.09\cdot 10^{-1}}$
LRU $4.39\cdot 10^{-1}$ $4.38\cdot 10^{-1}$ $4.52\cdot 10^{-1}$
S4D $\boldsymbol{2.95\cdot 10^{-1}}$ $\boldsymbol{2.79\cdot 10^{-1}}$ $2.93\cdot 10^{-1}$
Neutron’s LSTM $9.44\cdot 10^{-1}$ $9.33\cdot 10^{-1}$ $6.57\cdot 10^{-1}$
Filter ED $9.58\cdot 10^{-1}$ $9.38\cdot 10^{-1}$ $6.32\cdot 10^{-1}$
LRU $\boldsymbol{7.89\cdot 10^{-1}}$ $\boldsymbol{7.51\cdot 10^{-1}}$ $\boldsymbol{5.98\cdot 10^{-1}}$
S4D $9.13\cdot 10^{-1}$ $9.01\cdot 10^{-1}$ $6.28\cdot 10^{-1}$
[…]
Generally, performance metrics are relatively close across models across most effects examined in this study. Although the S4D models slightly outperform other models based on the metrics alone, it appears to be the most suitable architecture analog effect modeling, at least considering those within the scope of this study. The ED, LRU, and particularly the S4D models are more effective at encoding past signal information within their internal states. This is suggested by the case of the cases of optical compressors, where they significantly outperform other models. Conversely, the LSTM models appear to perform well only when emulating effects with shorter and less critical time dependencies, such as overdrive, equalizer, and filters. Moreover, LSTM models tend to predict energy in the low-frequency spectrum more accurately.
5 Conclusion
The computational challenges of simulating the nonlinearities of analog circuits, described by complex differential equations, are a significant obstacle for real-time applications. The complexity of these nonlinearities makes them difficult to model accurately, but they also contribute to the unique sound characteristics of the devices, making them appealing for both academic and industrial purposes. Modeling these nonlinear aspects is a difficult task that requires simplifications and assumptions to ensure stability and efficiency in the numerical schemes. Artificial neural networks (ANN) have emerged as a popular data-driven approach for this modeling process. While they offer a convenient and automated approach, they may struggle with the computational complexity and flexibility required for handling control parameters. In some cases, hybrid solutions that incorporate physics knowledge or measurements from internal components may provide advantages. However, there are situations where black-box models are necessary and feasible. Recurrent neural networks (RNNs) utilize internal states as memory mechanisms, allowing them to encode the signal history. This feature enables more interactive solutions as their memory capacity is not solely dependent on the input amount. In this paper, we present a comparative analysis that examines the performance of the widely used Long Short-Term Memory (LSTM) models against recently proposed recurrent models, namely Linear Recurrent Unit (LRU) and State-Sapce models (SSM), which have demonstrated remarkable results in sequence modeling. To carry out this study, we design four models that incorporate parameter conditioning and emphasize low-latency response. The control parameters are integrated into the models using the Feature-wise Linear Modulation (FiLM) method. Our selection of effects includes overdrive, saturation, low pass filter, equalization, and compressors. We evaluate these models using various metrics, which provide valuable insights for future research and development. Specifically, our metrics consider the root-mean-squared energy envelopes, the transients, and the frequency content.
LSTM-based networks excel in replicating distortion and equalizer effects, showing their ability to encode the audio dependencies enough for these specific effects effectively. On the other hand, LSTM, when designed in an encoder-decoder fashion, LRUs and, in particular, SSMs, outperform other models when emulating saturation and compression. In the case of compression, they demonstrate their advantage in accurately capturing the signal’s history. Generally, SSMs have also proved to be an appropriate candidate technique in VA, although the gap with the other models is not always so pronounced.
In this study, we did not carry out architecture—and effect-specific optimizations, which can definitely boost the accuracy, such as tuning hyperparameters and changing the size of the various layers. We limited computational complexity and input-output latency and focused the study on models that can be theoretically implemented in consumer-grade devices for live audio applications.
Conditioning techniques and performances should be further investigated in future work. The current conditioning blocks rely on affine transformation. Still, devices such as distortion or saturation could benefit from more complex transformations to account for the possible nonlinear influence of the parameters.

### v5

@@
-Comparative Study of Recurrent Neural Networks for Virtual Analog Audio Effects Modeling
+Comparative Study of State-based Neural Networks for Virtual Analog Audio Effects Modeling
 Abstract
@@
 Artificial neural networks are a promising technique for virtual analog modeling.
-They have rapidly gained popularity for the emulation of analog audio effects circuits, particularly recurrent networks.
+They have rapidly gained popularity for the emulation of analog audio effects, particularly recurrent networks.
 While neural approaches have been successful in accurately modeling distortion circuits, they require architectural improvements that account for parameter conditioning and low-latency response.
@@
 When considering long time-variant characteristics, the State-Space model demonstrates the greatest capability to track history.
-The Long Short-Term Memory networks present more tendency to introduce audio artifacts.
-keywords:
-Nonlinear Modeling, Virtual Analog, Audio Effect, Machine Learning
+The Long Short-Term Memory networks tend to introduce audio artifacts.
 1 Introduction
-The emulation of analog musical devices has become incredibly pervasive in digital commercial products.
-The distinct sound characteristics of both contemporary and old analog audio devices have captivated producers and musicians.
-nalog effect exists as standalone units as well as integrated with sound synthesizers.
-There are different types of popular effects such as filters, distortion, overdrive, compressor/limiter, as well as delay-based effects like chorus, flanger, phaser, and reverb.
-Analog circuits found in audio effects depend on intricately designed circuits that combine discrete passive and active electronic components to generate and modify electrical signals.
-Key components, such as operational amplifiers and diodes, are utilized in a variety of analog effects.
-These components present nonlinear behaviors that contribute to the unique sound of these devices.
-Accurate modeling of such nonlinearities is essential to replicate the behavior of analog audio devices.
-Consequently, developing digital audio effects capable of simulating analog counterparts has become a crucial and valuable area of study within digital audio signal processing, often referred to as Virtual Analog (VA) modeling.
-This article presents a comparative study that applies recent advancements in machine learning to audio effects modeling.
-Machine learning has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks.
-Although hybrid solutions (referred to as gray-box approaches) have also been proposed, which offer their advantages  [1], black-box approaches can still be appealing in some contexts.
-Gray-box approaches can incorporate physics knowledge  [2] or measurements from systems  [3].
+The emulation of analog musical devices has become pervasive in digital products due to their distinct sound, appealing to producers and musicians.
+Analog circuits use components like operational amplifiers and diodes, whose nonlinear behaviors contribute to creating a unique sound.
+Replicating these devices, called Virtual Analog (VA) modeling, is now a crucial area in digital audio signal processing.
+There are two possible approaches for emulating analog audio effects: physics-based and data-driven.
+Physics-based approaches discretize existing continuous-time models.
+Data-driven approaches identify a discrete model directly from discrete measurement data.
+Machine learning, as a data-driven method, has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks.
+Hybrid solutions, incorporating knowledge of the modeling tasks, also exist and can offer advantages, but data-driven approaches can still be appealing in some contexts.
 While the former may not always be accurate enough, the latter may not be feasible for certain tasks.
-Universal black-box models have already been proposed  [4], but these usually involve large networks that are not suitable for low-latency and real-time applications.
+Data-driven models, also called Black-box, being applicable to different audio modeling tasks, have already been proposed  (Ramírez et al.,, 2020a), but these usually involve large networks that are not suitable for low-latency and real-time applications.
 In addition, while machine learning modeling of nearly all types of analog effects has been investigated, they often do not provide a satisfactory response when continuously manipulating conditioning parameters, which is needed to offer users runtime control over the effect’s parameters.
-In this study, we compare different recurrent artificial neural network architectures for modeling a selection of audio effects.
-In particular, we considered overdrive, saturator, low-pass filter, equalizer, and compressor.
-We selected audio effects with different complexities and time dependencies to investigate how these architectural characteristics influence the results when learning directly from raw audio.
-For black-box modeling purposes, Recurrent Neural Networks (RNNs) are often preferred among the different types of artificial neural networks.
-RNNs utilize internal states, which function as memory, making them particularly suitable in scenarios where time dependencies are crucial.
-Especially when the inputs are limited to a small segment of digital audio samples and the network does not have a large amount of input information to infer, as can happen with feed-forward architectures.
+This article presents a comparative study that applies recurrent and state-based neural network architectures to model audio effects with different complexities.
+Neural networks utilizing internal states and a recurrent formulation are often preferred among the different types of artificial neural networks.
+The most known type is the Recurrent Neural Networks (RNNs).
+A Similar formulation is present in the State-Space Networks (SSMs).
+The internal states function as memory, making them particularly suitable in scenarios where time dependencies are crucial.
+Especially when the inputs are limited to one or a small segment of digital audio samples.
+For this reason, in this study, we consider Convolutional Neural Networks (CNNs) (LeCun et al.,, 1998) only when integrated with RNNs  (Simionato & Fasciani,, 2022).
+The aim is to work with small input vectors while also considering long-term temporal dependencies.
+Therefore, we focus on state-based structures.
 Analog audio effects are dynamic physical systems that exhibit time-dependent characteristics: their output depends on previous internal states.
-Therefore, dynamic physical systems possess internal states that describe their physical evolution.
-This similarity to stateful machine learning architectures is often appreciated, as it allows for effective modeling of these systems.
-The primary objective of this comparative study is to explore the capabilities of recurrent architectures in learning from raw digital audio and effectively modeling audio effects with different complexities.
-We aim to determine whether these architectures can learn various types of audio processing and be conditioned by control parameters found in audio effects.
-Additionally, we investigate whether novel machine learning architectures, which have shown remarkable results in sequence modeling, also perform well in the field of VA.
-To address these questions, we apply models based on three different recurrent networks to a variety of audio effects: the popular Long Short-Term Memory (LSTM)  [5], the State-Space models (SSMs)  [6], and the Linear Recurrent Unit (LRU)  [7].
-All models consider conditioning on control parameters using the Feature-wise Linear Modulation (FiLM) method  [8] and employ the Gated Linear Unit (GLU)  [9].
-For a fair comparison, we ensure that different models have similar numbers of layers and trainable parameters, as well as utilize the same amount of input audio samples.
-Moreover, we design these architectures with a limited computational complexity, measured in the number of floating points operations per output audio sample, which is $1.500$ FLOPs per sample or $72$ MFLOPs per second at $48$ KHz, ensuring theoretical real-time computational capabilities on consumer-grade devices.
-The designed architectures also present limited input-output audio latency, which we have limited to a maximum of $1.33$ ms.
-In this study, we consider Convolutional Neural Networks (CNNs) [10] only when integrated with RNNs  [11].
-However, when working with small input vectors and considering long-term temporal dependencies, an RNN-based structure with the ability to encode the signal’s history using the states is a particularly effective feature.
-Therefore, in this study, our focus is on models based on recurrent layers due to their effectiveness in capturing temporal dependencies, and we do not include convolutional-based networks, which are typically more suited for spatial pattern recognition.
+Therefore, dynamic physical systems possess internal states that describe their physical evolution, similar to stateful machine learning algorithms.
+We designed five architectures based on the popular Long Short-Term Memory (LSTM)  (Hochreiter & Schmidhuber,, 1997), the State-Space models (SSMs)  (Gu et al.,, 2022b; Gu & Dao,, 2023), and the Linear Recurrent Unit (LRU)  (Orvieto et al.,, 2023).
+All models consider conditioning on control parameters using the Feature-wise Linear Modulation (FiLM) method  (Perez et al.,, 2018) and employ the Gated Linear Unit (GLU)  (Dauphin et al.,, 2017).
+The architectures have a limited computational complexity, measured in the number of floating points operations per output audio sample, ensuring theoretical real-time computational capabilities on consumer-grade devices.
+The designed architectures also present limited intrinsic input-output audio samples latency, which we have limited to a maximum of $64$ samples ($1.33$ ms).
+Specifically, we refer to the intrinsic latency given by the architecture constraints, neglecting the latency given by the external DSP processing.
+This latency is due to the lag between the time index of the input and output samples.
 Section  2 reviews the recent advancements in modeling analog devices, with a specific focus on the benefits and drawbacks of the recent deep learning methods against other data-driven methods.
@@
 2 Background
-There are two possible approaches for approximating a continuous-time system with a discrete-time model: physics-based and data-driven.
-Physics-based approaches discretize existing continuous-time models.
-Data-driven approaches identify a discrete model directly from discrete measurement data.
-Physics-based modeling is used when comprehensive knowledge about the system is available.
-In the case of analog audio effects, this requires knowing the schematics of the circuits and specific electronic components populating the system.
-In addition, the circuit and components need to be mathematically modeled and solvable.
-A numerical scheme has to be formulated, and the obtained discrete equations have to be solved iteratively to produce the sound of the digital output.
-In the case of an implicit ordinary differential equations system, the Newton-Raphson (NR) iterative method is generally utilized to find the solutions.
-NR converges quadratically, although this is not always guaranteed.
-As a consequence, physics-based models generally present a high computational cost.
-Data-driven modeling is used when information on the inner circuitry of the audio effect is not available.
-The model is considered as a mapping between input-output data: input-output measurements are taken as a reference to find a suitable mapping that approximates the system response.
-There are three main components in the procedure leading to a data-driven model: the dataset, the set of candidate modeling techniques, and the validation procedure.
-The selection of input signals for which output is measured, the configuration of the effect’s control parameters, and the data collection process have a major impact on the quality of the resulting model.
-Data-driven approaches rely on measurements that are static representations of the system behavior, leading to higher difficulty in accurately emulating its dynamical aspects.
-Common examples of data-driven approaches are the block-oriented Hammerstein model, the Wiener model, and their combinations  [12].
-In the Wiener model, a linear model is followed by a static nonlinearity, while in the Hammerstein model is the opposite.
-These models have a convenient representation, and they are created by estimating the model coefficients from real data and with some error minimization techniques.
-A proposed method to classify whether the nonlinear system should rather be modeled by a Wiener or a Hammerstein model is in  [13].
-These approaches, for the reason of simplification, model linear and nonlinear behavior separately.
-Using this strategy, two block types can be identified: linear dynamic blocks and static nonlinearities.
-Another data-driven approach for modeling analog circuits involves utilizing artificial neural networks (ANNs)  [14].
-ANNs are composed of interconnected units, where each connection can transmit a signal to other artificial neurons.
-These neurons receive and process signals and can transmit them to connected neurons.
-A real number represents the signal transmitted through a connection, and the output of each neuron is determined by a nonlinear function applied to the sum of its weighted inputs.
-Neural networks are trained by optimizing the network’s parameters to minimize the error between the predicted output and the actual target values within a given dataset.
-Gradient-based techniques, such as backpropagation, are commonly employed to estimate the network’s weights.
-Also, in this scenario, it can be beneficial to incorporate a specific block that focuses on the ”unknown” behavior while separately utilizing a different block to handle the ”known” aspects that have already been accurately modeled.
-2.1 Virtual Analog Modeling
-Among physics modeling for analog effects, a popular method is based on the wave-digital principle [15].
-Wave digital filters (WDF) are constructed by interconnecting simple discrete-time models of individual electronic components (inductors, capacitors, and resistors).
-Wave digital filters are digital filters based on physical modeling principles.
-Wave digital (WD) techniques were extensively investigated [16, 17].
-Starting with linear elements, WD was expanded to handle nonlinear elements in [18].
-More recent advances allowed the realization of multidimensional simulations with multiple nonlinearities  [19].
-In [20], authors showed how its quadratic convergence strongly depends on the free parameters and proposed a novel formulation using WD formalism.
-Recent VA works using this approach include vacuum tube based-circuits as the preamplifier circuit of an amplifier [21], cascaded amplifiers [22], envelope filter guitar effect  [23], booster guitar pedal  [24], and a ring modulator  [25].
-Other approaches consider directly deriving the finite-difference time-domain (FDTD) scheme from the continuous physical differential equations.
-In  [26], this method is used to model an analog tape machine, while in  [27], authors simulate the sonic effect of the Hammond tonewheel organ.
-Port-Hamiltonian approach [28] uses the state-space description of power exchange between energy-storing parts, dissipative parts, and external sources in a circuit.
-Port Hamiltonian formalism has the advantage of preserving the power balance, a beneficial aspect in nonlinear cases  [29].
-A port-Hamiltonian formulation is found in  [30], in which a nonlinear passive peaking EQ mode is modeled.
-Physics-based methods are popular techniques in circuit modeling.
-They are more interpretable and flexible (since parametric) but efficiency in this case is not straightforward.
-Physics modeling can be computationally expensive since they are usually complex formulas and rely on the implicit scheme.
-In addition, the physics knowledge of the device and in particular the mathematical description may not be challenging to derive which can lead to introducing approximations in the simulated model.
-Among data-driven block-oriented models, other popular approaches are the block-oriented Hammerstein model, the Wiener model, and their combinations [12].
-In the Wiener model, a linear model is followed by a static nonlinearity, while in the Hammerstein model is the opposite.
-These models have a convenient representation, and they are created by estimating the model coefficients from real data and with some error minimization techniques.
-A proposed method to classify whether the nonlinear system should rather be modeled by a Wiener or a Hammerstein model is in [13].
-Recent examples are for guitar distortion pedal  [31]  [32]  [33], amplifier  [34]  [35], distortion circuit  [36].
+Physics-based modeling for analog audio effects requires comprehensive knowledge of the circuit schematics and components.
+These circuits need to be mathematically modeled and solved.
+A numerical scheme is formulated, and the discrete equations are solved iteratively to generate the digital audio output.
+Data-driven modeling is used when the inner circuitry of the audio effect is unknown.
+It treats the model as a mapping between input-output data: input-output measurements are used to approximate the system response.
+The procedure involves three main components: the dataset, candidate modeling techniques, and the validation process.
+The quality of the resulting model heavily depends on the selection of input signals, configuration of control parameters, and data collection.
+Data-driven approaches rely on measurements that are static representations of the system behavior, leading to greater difficulty in accurately emulating its dynamic aspects.
+Common examples of data-driven approaches include the block-oriented Hammerstein model, the Wiener model, and their combinations (Schoukens & Tiels,, 2017).
+In the Wiener model, a linear dynamic system is followed by a static nonlinearity, whereas in the Hammerstein model, the static nonlinearity precedes the linear dynamic system.
+These models offer a straightforward representation and are derived by estimating the model coefficients from real data using error-minimization techniques.
+A proposed method to classify whether a Wiener or a Hammerstein model should rather model the nonlinear system is in (Primavera et al.,, 2018).
+Recent examples are for guitar distortion pedal  (Eichas et al.,, 2015)  (Köper & Holters,, 2020)  (Darabundit et al.,, 2019), amplifier  (Jr & Bodmann,, 2016)  (Eichas et al.,, 2017), distortion circuit  (Eichas & Zölzer,, 2016).
+A more recent data-driven technique involves utilizing artificial neural networks (ANNs).
+ANNs consist of interconnected units called neurons, where each connection can transmit a signal.
+Each neuron processes these signals and can send them to other connected neurons.
+The transmitted signal is represented by a real number, and the neuron’s output is determined by applying a nonlinear function to the sum of its weighted inputs.
+Training ANNs involve optimizing the network’s parameters to minimize the error between predicted and actual target values in a dataset.
+Gradient-based techniques, like backpropagation, are commonly used to optimize the network’s weights.
+2.1 Neural Networks for Virtual Analog Modeling
 In the last two decades, ANNs have been extensively used for modeling analog audio devices.
-In particular, traditional ML techniques, initially developed for different purposes, have been investigated for creating models that emulate audio effects taking and processing raw audio samples, such as traditional DSP techniques.
-In an early attempt using machine learning for audio effect modeling  [37], a multilayer feedforward network was utilized to learn a high-pass filter and the Ibanez Tube-Screamer distortion effect.
-The author experimented with a network having one input layer, two hidden layers with sigmoid neurons, and one output layer with a linear activation function.
-The hidden and output layers present the same size as the input layer, and two different input sizes were considered in the experiments: $512$ and $60$ samples recorded at $44.1$-kHz sampling rate.
-The training was performed using the MSE as loss function and feeding the network with overlapped segments, in particular, the hop size was half of the input size.
-The high-pass filter training was carried out on the frequency domain and led to unsatisfactory results, while the distortion effect was trained in the time domain and demonstrated the ability of the network to learn distortion effects.
-A spurious frequency was found to occur due to the overlapping process since the transition among frames was not sufficiently smooth.
-Feedforward networks are the simplest type of network, they do not incorporate mechanisms to learn time dependencies but their ability to encode the history of the signal only depends on the amount of input samples they take as input.
-This aspect limits their use to tasks not involving significant time dependencies.
-Other attempts used convolutional-based networks to model the preamplifier circuit from the Fender Bassman 56F-A vacuum-tube amplifier [38].
-The dataset has been built using a circuit simulator and the network consists of a stack of dilated causal convolution and fully connected layers.
-The first fully connected layer uses the gated activation function.
-The second layer has a hyperbolic tangent activation.
-Finally, a linear layer outputs the predicted sample.
-The error-to-signal ratio is the measure used as loss.
-The selected number of convolutional layers is $10$ with a filter width of three and the dilation pattern $d_{k}$ = ${1,2,4,...,512}$.
-This configuration leads to a receptive field of $2046$ samples, corresponding to $46$ ms of latency between the input and corresponding output sample, as it uses a sampling rate of $44.1$-kHz.
-The model presents the fully connected layers with $8$ and the convolutional layers with $16$ units each.
-The number of samples used as input is set equal to the receptive field.
+In particular, traditional ML techniques, initially developed for different purposes, have been investigated to create models that emulate audio effects when taking and processing raw audio samples, such as traditional DSP techniques.
+In an early attempt using machine learning for audio effect modeling  (Mendoza,, 2005), a multilayer feedforward network was utilized to learn a high-pass filter and the Ibanez Tube-Screamer distortion effect.
+The high-pass filter training led to unsatisfactory results, while the distortion effect demonstrated the ability of the network to learn distortion effects.
+Other attempts used convolutional-based networks to model the preamplifier circuit from the Fender Bassman 56F-A vacuum-tube amplifier  (Damskägg et al.,, 2019a).
+The network consists of a stack of dilated causal convolution and fully connected layers.
 The dataset was simulated with different values of the gain, which is a variable resistor, and used to condition the model.
-The conditioning is realized in the convolutional layers, and using the gated activation [39].
+The conditioning is realized in the convolutional layers and using the gated activation  (Oord et al.,, 2016).
 The work shows that CNNs accurately emulate the reference device better than a fully connected feedforward network and can respond to user control changes.
-Using the same architecture, the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals have been modeled as well  [40].
-The datasets in this case are collected directly from the devices.
-Here, the authors investigated more configurations: considering different dilation patterns, activation functions, such as rectified linear unit (ReLU), tanh, and gated activation, numbers of convolutional layers ($10$, $18$, $24$) and number of convolution channels ($2$, $4$, $8$, $16$, and $32$).
+Using the same architecture, the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals have been modeled as well  (Damskägg et al.,, 2019b).
+Here, the authors investigated more configurations, considering different dilation patterns and activation functions.
 In this work, the conditioning is not considered.
-A model with gated activation, $18$ layers, and $16$ convolution channels resulted in lower ESR and real-time capability.
-Similarly to a fully connected feedforward network, convolutional networks’ ability to track the past information in the signal depends on the receptive field; a longer receptive field requires a greater number of layers.
-To overcome this problem, Temporal Convolutional networks (TCN) have been proposed for optical dynamic range compressors such as Teletronix LA-2A Leveling Amplifier [41], which present longer time-variant characteristics than distortion units.
+Similarly to a fully connected feedforward network, in which networks’ ability to track the past information depends on the amount of input received, the convolutional network performance depends on the receptive field; a longer receptive field requires a greater number of layers.
+To overcome this problem, Temporal Convolutional networks (TCN) have been proposed for optical dynamic range compressors such as Teletronix LA-2A Leveling Amplifier (Steinmetz et al.,, 2022), which present longer time-variant characteristics than distortion units.
 TCNs’ strength consists of dilated convolutional networks with rapidly growing dilation factors ($2^{2}$ and $02^{1}$ 0).
 Increasing the dilation factors increases, in turn, the receptive field and the ability of the networks when dealing with long dependencies while, in turn, not increasing the number of layers.
-In this case, the TCNs have been applied to $44.1$ kHz sampled audio files.
-A TCN block consists of a convolutional layer, batch normalization, conditioning using FiLM, and PRelu activation function.
-Authors compared networks with $4$, $5$, and $10$ TCN blocks, $32$ convolutional channel for each convolutional layer, and $5$, $13$, $15$ as the kernel size.
-The resulting models present receptive fields of $100$, $300$, and $1000$ ms, respectively.
+The authors compared TCN networks with presenting receptive fields of $100$, $300$, and $1000$ ms, respectively.
 The model with a receptive field of $300$ ms provided the best performances.
@@
 On the other hand, convolutional architectures present an intrinsic latency equal to the length of the receptive field, which is detrimental for live audio applications.
-A model, that used a combination of different deep learning architectures, was investigated in  [42].
+A model that combines different deep learning architectures was investigated in  (Ramírez et al.,, 2019).
 In this case, the model is divided into three parts: an adaptive front-end, latent space, and synthesis back-end.
 The front-end performs time-domain convolutions with the raw audio to map it into a latent space.
-It also generates a residual connection, which facilitates the reconstruction of the waveform by the back-end.
 The adaptive front-end consists of a convolutional encoder.
-It contains two CNN layers, one pooling layer, and one residual connection.
-The first layer has $128$ filters and $64$ channels and is followed by the absolute value as a nonlinear activation function.
-The second layer has $128$ filters of size $128$.
-The latent-space DNN contains two layers.
-The first layer is based on locally connected dense layers of $64$ units and the second layer consists of a fully connected layer of $64$ units.
-Both of these layers are followed by the softplus function.
-The DNN modifies the latent representation of the input audio into a new latent representation which is fed into the synthesis back-end.
+The latent-space DNN modifies the latent representation of the input audio into a new latent representation, which is fed into the synthesis back-end.
 The synthesis back-end consists of an unpooling layer, a deep neural network with smooth adaptive activation functions (DNN-SAAF), and a single CNN layer.
-DNN-SAAF consists of four fully connected dense layers of 128, 64, 64, and 128 hidden units, respectively.
-All dense layers are followed by the softplus function except the last layer.
-The SAAF [43] is used as the nonlinearity for the last layer.
-The model is first trained without the latent-space DNN and DNN-SAAF, and then these modules are incorporated for a second training.
-The loss function to be minimized is the mean absolute error (MAE).
-In both training procedures, the input and target audio are sliced into frames of $1024$ samples with a hop size of $64$ samples.
-Distortion, overdrive, and EQ effects have been modeled using this method, with a sampling rate equal to $16$ kHz.
-In  [4] the previously described architecture is further explored and compared with different variations and with the CNN architecture based on  [38].
+The SAAF  (Hou et al.,, 2017) is used as the nonlinearity for the last layer.
+Distortion, overdrive, and EQ effects have been modeled using this method.
+In  (Ramírez et al.,, 2020a), the previously described architecture is further explored and compared with different variations and with the CNN architecture based on  (Damskägg et al.,, 2019a).
 The variations mainly consist of incorporating Bi-LSTMs or CNNs into the latent space.
@@
 The highest score in the listening test was obtained by the variation incorporating Bi-LSTM.
-The input frame size was $4096$ sampled with a hop size of $2048$ samples.
-Another variation based on Bi-LSTM and SAAF as latent-space modules is used for plate and spring reverberators  [44].
-Here, the loss function to be minimized is the sum of MAE in the time domain and MSE of the log power magnitude spectra of the target and output.
+Another variation based on Bi-LSTM and SAAF as latent-space modules is used for plate and spring reverberators  (Ramírez et al.,, 2020b).
 These architectures represent a general-purpose deep learning framework for modeling audio effects, which has been shown to be capable of accurately modeling different types of effects.
 On the other hand, the models are large networks with high computational complexity, leading to difficulty when considering efficient implementation.
-In addition, this architecture lacks conditioning possibility and focuses on modeling a static representation of the devices (i.e., fixed parameters).
+In addition, using the Bi-LSTM layer makes the architecture non-causal, not allowing real-time applications.
+This architecture also lacks conditioning possibility and focuses on modeling a static representation of the devices (i.e., fixed parameters).
 The last group of machine-learning black-box modeling methods we review in this section are those based on RNNs.
-These have been applied to model a vacuum-tube amplifier  [45], where authors investigated different networks and input sizes in two scenarios: slightly distorted and overdriven signals.
-The networks consisted of $5$ or $10$ units, while the networks’ input included a variable number of audio samples.
+These have been applied to model a vacuum-tube amplifier  (Covert & Livingston,, 2013), where authors investigated different networks and input sizes in two scenarios: slightly distorted and overdriven signals.
+The networks’ input included a variable number of audio samples.
 The network is autoregressive: it generates the output sample at the current time step using current and past input samples, but also output samples generated at previous time steps.
-The number of past samples can differ from the input and output signal, and in this case, the experiments considered $5$, $10$, or $50$ input samples and $5$ or $10$ output samples.
-The work presents preliminary experiments and although the models resulted in poor performance, it demonstrates the feasibility of modeling vacuum-tube amplifiers with RNNs.
-Larger networks generated lower prediction errors, and the overdrive case resulted in more challenges in modeling accurately.
-RNNs, in particular LSTMs, have been also compared to CNNs when applied to the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals  [46].
+The work presents preliminary experiments, and although the models resulted in poor performance, the work demonstrates the feasibility of modeling vacuum-tube amplifiers with RNNs.
+RNNs, in particular LSTMs, have also been compared to CNNs when applied to the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals  (Wright et al.,, 2019).
 The architecture consists of one RNN layer and one linear fully connected layer as an output layer with one unit.
-LSTM and GRU are compared as RNN layers varying the number of units ($32$, $64$, and $96$).
+LSTM and GRU are compared as RNN layers varying the number of units.
 The input signal is extended to include an additional value representing the user control’s parameter.
 The results show that RNNs achieve similar accuracy to the CNN model while requiring significantly less computational cost and latency since requiring only the current input sample to predict the current output.
-LSTM performed better than GRU networks, and the error decreased when increasing the number of units.
-In addition, listening tests showed LSTM networks with $32$ units performing similarly to TCN with $300$ ms of latency  [41].
-Another RNN-based model utilized for the Klon Centaur guitar pedal circuit is presented in  [47], which is also compared with a WDF model of the same circuit.
-ML implementation resulted in better computational performance than the WDF model but sounded very similar although the ML implementation has slightly damped high frequencies.
-The architecture consists of a GRU layer of $8$ units and one linear fully connected layer with one unit as the output layer.
+LSTM performed better than GRU networks, and the error decreased when the number of units was increased.
+Another RNN-based model utilized for the Klon Centaur guitar pedal circuit is presented in  (Chowdhury,, 2020), which is also compared with a WDF model of the same circuit.
+ML implementation resulted in better computational performance than the WDF model but sounded very similar, although the ML implementation slightly damped high frequencies.
+The architecture consists of a GRU layer and one linear fully connected layer with one unit as the output layer.
 Also, here, a single variable control parameter of the audio effect is added as an extra input to condition the inference of the networks.
-A delay-based effect is investigated in  [48], where a tape-based delay is modeled using GRU networks using a $44.1$kHz sample rate.
+A delay-based effect is investigated in  (Mikkonen et al.,, 2023), where a tape-based delay is modeled using GRU networks.
 Here, the delay trajectory is analyzed and extracted using impulse train signals.
-This trajectory is used to demodulate the signal before training the models or used to directly guide a delay-line built in a differentiable way and based on  [1] work.
-The second strategy produced smaller prediction errors.
-The authors include the modeling of the tape’s noise and the trajectory using diffusion techniques to approximate the probabilistic distribution.
-The hidden size of the GRU layer was set to $64$, and attempts with larger sizes led to no improvements.
-RNNs rely on their internal memory mechanism to encode the history of the signal and system and allow the design of models to take a small number of input samples.
-In addition, they are compatible with real-time applications.
-On the other hand, they can still suffer from remembering significant long-term dependencies.
+This trajectory is used to demodulate the signal before training the models or used to directly guide a delay-line built in a differentiable way and based on  (Engel et al.,, 2020) work.
+RNNs rely on their internal memory mechanism to encode the history of the signal, but they can still suffer from remembering significant long-term dependencies.
 Indeed, LSTMs alone have not been successful for LFO delay-based effects, such as flanger and phaser.
-In  [49] authors found it beneficial to use as extra input the LFO signal representing the frequency response of a time-varying system over time.
-The LFO signal is first extracted by analyzing the recording at $44.1$kHz and then used to help the LSTM during the learning process.
+In  (Wright & Välimäki,, 2021), authors found it beneficial to use as extra input the LFO signal representing the frequency response of a time-varying system over time.
+The LFO signal is first extracted by analyzing the recording and then used to help the LSTM during the learning process.
 The extra input has shown to be beneficial to learning the LFO-based effect and can be freely controlled after model training, although in the study is considered only static parameter values and not LFO signal changing depth and width over time.
-The error-to-signal ratio (ESR) is used as a loss function and the model is an LSTM layer followed by a linear fully connected layer with one unit.
-The size of the LSTM layer is $16$ or $32$ or $48$.
-A sequence-to-sequence-based model, based on LSTMs, was designed and proposed for the case of another optical compressor, such as the TubeTech CL 1B compressor  [11], to help LSTMs’ memory.
+A sequence-to-sequence-based model, based on LSTMs, was designed and proposed for the case of another optical compressor, such as the TubeTech CL 1B compressor  (Simionato & Fasciani,, 2022), to help LSTMs’ memory.
 The network is based on two LSTM layers: one acts as an encoder, and the other acts as a decoder.
 The first takes a segment of past input samples together with the controls’ parameters, and the decoder takes the current input sample (in case of single output) or samples (in case of multi-sample output).
-The encoder, on processing the input, computes the internal states.
+The encoder, when processing the input, computes the internal states.
 These states are used as initial states for the second LSTM.
@@
 The sequence-to-sequence-based model is compared to LSTMs and fully connected networks, resulting in significantly better accuracy.
-The dataset, in this case, was at $48$kHz sampling rate.
-The authors initially considered two control parameters, threshold and ratio, while attack and release parameters were fixed.
-In  [50], the latter parameters were included.
+The authors initially considered two control parameters, threshold, and ratio, while attack and release parameters were fixed.
+In  (Simionato & Fasciani,, 2023), the latter parameters were included.
 The model’s encoder was replaced with a convolutional layer, and the model was compared to TCNs.
@@
 Other works consider hybrid approaches.
-In  [3], fully connected neural networks exploit measurements within the circuits.
-Here, the structure includes $2$ to $3$ layers embedded within a discrete-time state-space system and explored for diode-based guitar distortion circuits and the lowpass filter of the Korg MS-20 synthesizer.
-The method required measurements from the devices’ circuits, which are not always accessible, and is moving away from black-box models.
-Continuing towards gray box approaches,  [51] is based on the Differentiable Digital Signal Processing (DDSP) framework and applies this concept to RC lowpass filter.
-Subsequently, the idea is applied to the parametric tone section of a guitar amplifier and a nonlinear guitar overdrive circuit in [52].
+In  (Parker et al.,, 2019), fully connected neural networks exploit measurements within the circuits.
+Here, the network is embedded within a discrete-time state-space system and explored for diode-based guitar distortion circuits and the lowpass filter of the Korg MS-20 synthesizer.
+The method required measurements from the devices’ circuits, which are not always accessible.
+Continuing towards gray box approaches,  (Kuznetsov et al.,, 2020) is based on the Differentiable Digital Signal Processing (DDSP) framework and applies this concept to RC lowpass filter.
+Subsequently, the idea is applied to the parametric tone section of a guitar amplifier and a nonlinear guitar overdrive circuit in (Esqueda et al.,, 2021).
 The technique learns the parameters of a circuit model from measurements of real devices using a Wiener-Hammerstein model consisting of two IIR filters and a feedforward artificial neural network.
 The IIR filters are, in this way, trainable and derived from the analysis of the circuit.
-This approach is also used in [53].
+This approach is also used in (Nercessian et al.,, 2021).
 In this case, a cascade of biquad filters is trained, employing a Fourier transform, to model the BOSS MT-2 pedal distortion pedal.
 The model includes four tunable parameters.
-Differentiable Digital Filters have been applied for the case of a diode clipper circuit [54], showing higher accuracy and similar computation time when compared to traditional physics-based models.
-In [55], the authors proposed to learn ordinary differential equations governing the first-order and the second-order diode clipper using fully connected layers and RNNs.
-The formulation similarly uses the discrete-time state-space as in [3].
-Neural networks have also been investigated in the case of wave digital filter formulations [2].
+Differentiable Digital Filters have been applied for the case of a diode clipper circuit (Chowdhury & Clarke,, 2022), showing higher accuracy and similar computation time when compared to traditional physics-based models.
+In (Wilczek et al.,, 2022), the authors proposed to learn ordinary differential equations governing the first-order and the second-order diode clipper using fully connected layers and RNNs.
+The formulation similarly uses the discrete-time state-space as in (Parker et al.,, 2019).
+Neural networks have also been investigated in the case of wave digital filter formulations (Darabundit et al.,, 2022).
 A tube amplifier circuit is simulated using multiport nonlinearities inside a wave digital filter and fully connected layers to learn features from a Kirchhoff-domain dataset.
-Similarly, in  [56], authors combine recurrent neural networks (RNNs) with the Kirchhoff nodal analysis to create a grey-box guitar amplifier model.
-Lastly, an example of gray box modeling in the case of the LA-2A leveling amplifier is [57], where it is modeled starting from the structure of a traditional digital compressor but using fully connected and RNN to predict the compressor’s parameters.
-Gray box methods have the advantage of helping the learning process with knowledge, potentially leading to more accurate and less computationally demanding models.
-On the other hand, internal measurement of circuit components is not always feasible because it is invasive or difficult to achieve.
-Infusing external knowledge could not be trivial.
-For example, in the case of an optical compressor, the attack and release time are governed by a lighting system whose behavior is difficult to model.
+Similarly, in  (Miklanek et al.,, 2023), authors combine recurrent neural networks (RNNs) with the Kirchhoff nodal analysis to create a grey-box guitar amplifier model.
+Lastly, an example of gray box modeling in the case of the LA-2A leveling amplifier is (Wright & Välimäki,, 2022), where it is modeled starting from the structure of a traditional digital compressor but using fully connected and RNN to predict the compressor’s parameters.
 3 Methods
-This study specifically investigates recurrent neural networks for virtual analog modeling.
-In particular, we aim to identify which type of recurrent layer or architectural configuration provides accurate modeling of different types of audio effects, if any.
-These architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states rather than solely relying on the input received at each iteration.
-In addition, we restrict the experimentation to relatively small networks, which present low computational complexity and low input-output, allowing implementation on real-time consumer-grade digital audio systems.
-We have developed models that exhibit minimal input-output delay and are well-suited for the typical audio processing workflow that operates on data in blocks.
-The designed architectures primarily consist of a recurrent layer along with a conditioning block.
-This architectural choice helps us incorporate control parameters.
+This study aims to identify which type of recurrent layer or architectural configuration accurately models different types of audio effects, if any.
+These insights can be used to understand the goodness and weakness of these models and determine which can better suit each audio effect.
+The states-based architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states rather than solely relying on the input received at each iteration.
+In addition, we restrict the experimentation to relatively small networks fed by a small number of input segments, which present low computational complexity and low input-output latency, allowing implementation on real-time consumer-grade digital audio systems.
 The designed models are applied to different audio effects and finally evaluated using different metrics.
-The primary objective of this study is to explore the capabilities of the different recurrent models in learning from raw digital audio and effectively modeling audio effects.
-These insights can be used to understand the goodness and weakness of these models and determine which can better suit each audio effect.
-The models presented here, having limited parameters, are not expected to lead to perfectly accurate emulation but a good starting point.
-After selecting several types of analog audio effects to be included in this comparative study, we recorded the dataset using analog effects within our availability.
-We built the dataset using software plugins emulating specific analog effects.
-Datasets for hardware or software effects were collected using a tool that we have developed for this specific purpose   [^1] https://github.com/stefanofasciani/DGMD.
-In the rest of this section, we first detail the four architectures used in this comparative study; then we present the the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets.
+To be noticed, the models presented here, having limited parameters, are not expected to lead to perfectly accurate emulation but a good starting point.
+In the rest of this section, we first detail the four architectures used in this comparative study; then, we present the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets.
 Finally, explain the process we used to train and test the models.
@@
 The four architectures are illustrated in Figure 1.
-ED is based on the concept of internal state sharing  [11].
+ED is based on the concept of internal state sharing  (Simionato & Fasciani,, 2022).
 These architectures follow the same design, consisting of a linear, fully connected layer, a recurrent layer, and a conditioning block.
@@
 The input of all architectures is an array, including the $64$ most recent input sample, used to generate the single current output sample.
+This implies that a new input sample enters the buffer with each inference cycle, while the oldest is discarded.
+Therefore, the set of the $64$ most recent input samples constitutes the network input, which, after initial compression by the FC layer, is used to update the states in the recurrent layers.
+This approach aids the network in making predictions while maintaining a stateful design using truncated backpropagation through time.
+Predicting one sample at a time helps minimize the audible artifacts commonly associated with machine-learning models of audio effects, which arise from slight amplitude mismatches at the boundaries of consecutive output segments.
 Consequently, the model must perform predictions at an audio rate, one for each output audio sample.
-The models map a vector of $64$ sample in one scalar output sample:
-$y_{n}=g(\boldsymbol{x_{n}})=g(x_{n},x_{n-1},...,x_{n-62},x_{n-63})$ (1)
+The models map a vector of $64$ sample in one scalar output sample, give the internal states $\boldsymbol{h}$:
+$y_{n}=g(\boldsymbol{x_{n}})=g(x_{n},x_{n-1},...,x_{n-62},x_{n-63};\boldsymbol{h})$ (1)
 where $n$ is the discrete time index, and $g$ is the overall function describing the models.
-Before being fed into the recurrent layer (LSTM, LRU, or S4D), the input samples array $\boldsymbol{x_{n}}$ is processed by a fully connected layer that performs a linear projection:
+Before being fed into the recurrent layer (LSTM, LRU, S4D, or S6), the input samples array $\boldsymbol{x_{n}}$ is processed by a fully connected layer that performs a linear projection:
 $\boldsymbol{u_{n}}=\boldsymbol{W_{l}}\boldsymbol{x_{n}}+\boldsymbol{b_{l}}$ (2)
@@
 $\boldsymbol{W_{l}}$ and is the weights’ matrices, and $\boldsymbol{b_{l}}$ the vector bias terms.
-Figure 1: The four compared architectures: LSTM, ED, LRU, and S4D.
+Figure 1: The four compared architectures: LSTM, ED, LRU, S4D, and S6.
 All architectures present the same structure: a fully connected layer, a recurrent layer, and a conditioning layer.
 For the case of ED, the input vector is equally split into two vectors, $\boldsymbol{x_{e}}$ and $\boldsymbol{x_{d}}$.
-The first vector is used to feed two convolutional layers, which are used to compute the states $[h,c]$.
-These states are combined with the internal states of the LSTM layer included in the decoder.
-Specifically, a sigmoid function is applied to the LSTM’s states, which are then element-wise multiplied to $[h,c]$.
-The result of this operation will be the new LSTM’s internal states.
-The input, identical in all four cases, is an array of the last $64$ input samples $[x_{n},...,x_{n-63}]$, which produces the output sample $y_{n}$.
-The inputs are firstly linearly projected and then fed to the recurrent layer.
-Finally, the current output sample is generated through the output layer.
 The first two architectures use LSTMs as the recurrent layer, designed to improve the memory capacity of the vanilla RNN by incorporating internal mechanisms known as gates, regulating the flow of information.
@@
 Finally, $\boldsymbol{W}$ and $\boldsymbol{U}$ are the weights’ matrices, and $\boldsymbol{b}$ the vector bias terms.
+After the LSTM layer, a linear FC layer reduces the dimensionality down to $4$, ensuring consistency with all other models before proceeding to the conditioning block.
 In the case of ED, the input split into two equal-size parts, $\boldsymbol{x_{d}}$ containing $[x_{n},...,x_{n-31}]$ and $\boldsymbol{x_{e}}$ containing containing $[x_{n-32},...,x_{n-63}]$.
@@
 Additionally, to further enhance stability, the LRU incorporates a complex-valued diagonal recurrent matrix initialization and an exponential parameterization.
-At the same time, a normalization scheme for the hidden states is used to boost the efficiency and accuracy [7].
+At the same time, a normalization scheme for the hidden states is used to boost the efficiency and accuracy (Orvieto et al.,, 2023).
 Lastly, the SSMs operate by considering a continuous-time representation of the state-space formulation that is discretized to obtain the following equations:
@@
 $\boldsymbol{o_{n}}$ $=\boldsymbol{\bar{W}_{C}}\boldsymbol{h_{n}}+\boldsymbol{\bar{W}_{D}}\boldsymbol{u_{n}}.$ (5)
-The remarkable ability of SSMs to capture long-range dependencies stems from their utilization of a specific state matrix $\boldsymbol{\bar{W_{A}}}$ known as the ”HiPPO matrix” [58].
+In our study, we employ the diagonal state space model variant S4D  (Gu et al.,, 2022a), and newer S6 variant (Gu & Dao,, 2023).
+The remarkable ability of S4D to capture long-range dependencies stems from their utilization of a specific state matrix $\boldsymbol{\bar{W_{A}}}$ known as the "HiPPO matrix" (Gu et al.,, 2020).
 This matrix is designed to encode all the past input history in $\boldsymbol{h_{n}}$, finding a map from $\boldsymbol{u_{n}}$ to a higher dimensional space $\boldsymbol{h_{n}}$ that represents the compression of the history.
@@
 In this way, the state $\boldsymbol{h_{n}}$ encodes the history of the signal.
-In our study, we employ the diagonal state space model variant, S4D model  [59], which is defined by parameterizing its state matrix as a diagonal matrix, boosting the model’s efficiency.
+In the diagonal form, the S4D model is defined by parameterizing its state matrix as a diagonal matrix, boosting the model’s efficiency.
 By doing so, the basis kernels have closed-form formulas represented by normalized Legendre polynomials $L_{n}(t)$, resulting in the SSM decomposing the input signal $\boldsymbol{u_{n}}$ onto an infinitely long set of basis functions that are orthogonal with respect to an exponentially decaying measure.
-This characteristic gives the SSM the ability to model long-range dependencies effectively.
+In contrast, the S6 layer introduces input-dependent matrices.
+Specifically, $\boldsymbol{\bar{W_{B}}}$ and $\boldsymbol{\bar{W_{C}}}$ become dependent on the input and computed using a linear FC layer, while $\boldsymbol{\bar{W_{A}}}$ and $\boldsymbol{\bar{W_{D}}}$’s parameters are independent of the input and learned during the training process.
 The recurrent layer’s output size will equal the number of hidden units of the layer.
@@
 The conditioning block is then applied just before the output layer, which consists of a linear fully connected layer with one unit, which is described in the following subsection.
-Initial experiments indicate improved performance when the conditioning block is placed after the recurrent layers rather than before.
+Placing the conditioning block after the recurrent layers rather than before has been found beneficial  (Fasciani et al.,, 2024a).
 This suggests that it is more beneficial for the networks to use the information given by the control parameters to project the output of the recurrent layer and determine the extent to which this information influences the final output rather than influencing the inference of the recurrent layer based on the control parameters.
@@
 3.1.1 Conditioning Block
-The conditioning block consists of the FiLM method and the GLU layer, as can be seen in Figure 3.
-Figure 3: The Conditioning block operates as follows: the vector with conditioning values $\boldsymbol{p}$ is first projected using a linear fully connected layer.
-This projection results in an output vector with twice the length $\boldsymbol{p}$.
-This output vector is then divided into two halves, labeled $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$.
-These vectors, $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$, are utilized to perform an affine transformation on the vector $\boldsymbol{o_{n}}$, coming from the recurrent layer.
-Subsequently, the transformed vector passes through another linear fully connected layer, producing another output vector that is double the length of its input.
-This output is also split into two halves, denoted as $\boldsymbol{q_{1}}$ and $\boldsymbol{q_{2}}$.
-To complete the process, the softsign function is applied to $\boldsymbol{q_{2}}$, and the result is multiplied element-wise with $\boldsymbol{q_{1}}$ to produce $\boldsymbol{o^{c}_{n}}$
+The conditioning block consists of the FiLM method and the GLU layer, as in (Fasciani et al.,, 2024a), and is illustrated in Figure 3.
+Figure 3: Conditioning block based on the FiLM method and the GLU layer.
+The GLU layer features the softsign activation, $\boldsymbol{p}$ denotes the conditioning vector, and $\boldsymbol{o_{n}}$ represents the network’s internal representation of the input.
 FiLM is a technique that applies an affine transformation to a vector based on conditioning values.
@@
 $\boldsymbol{q_{2}}$ $=\boldsymbol{\dot{o}_{n}}\boldsymbol{K_{2}}+\boldsymbol{d_{2}}$
-$\boldsymbol{\dot{o}_{n}}$ $=(\boldsymbol{\theta}\boldsymbol{o_{n}}+\boldsymbol{\gamma}).$ (8)
+$\boldsymbol{\dot{o}_{n}}$ $=(\boldsymbol{\theta}\boldsymbol{o_{n}}+\boldsymbol{\eta}).$ (8)
 The GLU layer determines the flow of information through the network, acting as a logical gate.
@@
 This allows for more flexibility.
+Lastly, the presented models are designed to have limited computational complexity for the inference process.
+Specifically, we set the limit to $1.500$ FLOPs per sample or $72$ MFLOPs per second at $48$ KHz.
+By doing so, the models feature $\sim 750$ internal parameters.
+Table 1 reports the FLOPs per sample for each model.
+Table 1: FLOPs per sample for each model considered in this study.
+Model Recurrent Layer Conditioning Block Total
+LSTM $1040$ $120$ $1160$
+ED $928$ $120$ $1048$
+LRU $692$ $120$ $812$
+S4D $792$ $120$ $912$
+S6 $864$ $120$ $984$
 3.2 Losses and Metrics
@@
 This allows for a fair comparison between the trained models, as they all attempt to minimize the same loss function.
-Therefore, we selected the MSE for this comparative study as it helps to capture the dynamic aspects of the output signal more effectively [50].
+Therefore, we selected the MSE for this comparative study as it helps to capture the dynamic aspects of the output signal more effectively  (Simionato & Fasciani,, 2023).
 In addition, our quantitative evaluation is based on various metrics to assess the models from different perspectives.
-In the following expressions, $y_{n}$ and $\hat{y}_{n}$ represent the target and predicted signal, and the symbol $N$ represents the overall number of samples in the output signal.
-The first metric is the root-mean-square energy (RMSE) error, defined as:
-$M_{RMSE}=\sqrt{\frac{1}{N}\sum_{n=1}^{N}(|y_{n}|^{2}-|\hat{y}_{n}|^{2})}$ (9)
+In the following expressions, $y_{n}$ and $\hat{y}_{n}$ represent the target and predicted signal and the symbol $N$ represents the overall number of samples in the output signal.
+The first metric is the root-mean-square energy (RMSE) error.
 RMSE error provides insights into the energy deviation between the target and prediction.
-The RMSE envelope is calculated using a frame length $W$ and hop size that is both equal to the minibatch size.
 The losses and metrics presented so far calculate the average difference between the target and prediction, which may hide large errors that occur only during signal transients.
@@
 In this case, we may obtain a relatively low error even with all transients poorly modeled.
-Therefore, we have included a metric that considers the differences between pairs of spectrums computed on consecutive overlapping windows, as used in the library  [^2] https://github.com/magenta/ddsp developed after  [1]:
+Therefore, we have included a metric related to the spectral flux, which considers the differences between pairs of spectrums computed on consecutive overlapping windows, as used in the library  [^1] https://github.com/magenta/ddsp developed after  (Engel et al.,, 2020):
 $\begin{split}M_{SF}=\Big{|}\Big{|}\Big{(}\Big{|}|STFT(y_{n})|-|STFT(y_{n-1})|\Big{|}\Big{)}\\ -\Big{(}\Big{|}|STFT(\hat{y}_{n})|-|STFT(\hat{y}_{n-1})|\Big{|}\Big{)}\Big{|}\Big{|}_{1}\end{split}$ (9)
 where the Short-Time Fourier Transform (STFT) is computed with a window size equal to $2048$ and hop size equal to $512$ samples.
-This metric emphasizes the ability to accurately predict portions of the output sound with rapid changes.
-To account for the frequency response, we use the multi-resolution STFT  [1]:
+This metric emphasizes the ability to predict portions of the output sound with rapid changes.
+To account for the frequency response, we use the multi-resolution STFT  (Engel et al.,, 2020):
 $\begin{split}M_{STFT}&=\frac{1}{N}\sum_{m}|||STFT_{m}(y_{n})|-|STFT_{m}(\hat{y}_{n})|||_{1}\\ &+\frac{1}{N}\sum_{m}||\log(|STFT_{m}(y_{n})|)-\log(|STFT_{m}(\hat{y}_{n})|)||_{1}\end{split}$ (10)
@@
 This metric quantifies the error between the spectra employing different resolutions.
-We compute two separate multi-resolution STFT errors: $M_{STFT-1}$ with FFT window lengths of $m=[32,64,128]$, and $M_{STFT-2}$ with $m=[256,512,1024]$.
-This differentiation allows us to evaluate errors occurring in windows with high frequency but low time resolution separately from errors in windows with low frequency but high time resolution.
-For both cases, the hop size is $25\%$ of the window length.
-Similarly, we compare the MFCCs, to compute a metric linked to human sound perception:
+We compute the metrics using $m=[256,512,1024]$ as resolutions.
+Similarly, we compare the MFCCs to compute a metric linked to human sound perception:
 $M_{MFCC}=\frac{1}{N}|||MFCC(y_{n})|-|MFCC(\hat{y}_{n})|||_{1}$ (11)
@@
 Table 2 summarizes all the datasets included in the study.
-Delay-based effects require a different approach than a direct black box  [49],  [48].
+Delay-based effects require a different approach than a direct black box  (Wright & Välimäki,, 2021),  (Mikkonen et al.,, 2023).
 For this reason, they are not included in this study.
 The dataset we use is derived from the following hardware devices: Behringer OD300 overdrive pedal  [^2] https://www.behringer.com/product.html?modelCode=P0608, Behringer Neutron  [^3] https://www.behringer.com/product.html?modelCode=P0CM5’s overdrive module, Behringer Neutron’s filter module in low-pass mode, TC Electronic Bucket Brigade Analog Delay used as a saturator  [^4] https://www.tcelectronic.com/product.html?modelCode=P0EBV, CL 1B TubeTech  [^5] http://www.tube-tech.com/cl-1b-opto-compressor/and Teletronix LA-2A  [^6] https://www.uaudio.com/hardware/la-2a.html optical compressors.
-We have also collected data from two software VA plugins: the Helper Saturator  [^8] https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ  [^9] https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
+We have also collected data from two software VA plugins using  (Fasciani et al.,, 2024b): the Helper Saturator  [^7] https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ  [^8] https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
 Overdrive
 Both the overdrive devices feature the following control parameters: volume, distortion level, and tone.
-The OD300 is a distortion pedal typically used with electric guitars, which features two different types of distortion, labeled overdrive and distortion modes.
-To record the dataset the unit was set to overdrive mode.
-The device was sampled at $5$ different and equally spaced values for the distortion and tone knobs, between the minimum and maximum values, resulting in a $25$ different combination.
-A separate output signal signal is recorded for each combination of the control parameters.
+The OD300 is a distortion pedal that features two different types of distortion, labeled overdrive and distortion modes.
+To record the dataset, the unit was set to overdrive mode.
+The device was sampled at $5$ different and equally spaced values for the distortion and tone knobs.
 Neutron’s overdrive is a module integrated into the Neutron synthesizer.
@@
 We recorded the dataset’s varying distortion and tone as we did for the OD300, but sampling $10$ equally spaced values for each parameter.
-In this case, we recorded the output for $100$ different combinations of control parameters.
-In both devices, the volume level controls an attenuator, and we fixed this parameter to the maximum (i.e.
-no attenuation).
+In both devices, the volume level controls an attenuator, and we fixed this parameter to the maximum (i.e., no attenuation).
 Filter
@@
 The filter can be set as low-pass, high-pass, and band-pass.
-We isolated the module and recorded the dataset at $10$ equally spaced values for the cutoff and $5$ for the resonance.
-The cutoff parameter is between the minimum and maximum, while the resonance parameter is between the minimum to the mid value.
-having $100$ different combinations and selecting the low-pass filter type.
+We isolated the module and recorded the dataset at $10$ equally spaced values for the cutoff and $5$ for the resonance, selecting the low-pass filter type.
 The filter was disconnected from the ADSR envelope module.
@@
 The software presents four parameters: lowpass and highpass cutoffs, saturation type switch (tape or tube), and saturation level.
-We recorded the dataset for $10$ different values of the saturation level between minimum and maximum.
+We recorded the dataset for $10$ different values of the saturation level.
 The saturation switch was set to tape mode.
 Lowpass and highpass cutoffs were set to $20$ and $20000$ Hz, respectively.
-The other device under this category is the TC Electronic Bucket Brigade Analog Delay pedal.
-The pedal is equipped with the well-known bucket brigade circuit, which consists of a series of capacitors that transmit the signal.
-By regulating the speed at which the capacitors fill and empty, a clock controls and slows down the signal, resulting in a delay as it passes through the bucket brigade device (BBD) circuit.
-The signal is sampled by storing an electric charge in a capacitor at each clock pulse.
-The sampling creates harmonic content and saturation.
-After traversing the BBD, the signal undergoes another low-pass filtering process, attenuating the new harmonics, and is expanded to match the dynamic range of the original input signal as closely as possible.
-Before entering the BBD section, the signal is typically manipulated as well, compressing and filtering using an anti-aliased filter.
-This device features four knobs: time, feedback, volume, and depth, as well as a mode switch.
-The time knob determines the length of the delay, adjustable up to $600$ ms.
-The feedback knob controls the number of repetitions of the delayed signal.
-The volume knob adjusts the volume of the delayed signal, while the depth knob defines the level of modulation.
-Activating the mode switch enables the modulation effect.
-For our study, we primarily want to capture the processed signal without the delay component, which can be digitally simulated  [60].
-All knobs were set to their minimum values except for the volume knob, while the mode switch remained deactivated.
 Equalizer
-The software plug-in Universal Audio Pultec Passive EQ emulates the tube-based equalizer Pultec EQP-1A.
+The software plug-in Universal Audio Pultec Passive Equalizer emulates the tube-based equalizer Pultec EQP-1A.
 The effect presents low-frequency ([20, 30, 60, 100] Hz) and high-frequency ([3, 4, 5, 8, 10, 12, 16]kHz) switches to select the frequencies to boost or cut using the respective boost and attenuator knobs ([0,10]).
@@
 The equalizer boosts and cuts the selected frequency bands, giving some saturation simultaneously.
-In this case, we recorded the dataset varying the boost and attenuator knobs at $5$ different equally-spaced values up to half of their range, setting $60$ Hz as the low-frequency switch value and 00 (sharp) as the bandwidth, obtaining $25$ different recordings.
+In this case, we recorded the dataset varying the boost and attenuator knobs at $5$ different equally-spaced values up to half of their range, setting $60$ Hz as the low-frequency switch value and 00 (sharp) as the bandwidth.
 Compressor/Limiter
@@
 The CL 1B compressor dataset presents four parameters: attack, release, ratio, and threshold.
-We consider the threshold values to range from 00 to $40$ dBu, ratio values from $2$:$1$ to $10$:$1$, attack time from $0.5$ to $300$ ms, and release time from $0.05$ to $10$ seconds.
 We consider three equally spaced values for the ratio, attack, and release time and four values for the threshold.
-The resulting number of parameter combinations is $108$.
 On the other hand, LA-2A presents the peak reduction knob indicating the amount of compression ([00,$100$]) and a mode switch (limiter, compression).
-We consider ten equally spaced values for the peak reduction, resulting in $20$ parameter combinations.
+We consider 10 equally spaced values for the peak reduction.
 Table 2: Datasets with related parameters and ranges considered in the study.
 Type Parameters Range Combinations
-OD300111Hardware. Overdrive Level [min, max] $25$
+OD300111Recorded from hardware devices. Overdrive Level [min, max] $25$
 Tone [min, max]
-Neutron’s Overdrive Level [min, max] $100$
-OD module [^1] Hardware. Tone [min, max]
-Neutron’s Low pass Cutoff [min, max] $100$
-Filter module [^1] Hardware. filter Resonance [min, max]
+Neutron Overdrive Level [min, max] $100$
+Overdrive module [^1] Recorded from hardware devices. Tone [min, max]
+Neutron Low pass Cutoff [min, max] $100$
+Filter module [^1] Recorded from hardware devices. filter Resonance [min, max]
 Helper Saturator Saturation [min, max] $10$
-Saturator [^2] Software. Saturation Type Tape
+Saturator [^2] Recorded from software devices. Saturation Type Tape
 Pultec Equalizer Low frequency 60 Hz $25$
-Passive EQ  [^2] Software. Frequency booster [0,10]
+Passive EQ  [^2] Recorded from software devices. Frequency booster [0,10]
 Bandwidth 0
 Frequency Attenuator [0,10]
-LA-2A [^1] Hardware.  [^3] Downloaded. Optical Peak Reduction [0,100] $20$
+LA-2A [^1] Recorded from hardware devices.  [^3] Publicly available. Optical Peak Reduction [0,100] $20$
 Compressor Switch Mode [Compressor,Limiter]
-CL 1B [^1] Hardware. Optical Threshold [-40, 0] dB $108$
+CL 1B [^1] Recorded from hardware devices. Optical Threshold [-40, 0] dB $108$
 Compressor Ratio 1:[1, 10]
@@
 Release [0.005, 10] s
-[…]
 3.3.1 Data Collection
@@
 The right input channel of the audio interface was connected to the output of the device, and the input of the device to the right output channel of the audio interface.
-This allows recording effectively both the device’s input and output signals compensating for the minor sound coloring and latency of the audio interface.
+This allows effective recording of both the device’s input and output signals, compensating for the minor sound coloring and latency of the audio interface.
 The audio data was recorded at a sampling rate of $48$ kHz.
@@
 3.4 Experimenting and Learning
-The models are trained for $60$ epochs and using the Adam [61] optimizer with a gradient norm scaling of $1$  [62].
+The models are trained for $200$ epochs and using the Adam (Kingma & Ba,, 2014) optimizer with a gradient norm scaling of $1$  (Pascanu et al.,, 2013).
 The training was stopped earlier in case of no reduction of validation loss for $10$ epochs.
-We design a time-based schedule for the learning rate as follows:
-$lr=\frac{LR}{e^{4}},\text{ with }LR=3\cdot 10^{4}$ (13)
-where LR is the initial learning rate, and $e$ is the number of the epochs.
+We design a learning rate with an exponential decay according to the formula: $lr=LR*0.25^{e}$, where LR is the initial learning rate and it is set to $3\cdot 10^{-4}$, and $e$ is the number of the epochs.
 Test losses and evaluation metrics are computed using the model’s weights that minimize the validation loss throughout the training epochs.
-Finally, the input signal is split into segments of $2400$ samples (equivalent to $50$ ms), to be processed before updating the weights.
+Finally, the input signal is split into segments of $2400$ samples (equivalent to $50$ ms) to be processed before updating the weights.
+Considering the network architecture described in Section 3.1, in which a buffer of the $64$ most recent samples is used as input, training layers with states necessitates sequential input-output examples.
+Consequently, our implementation adopts an input shape of ($B$, $L$, $F$), which is set to ($B$, $1$, $64$), where $B$ denotes the arbitrary batch size, $L$ the sequence length, and $F$ the input dimensionality, often identified as the number of input features.
+This configuration considers the $64$ input samples as the vector of input features at the current sampling instant.
 We adopted the same learning schedule and minibatch size for all models across all datasets.
@@
 Different datasets and tasks may require slightly different learning rate schedules; an excessively slow or fast learning rate can lead to suboptimal solutions.
-The number of samples to process before updating the weights, referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
+The number of samples to process before updating the weights referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
 The composition of the training, validation, and test sets is identical across datasets.
@@
 4 Results
-The evaluation is based on the metrics described in Section 3.2 and presented in Table 3, and 4.
+The evaluation is based on the metrics described in Section 3.2 and presented in Table 4.
 The MSE loss and the epoch-related epoch resulting in lower validation loss during training are reported in Table 3.
@@
 Regarding the low-pass filter, the parameters are configured for minimum cutoff and maximum resonance.
-The plots show the RMSE in Figure  4, the spectral flux in Figure  5, the spectra in Figure  7, and the spectrograms in Figure  6.
+The plots show the RMSE in Figure  4, and the spectrograms in Figure  5.
 For datasets including conditioning parameters, the plots showing predictions versus the target are generated for a $2.5$ seconds input signal containing bass, guitar, and drum loop sound.
@@
 To ensure a clearer comparative visualization of the spectral flux, we use only the first $1.25$ seconds.
+Table 3: MSE test loss computed on the test set. The test sets include all conditioning values, if applicable. The last column reports the epoch at which the training met the early stopping condition.
+Dataset Model MSE Epoch Dataset Model MSE Epoch
+OD300 LSTM $\boldsymbol{2.46\cdot 10^{-4}}$ $144$ Neutron LSTM $\boldsymbol{2.27\cdot 10^{-3}}$ $4$
+Overdrive ED $3.57\cdot 10^{-4}$ $176$ Overdrive ED $2.79\cdot 10^{-3}$ $3$
+LRU $2.98\cdot 10^{-4}$ $170$ LRU $2.43\cdot 10^{-3}$ $100$
+S4D $3.23\cdot 10^{-4}$ $165$ S4D $2.62\cdot 10^{-3}$ $8$
+S6 $3.39\cdot 10^{-4}$ $172$ S6 $2.41\cdot 10^{-3}$ $28$
+Helper LSTM $4.56\cdot 10^{-4}$ $165$ Pultec LSTM $\boldsymbol{7.41\cdot 10^{-4}}$ $19$
+Saturator ED $4.41\cdot 10^{-4}$ $138$ Passive ED $1.11\cdot 10^{-3}$ $101$
+LRU $\boldsymbol{4.14\cdot 10^{-4}}$ $176$ Equalizer LRU $1.03\cdot 10^{-3}$ $150$
+S4D $4.40\cdot 10^{-4}$ $149$ S4D $1.06\cdot 10^{-3}$ $11$
+S6 $4.53\cdot 10^{-4}$ $188$ S6 $9.68\cdot 10^{-4}$ $8$
+LA-2A LSTM $1.11\cdot 10^{-4}$ $59$ CL 1B LSTM $1.78\cdot 10^{-4}$ $74$
+Optical ED $9.54\cdot 10^{-5}$ $90$ Optical ED $\boldsymbol{8.96\cdot 10^{-5}}$ $38$
+Compressor LRU $1.14\cdot 10^{-4}$ $146$ Compressor LRU $6.76\cdot 10^{-4}$ $168$
+S4D $8.43\cdot 10^{-5}$ $62$ S4D $1.22\cdot 10^{-4}$ $14$
+S6 $\boldsymbol{8.25\cdot 10^{-5}}$ $42$ S6 $1.01\cdot 10^{-4}$ $19$
+Neutron LSTM $4.07\cdot 10^{-3}$ $46$
+Filter ED $3.36\cdot 10^{-3}$ $9$
+LRU $\boldsymbol{2.43\cdot 10^{-3}}$ $188$
+S4D $2.75\cdot 10^{-3}$ $12$
+S6 $3.16\cdot 10^{-3}$ $7$
+Table 4: ESR, $M_{RMSE}$, $M_{SF}$, and $M_{STFT}$ metrics computed on the test set. The test sets include all conditioning values, if applicable.
+Dataset Model ESR $M_{RMSE}$ $M_{SF}$ $M_{STFT}$ $M_{MFCC}$
+OD300 LSTM $\boldsymbol{1.06\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $1.60\cdot 10^{-2}$ $3.56\cdot 10^{-1}$ $6.08\cdot 10^{-1}$
+Overdrive ED $1.54\cdot 10^{-1}$ $1.30\cdot 10^{-2}$ $1.19\cdot 10^{-2}$ $3.52\cdot 10^{-1}$ $5.80\cdot 10^{-1}$
+LRU $2.35\cdot 10^{-1}$ $1.65\cdot 10^{-2}$ $1.77\cdot 10^{-2}$ $4.55\cdot 10^{-1}$ $1.05$
+S4D $1.40\cdot 10^{-1}$ $\boldsymbol{1.20\cdot 10^{-2}}$ $\boldsymbol{8.50\cdot 10^{-3}}$ $3.11\cdot 10^{-1}$ $\boldsymbol{5.38\cdot 10^{-1}}$
+S6 $1.46\cdot 10^{-1}$ $1.25\cdot 10^{-2}$ $1.04\cdot 10^{-2}$ $\boldsymbol{2.78\cdot 10^{-1}}$ $5.61\cdot 10^{-1}$
+Neutron LSTM $\boldsymbol{2.50\cdot 10^{-1}}$ $\boldsymbol{1.86\cdot 10^{-2}}$ $5.74\cdot 10^{-3}$ $4.80\cdot 10^{-1}$ $3.49\cdot 10^{-1}$
+Overdrive ED $3.07\cdot 10^{-1}$ $2.36\cdot 10^{-2}$ $6.33\cdot 10^{-3}$ $5.64\cdot 10^{-1}$ $3.80\cdot 10^{-1}$
+LRU $2.70\cdot 10^{-1}$ $1.94\cdot 10^{-2}$ $6.64\cdot 10^{-3}$ $\boldsymbol{3.40\cdot 10^{-1}}$ $3.39\cdot 10^{-1}$
+S4D $2.88\cdot 10^{-1}$ $2.21\cdot 10^{-2}$ $5.71\cdot 10^{-3}$ $5.20\cdot 10^{-1}$ $3.59\cdot 10^{-1}$
+S6 $2.65\cdot 10^{-1}$ $1.89\cdot 10^{-2}$ $\boldsymbol{4.97\cdot 10^{-3}}$ $3.86\cdot 10^{-1}$ $\boldsymbol{3.05\cdot 10^{-1}}$
+Helper LSTM $2.67\cdot 10^{-1}$ $1.05\cdot 10^{-2}$ $2.78\cdot 10^{-2}$ $5.97\cdot 10^{-1}$ $4.72\cdot 10^{-1}$
+Saturator ED $2.58\cdot 10^{-1}$ $1.06\cdot 10^{-2}$ $1.29\cdot 10^{-2}$ $5.33\cdot 10^{-1}$ $4.14\cdot 10^{-1}$
+LRU $3.46\cdot 10^{-1}$ $\boldsymbol{9.72\cdot 10^{-3}}$ $1.42\cdot 10^{-2}$ $5.44\cdot 10^{-1}$ $\boldsymbol{2.46\cdot 10^{-1}}$
+S4D $\boldsymbol{2.57\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $\boldsymbol{1.22\cdot 10^{-2}}$ $5.41\cdot 10^{-1}$ $3.67\cdot 10^{-1}$
+S6 $2.65\cdot 10^{-1}$ $1.01\cdot 10^{-2}$ $1.31\cdot 10^{-2}$ $\boldsymbol{4.28\cdot 10^{-1}}$ $3.88\cdot 10^{-1}$
+Pultec LSTM $\boldsymbol{2.15\cdot 10^{-1}}$ $1.29\cdot 10^{-2}$ $8.85\cdot 10^{-3}$ $9.33\cdot 10^{-1}$ $2.06\cdot 10^{-1}$
+Passive ED $3.22\cdot 10^{-1}$ $1.24\cdot 10^{-2}$ $2.49\cdot 10^{-3}$ $4.94\cdot 10^{-1}$ $1.91\cdot 10^{-1}$
+Equalizer LRU $3.00\cdot 10^{-1}$ $\boldsymbol{1.16\cdot 10^{-2}}$ $1.99\cdot 10^{-3}$ $\boldsymbol{1.73\cdot 10^{-1}}$ $9.81\cdot 10^{-2}$
+S4D $3.08\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $\boldsymbol{1.79\cdot 10^{-3}}$ $4.09\cdot 10^{-1}$ $\boldsymbol{9.44\cdot 10^{-2}}$
+S6 $2.80\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $3.89\cdot 10^{-3}$ $2.87\cdot 10^{-1}$ $1.26\cdot 10^{-1}$
+LA-2A LSTM $1.03\cdot 10^{-1}$ $4.51\cdot 10^{-3}$ $1.12\cdot 10^{-2}$ $4.37\cdot 10^{-1}$ $3.10\cdot 10^{-1}$
+Optical ED $8.91\cdot 10^{-2}$ $4.10\cdot 10^{-3}$ $7.28\cdot 10^{-3}$ $2.30\cdot 10^{-1}$ $2.30\cdot 10^{-1}$
+Compressor LRU $1.06\cdot 10^{-1}$ $4.67\cdot 10^{-3}$ $7.84\cdot 10^{-3}$ $2.70\cdot 10^{-1}$ $2.93\cdot 10^{-1}$
+S4D $7.87\cdot 10^{-2}$ $\boldsymbol{3.51\cdot 10^{-3}}$ $1.00\cdot 10^{-2}$ $2.52\cdot 10^{-1}$ $2.53\cdot 10^{-1}$
+S6 $\boldsymbol{7.71\cdot 10^{-2}}$ $3.63\cdot 10^{-3}$ $\boldsymbol{6.26\cdot 10^{-3}}$ $\boldsymbol{1.84\cdot 10^{-1}}$ $\boldsymbol{2.12\cdot 10^{-1}}$
+CL 1B LSTM $2.92\cdot 10^{-1}$ $9.33\cdot 10^{-3}$ $5.21\cdot 10^{-3}$ $3.56\cdot 10^{-1}$ $2.26\cdot 10^{-1}$
+Optical ED $\boldsymbol{1.46\cdot 10^{-1}}$ $\boldsymbol{4.05\cdot 10^{-3}}$ $\boldsymbol{4.18\cdot 10^{-3}}$ $4.17\cdot 10^{-1}$ $\boldsymbol{2.09\cdot 10^{-1}}$
+Compressor LRU $1.88\cdot 10^{-2}$ $6.38\cdot 10^{-3}$ $5.49\cdot 10^{-3}$ $4.38\cdot 10^{-1}$ $4.52\cdot 10^{-1}$
+S4D $2.00\cdot 10^{-1}$ $5.02\cdot 10^{-3}$ $6.12\cdot 10^{-3}$ $\boldsymbol{2.79\cdot 10^{-1}}$ $2.93\cdot 10^{-1}$
+S6 $1.67\cdot 10^{-1}$ $4.41\cdot 10^{-3}$ $6.23\cdot 10^{-3}$ $4.99\cdot 10^{-1}$ $2.36\cdot 10^{-1}$
+Neutron LSTM $1.06$ $3.15\cdot 10^{-2}$ $3.03\cdot 10^{-2}$ $9.33\cdot 10^{-1}$ $6.57\cdot 10^{-1}$
+Filter ED $8.77\cdot 10^{-1}$ $2.18\cdot 10^{-2}$ $2.21\cdot 10^{-2}$ $9.38\cdot 10^{-1}$ $6.32\cdot 10^{-1}$
+LRU $\boldsymbol{6.58\cdot 10^{-1}}$ $\boldsymbol{1.66\cdot 10^{-2}}$ $\boldsymbol{1.95\cdot 10^{-2}}$ $\boldsymbol{7.51\cdot 10^{-1}}$ $\boldsymbol{5.98\cdot 10^{-1}}$
+S4D $7.18\cdot 10^{-1}$ $1.97\cdot 10^{-2}$ $2.00\cdot 10^{-2}$ $9.01\cdot 10^{-1}$ $6.28\cdot 10^{-1}$
+S6 $8.23\cdot 10^{-1}$ $2.17\cdot 10^{-2}$ $2.18\cdot 10^{-2}$ $8.94\cdot 10^{-1}$ $6.27\cdot 10^{-1}$
 Figure 4: Comparison of the RMSE between the target and predicted output.
@@
 When the model includes conditioning, the results refer to the highest conditioning values, representing the most challenging scenario.
-In the OD300 overdrive dataset, the LSTM model presents the lowest MSE, MAE, and ESR, indicating better general amplitude matching between target and prediction.
-The S4D model is also accurate, especially when considering the transients and the RMS envelope.
-This can be observed from the lower error in terms of RMSE and spectral flux.
-On the other hand, in the case of maximum distortion, as seen in Figures 4 and 5, this superiority is less perceived.
-ED and LRU are left behind except when looking $M_{SF}$ where ED show better performance than LSTM.
-Notably, the spectrograms shown in Figure 6 illustrate that S4D model generates more harmonics in the higher frequency region than LSTM, and especially than LRU whose produces a cleaner signal.
-This could be perceptually advantageous compared to the other model that poorly predicts such region of the spectrum, as summarized by the $M_{MFCC}$ errors.
-From Figure 7 it is evident how the LSTM-based produces the best matching for energy at low frequencies.
-Figure 5: Comparison of the spectral flux between the target and predicted output. The results refer to the highest conditioning values, representing the most challenging scenario.
-Models trained on the Neutron OD dataset exhibit larger errors, suggesting that this dataset presents a more challenging modeling case for the architectures being investigated.
-Both LSTM and LRU models yield similar results in terms of average amplitude and frequency errors.
-The LSTM models show slightly better performance in the time domain, while the LRU excel in the frequency domain.
-As observed in Figure 4, both LSTM and LRU model achieves a better match with the target.
-However, the LSTM model tends to produce more energy, whereas the LRU model yields less energy than the target.
-The S4D model surpasses previous models in terms ofSF, despite exhibiting a larger mismatch at maximum distortion settings, as visible in Figure  5.
-The spectrograms indicate that the S4D model exhibits a slightly better match than the LSTM model, even though the latter exhibits smaller errors.
-Upon informal listening, the S4D model tends to add less distortion necessary, whereas the LSTM-based models tend to do the opposite.
-Lastly, the LRU model exhibits a later stopping epoch than the other three cases.
+In the OD300 and Neutron’s overdrive dataset, the LSTM model presents the lowest MSE and ESR, indicating better general amplitude matching between target and prediction.
+On the other hand, the S4D and S6 present better accuracy when considering the transients, the RMS envelope, and frequency-related content.
+This can be observed from the general lower error in terms of RMSE, spectral flux, STFT, and MFCC.
+In the case of maximum distortion, we can notice LRU prediction having significantly lower energy than other models.
+Notably, the spectrograms shown in Figure 5 illustrate that S4D and S6 models better match the energy in the harmonics at higher frequency regions.
+LRU fails to generate the harmonics in the OD300 case while ED presents the opposite behavior in Neutron’s overdrive.
+Models trained on the Neutron OD dataset exhibit larger errors, suggesting that this dataset presents a more challenging modeling case for the investigated architectures.
+This challenge can also be seen in the figures.
+The overdrive appears to be cutting the higher frequencies, behaving also as a low-pass filter.
 Figure 5: Comparison of spectrograms of the target and predicted output.
@@
 The results refer to the highest conditioning values, representing the most challenging scenario.
-For models trained on the Saturator dataset, the LRU model presents lower MSE, MAE, and $M_{RMSE}$ values.
-Meanwhile, the S4D model outperforms the others with respect to ESR and $M_{SF}$ metrics.
-The ED model is slightly less accurate but performs similarly to S4D.
-In terms of frequency domain errors, LRU also produces better results.
-As evidenced in Figures  4 and  5, the S4D and ED models achieve closer matching to the target, whereas the LRU model exhibits significantly lower energy, and the LSTM model introduces artifacts, which can be clearly heard when listening to the audio output and negatively affect its performance.
-These artifacts can also be seen in Figure 5 and  6.
-Despite these issues, the S4D and ED models still provide a better match with the target spectrograms.
-The same trend is observed in Figure 7, suggesting that the LRU model performs better in scenarios where minimal sound alteration is required.
-[…]
-Figure 7: Comparison of the frequency content between the target and predicted output.
-The FFT is computed using windows of $512$ samples.
-The results refer to the highest conditioning values, representing the most challenging scenario.
+For models trained on the Saturator dataset, the LRU model presents lower MSE and $M_{RMSE}$ values, but when considering other metrics, no models appear clearly outperforming.
+When inspecting Figures  4, LRU is again left behind, while other models present closer performances.
+In contrast, LRU achieves closer matching in the spectrograms, visible in Figure 5.
+Other models tend to generate more harmonic, especially LSTM that introduces artifacts, which can be clearly heard when listening to the audio output and negatively affect its performance.
+When training models with the Equalizer dataset, LSTM performs better in the time domain, having lower MSE and ESR errors while performing poorly in the frequency domain.
+Figures  4 show close performance among the models, while looking at the spectrograms S6 present better matching.
 The optical compressor features a time-variant response with pronounced temporal dependencies, largely influenced by its variable attack and release parameters.
 This is especially noticeable in the case of the CL 1B compressor, which represents a more challenging task for neural network modeling.
-Among the tested models, the LSTM is the least accurate, exhibiting higher loss in both the LA-2A and CL 1B cases.
-In contrast, the S4D and ED models demonstrate superior accuracy for these tasks.
-Specifically, S4D outperforms the ED in the LA-2A case, while the textbfED excels in the CL 1B case when evaluated using time domain-based metrics.
-However, the trend reverses in the frequency domain, although ED consistently shows lower $M_{SF}$ and $M_{MFCC}$ values for both cases.
-In the LA-2A case, we observe an overall good match for RMSE in Figure 4 and for transient phases in Figure 5.
-The LSTM and LRU models display slightly more mismatches in both plots.
-Upon examining Figure  6, spectrograms related to the LRU model appear similar to the target output than those of the S4D model, despite a discernible decrease in accuracy observed in Figure 7.
-In the case of the compressor, the capability to encode past information of the signal is critical, and S4D exhibits overall better accuracy.
-The ED and LRU models are designed to improve the temporal tracking abilities of basic LSTM, as confirmed by the results for both compressor devices.
-This improvement is particularly evident with the CL 1B when setting long attack and release time through the conditioning control parameters.
+Among the tested models, the LSTM and LRU are the least accurate in the time domain metrics, exhibiting higher loss in both the LA-2A and CL 1B cases.
+In contrast, in the frequency domain, the gap is less prominent.
+The S6 and ED models demonstrate superior accuracy for these two tasks.
+Specifically, S6 outperforms the ED in the LA-2A case, while the ED excels in the CL 1B case.
+S4D presents similar performance than S6.
+In the LA-2A case, we observe an overall good match for RMSE in Figure 4.
+The LRU model displays slightly more mismatches in both plots.
+Upon examining Figure  5, spectrograms related to the S6 model have the best match with the target output, while the other models present similar mismatches when looking at the higher frequencies.
+More difficult is the CL 1B case, where LRU presents the best envelope matching in Figure 4, and Figure  5.
+In the case of the compressor, the capability to encode past information of the signal is critical, LSTM is confirmed to have a worse temporal tracking ability than the other models, which present design improvements.
+These improvements are particularly evident with the CL 1B when setting long attack and release time through the conditioning control parameters.
 These reach up to $300$ ms and $5$ seconds, respectively.
-In Figure 4 and  5, the difference across models is subtler, yet the LRU model consistently shows smaller RMS errors than other models and a more accurate match in the frequency domain as visible in Figure  6 and  7.
 Modeling the low-pass filter is inherently a more complex task than the equalizer, given its more pronounced alterations to the frequency content of the processed signal.
 Nevertheless, informal listening evaluations of the predicted outputs indicate that all models struggle with accurately emulating the filter behavior; instead of eliminating high-frequency contents, they attenuate the whole signal, as apparent in Figure 5.
-However, the LRU model outperforms the others when considering the quantitative metrics.
-The LSTM model, in particular, tends to incorrectly predict excessive energy in the lower spectrum, as shown in Figure 7.
-Table 2: MSE test loss computed on the test set. The test sets include all conditioning values, if applicable. The last column reports the epoch at which the training met the early stopping condition.
-Dataset Model MSE Epoch Dataset Model MSE Epoch
-OD300 LSTM $\boldsymbol{2.46\cdot 10^{-4}}$ $144$ Neutron’s LSTM $\boldsymbol{2.27\cdot 10^{-3}}$ $4$
-ED $3.57\cdot 10^{-4}$ $176$ OD ED $2.79\cdot 10^{-3}$ $3$
-LRU $2.98\cdot 10^{-4}$ $170$ LRU $2.43\cdot 10^{-3}$ $100$
-S4D $3.23\cdot 10^{-4}$ $165$ S4D $2.62\cdot 10^{-3}$ $8$
-LA-2A LSTM $1.11\cdot 10^{-4}$ $59$ CL 1B LSTM $1.78\cdot 10^{-4}$ $74$
-ED $9.54\cdot 10^{-5}$ $90$ ED $\boldsymbol{8.96\cdot 10^{-5}}$ $38$
-LRU $1.14\cdot 10^{-4}$ $146$ LRU $6.76\cdot 10^{-4}$ $168$
-S4D $\boldsymbol{8.43\cdot 10^{-5}}$ $62$ S4D $1.22\cdot 10^{-4}$ $14$
-Helper LSTM $4.56\cdot 10^{-4}$ $165$ Pultec LSTM $\boldsymbol{7.41\cdot 10^{-4}}$ $19$
-Saturator ED $4.41\cdot 10^{-4}$ $138$ Passive ED $1.11\cdot 10^{-3}$ $101$
-LRU $\boldsymbol{4.14\cdot 10^{-4}}$ $176$ EQ LRU $1.03\cdot 10^{-3}$ $150$
-S4D $4.40\cdot 10^{-4}$ $149$ S4D $1.06\cdot 10^{-3}$ $11$
-Neutron’s LSTM $4.07\cdot 10^{-3}$ $46$
-Filter ED $3.36\cdot 10^{-3}$ $9$
-LRU $\boldsymbol{2.43\cdot 10^{-3}}$ $188$
-S4D $2.75\cdot 10^{-3}$ $12$
-[…]
-Table 3: MAE, ESR, $M_{RMSE}$, and $M_{SF}$ metrics computed on the test set. The test sets include all conditioning values, if applicable.
-Dataset Model MAE ESR $M_{RMSE}$ $M_{SF}$
-OD300 LSTM $\boldsymbol{1.14\cdot 10^{2}}$ $\boldsymbol{1.06\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $1.60\cdot 10^{-2}$
-ED $1.40\cdot 10^{-2}$ $1.54\cdot 10^{-1}$ $1.30\cdot 10^{-2}$ $1.19\cdot 10^{-2}$
-LRU $1.73\cdot 10^{-2}$ $2.35\cdot 10^{-1}$ $1.65\cdot 10^{-2}$ $1.77\cdot 10^{-2}$
-S4D $1.29\cdot 10^{-2}$ $1.40\cdot 10^{-1}$ $\boldsymbol{1.20\cdot 10^{-2}}$ $\boldsymbol{8.50\cdot 10^{-3}}$
-Neutron’s LSTM $\boldsymbol{2.02\cdot 10^{-2}}$ $\boldsymbol{2.50\cdot 10^{-1}}$ $\boldsymbol{1.86\cdot 10^{-2}}$ $5.74\cdot 10^{-3}$
-OD ED $2.54\cdot 10^{-2}$ $3.07\cdot 10^{-1}$ $2.36\cdot 10^{-2}$ $6.33\cdot 10^{-3}$
-LRU $2.14\cdot 10^{-2}$ $2.70\cdot 10^{-1}$ $1.94\cdot 10^{-2}$ $6.64\cdot 10^{-3}$
-S4D $2.40\cdot 10^{-2}$ $2.88\cdot 10^{-1}$ $2.21\cdot 10^{-2}$ $\boldsymbol{5.71\cdot 10^{-3}}$
-Helper LSTM $1.12\cdot 10^{-2}$ $2.67\cdot 10^{-1}$ $1.05\cdot 10^{-2}$ $2.78\cdot 10^{-2}$
-Saturator ED $1.15\cdot 10^{-2}$ $2.58\cdot 10^{-1}$ $1.06\cdot 10^{-2}$ $1.29\cdot 10^{-2}$
-LRU $\boldsymbol{1.04\cdot 10^{-2}}$ $3.46\cdot 10^{-1}$ $\boldsymbol{9.72\cdot 10^{-3}}$ $1.42\cdot 10^{-2}$
-S4D $1.13\cdot 10^{-2}$ $\boldsymbol{2.57\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $\boldsymbol{1.22\cdot 10^{-2}}$
-Pultec LSTM $1.36\cdot 10^{-2}$ $\boldsymbol{2.15\cdot 10^{-1}}$ $1.29\cdot 10^{-2}$ $8.85\cdot 10^{-3}$
-Passive ED $1.40\cdot 10^{-2}$ $3.22\cdot 10^{-1}$ $1.24\cdot 10^{-2}$ $2.49\cdot 10^{-3}$
-EQ LRU $\boldsymbol{1.32\cdot 10^{-2}}$ $3.00\cdot 10^{-1}$ $\boldsymbol{1.16\cdot 10^{-2}}$ $1.99\cdot 10^{-3}$
-S4D $1.36\cdot 10^{-2}$ $3.08\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $\boldsymbol{1.79\cdot 10^{-3}}$
-LA-2A LSTM $4.77\cdot 10^{-3}$ $1.03\cdot 10^{-1}$ $4.51\cdot 10^{-3}$ $1.12\cdot 10^{-2}$
-ED $4.32\cdot 10^{-3}$ $8.91\cdot 10^{-2}$ $4.10\cdot 10^{-3}$ $\boldsymbol{7.28\cdot 10^{-3}}$
-LRU $5.10\cdot 10^{-3}$ $1.06\cdot 10^{-1}$ $4.67\cdot 10^{-3}$ $7.84\cdot 10^{-3}$
-S4D $\boldsymbol{3.70\cdot 10^{-3}}$ $\boldsymbol{7.87\cdot 10^{-2}}$ $\boldsymbol{3.51\cdot 10^{-3}}$ $1.00\cdot 10^{-2}$
-CL 1B LSTM $1.06\cdot 10^{-2}$ $2.92\cdot 10^{-1}$ $9.33\cdot 10^{-3}$ $5.21\cdot 10^{-3}$
-ED $\boldsymbol{4.25\cdot 10^{-3}}$ $\boldsymbol{1.46\cdot 10^{-1}}$ $\boldsymbol{4.05\cdot 10^{-3}}$ $\boldsymbol{4.18\cdot 10^{-3}}$
-LRU $6.59\cdot 10^{-3}$ $1.88\cdot 10^{-2}$ $6.38\cdot 10^{-3}$ $5.49\cdot 10^{-3}$
-S4D $5.19\cdot 10^{-3}$ $2.00\cdot 10^{-1}$ $5.02\cdot 10^{-3}$ $6.12\cdot 10^{-3}$
-Neutron’s LSTM $3.90\cdot 10^{-2}$ $1.06$ $3.15\cdot 10^{-2}$ $3.03\cdot 10^{-2}$
-Filter ED $2.51\cdot 10^{-2}$ $8.77\cdot 10^{-1}$ $2.18\cdot 10^{-2}$ $2.21\cdot 10^{-2}$
-LRU $\boldsymbol{2.11\cdot 10^{-2}}$ $\boldsymbol{6.58\cdot 10^{-1}}$ $\boldsymbol{1.66\cdot 10^{-2}}$ $\boldsymbol{1.95\cdot 10^{-2}}$
-S4D $2.22\cdot 10^{-2}$ $7.18\cdot 10^{-1}$ $1.97\cdot 10^{-2}$ $2.00\cdot 10^{-2}$
-[…]
-Table 4: $M_{STFT-1}$, $M_{STFT-2}$, and $M_{MFCC}$ metrics computed on the test set. The test sets include all conditioning values, if applicable.
-Dataset Model $M_{STFT-1}$ $M_{STFT-2}$ $M_{MFCC}$
-OD300 LSTM $4.53\cdot 10^{-1}$ $3.56\cdot 10^{-1}$ $6.08\cdot 10^{-1}$
-ED $4.47\cdot 10^{-1}$ $3.52\cdot 10^{-1}$ $5.80\cdot 10^{-1}$
-LRU $4.65\cdot 10^{-1}$ $4.55\cdot 10^{-1}$ $1.05$
-S4D $\boldsymbol{4.19\cdot 10^{-1}}$ $\boldsymbol{3.11\cdot 10^{-1}}$ $\boldsymbol{5.38\cdot 10^{-1}}$
-Neutron’s LSTM $5.79\cdot 10^{-1}$ $4.80\cdot 10^{-1}$ $3.49\cdot 10^{-1}$
-OD ED $6.68\cdot 10^{-1}$ $5.64\cdot 10^{-1}$ $3.80\cdot 10^{-1}$
-LRU $\boldsymbol{3.56\cdot 10^{-1}}$ $\boldsymbol{3.40\cdot 10^{-1}}$ $\boldsymbol{3.39\cdot 10^{-1}}$
-S4D $6.40\cdot 10^{-1}$ $5.20\cdot 10^{-1}$ $3.59\cdot 10^{-1}$
-Helper LSTM $6.56\cdot 10^{-1}$ $5.97\cdot 10^{-1}$ $4.72\cdot 10^{-1}$
-Saturator ED $6.65\cdot 10^{-1}$ $\boldsymbol{5.33\cdot 10^{-1}}$ $4.14\cdot 10^{-1}$
-LRU $\boldsymbol{5.62\cdot 10^{-1}}$ $5.44\cdot 10^{-1}$ $\boldsymbol{2.46\cdot 10^{-1}}$
-S4D $6.73\cdot 10^{-1}$ $5.41\cdot 10^{-1}$ $3.67\cdot 10^{-1}$
-Pultec LSTM $1.10$ $9.33\cdot 10^{-1}$ $2.06\cdot 10^{-1}$
-Passive ED $6.74\cdot 10^{-1}$ $4.94\cdot 10^{-1}$ $1.91\cdot 10^{-1}$
-EQ LRU $\boldsymbol{1.82\cdot 10^{-1}}$ $\boldsymbol{1.73\cdot 10^{-1}}$ $9.81\cdot 10^{-2}$
-S4D $6.04\cdot 10^{-1}$ $4.09\cdot 10^{-1}$ $\boldsymbol{9.44\cdot 10^{-2}}$
-LA-2A LSTM $4.68\cdot 10^{-1}$ $4.37\cdot 10^{-1}$ $3.10\cdot 10^{-1}$
-ED $\boldsymbol{3.12\cdot 10^{-1}}$ $\boldsymbol{2.30\cdot 10^{-1}}$ $\boldsymbol{2.30\cdot 10^{-1}}$
-LRU $2.75\cdot 10^{-1}$ $2.70\cdot 10^{-1}$ $2.93\cdot 10^{-1}$
-S4D $3.19\cdot 10^{-1}$ $2.52\cdot 10^{-1}$ $2.53\cdot 10^{-1}$
-CL 1B LSTM $4.70\cdot 10^{-1}$ $3.56\cdot 10^{-1}$ $2.26\cdot 10^{-1}$
-ED $4.89\cdot 10^{-1}$ $4.17\cdot 10^{-1}$ $\boldsymbol{2.09\cdot 10^{-1}}$
-LRU $4.39\cdot 10^{-1}$ $4.38\cdot 10^{-1}$ $4.52\cdot 10^{-1}$
-S4D $\boldsymbol{2.95\cdot 10^{-1}}$ $\boldsymbol{2.79\cdot 10^{-1}}$ $2.93\cdot 10^{-1}$
-Neutron’s LSTM $9.44\cdot 10^{-1}$ $9.33\cdot 10^{-1}$ $6.57\cdot 10^{-1}$
-Filter ED $9.58\cdot 10^{-1}$ $9.38\cdot 10^{-1}$ $6.32\cdot 10^{-1}$
-LRU $\boldsymbol{7.89\cdot 10^{-1}}$ $\boldsymbol{7.51\cdot 10^{-1}}$ $\boldsymbol{5.98\cdot 10^{-1}}$
-S4D $9.13\cdot 10^{-1}$ $9.01\cdot 10^{-1}$ $6.28\cdot 10^{-1}$
-[…]
+However, when considering the quantitative metrics, the LRU model outperforms the others.
+The LSTM model, in particular, tends to incorrectly predict excessive energy on the RMS envelope, as visible in Figure 4, and to introduce artifact, as shown in Figure 5.
 Generally, performance metrics are relatively close across models across most effects examined in this study.
-Although the S4D models slightly outperform other models based on the metrics alone, it appears to be the most suitable architecture analog effect modeling, at least considering those within the scope of this study.
-The ED, LRU, and particularly the S4D models are more effective at encoding past signal information within their internal states.
+Although the S4D model slightly outperforms other models based on the metrics alone, it appears to be the most suitable architecture analog effect modeling, at least considering those within the scope of this study.
+The S6 model often performs very closely, while LRU presents the bigger variance across the different metrics and datasets.
+The ED shows its quality when modeling the compressors.
+The ED, LRU, and particularly the S4D and S6 models are more effective at encoding past signal information within their internal states.
 This is suggested by the case of the cases of optical compressors, where they significantly outperform other models.
 Conversely, the LSTM models appear to perform well only when emulating effects with shorter and less critical time dependencies, such as overdrive, equalizer, and filters.
-Moreover, LSTM models tend to predict energy in the low-frequency spectrum more accurately.
+Moreover, LSTM models tend to introduce artifacts in some cases.
 5 Conclusion
-The computational challenges of simulating the nonlinearities of analog circuits, described by complex differential equations, are a significant obstacle for real-time applications.
+The computational challenges of simulating the nonlinearities of analog effect, described by complex differential equations, are a significant obstacle for real-time applications.
 The complexity of these nonlinearities makes them difficult to model accurately, but they also contribute to the unique sound characteristics of the devices, making them appealing for both academic and industrial purposes.
@@
 This feature enables more interactive solutions as their memory capacity is not solely dependent on the input amount.
-In this paper, we present a comparative analysis that examines the performance of the widely used Long Short-Term Memory (LSTM) models against recently proposed recurrent models, namely Linear Recurrent Unit (LRU) and State-Sapce models (SSM), which have demonstrated remarkable results in sequence modeling.
-To carry out this study, we design four models that incorporate parameter conditioning and emphasize low-latency response.
+In this paper, we present a comparative analysis that examines the performance of the widely used Long Short-Term Memory (LSTM) models against recently proposed recurrent models, namely Linear Recurrent Unit (LRU) and State-Space models (SSM), which have demonstrated remarkable results in sequence modeling.
+To carry out this study, we design five models incorporating parameter conditioning and providing low-latency response.
 The control parameters are integrated into the models using the Feature-wise Linear Modulation (FiLM) method.
@@
 LSTM-based networks excel in replicating distortion and equalizer effects, showing their ability to encode the audio dependencies enough for these specific effects effectively.
-On the other hand, LSTM, when designed in an encoder-decoder fashion, LRUs and, in particular, SSMs, outperform other models when emulating saturation and compression.
+Conversely, LSTMs tend to introduce more artifacts than other models considered in this study.
+In contrast, LSTM, when designed in an encoder-decoder fashion, LRUs and, in particular, SSMs, outperform other models when emulating saturation and compression.
 In the case of compression, they demonstrate their advantage in accurately capturing the signal’s history.
 Generally, SSMs have also proved to be an appropriate candidate technique in VA, although the gap with the other models is not always so pronounced.
+Lastly, LRUs are less consistent in their quality across the different metrics.
 In this study, we did not carry out architecture—and effect-specific optimizations, which can definitely boost the accuracy, such as tuning hyperparameters and changing the size of the various layers.
 We limited computational complexity and input-output latency and focused the study on models that can be theoretically implemented in consumer-grade devices for live audio applications.
-Conditioning techniques and performances should be further investigated in future work.
-The current conditioning blocks rely on affine transformation.
-Still, devices such as distortion or saturation could benefit from more complex transformations to account for the possible nonlinear influence of the parameters.

### v6

@@
 Abstract
-Analog electronic circuits are at the core of an important category of musical devices, which includes a broad range of sound synthesizers and audio effects.
-The nonlinear features of their passive and active electronic components give analog musical devices a distinctive timbre and sound quality, making them highly desirable.
-The development of software that simulates analog musical devices, known as virtual analog modeling, is a significant sub-field in audio signal processing.
-Artificial neural networks are a promising technique for virtual analog modeling.
-They have rapidly gained popularity for the emulation of analog audio effects, particularly recurrent networks.
-While neural approaches have been successful in accurately modeling distortion circuits, they require architectural improvements that account for parameter conditioning and low-latency response.
-Although hybrid solutions can offer advantages, black-box approaches can still be advantageous in some contexts.
+Artificial neural networks are a promising technique for virtual analog modeling, having shown particular success in emulating distortion circuits.
+Despite their potential, enhancements are needed to enable effect parameters to influence the network’s response and to achieve a low-latency output.
+While hybrid solutions, which incorporate both analytical and black-box techniques, offer certain advantages, black-box approaches, such as neural networks, can be preferable in contexts where rapid deployment, simplicity, or adaptability are required, and where understanding the internal mechanisms of the system is less critical.
 In this article, we explore the application of recent machine learning advancements for virtual analog modeling.
-In particular, we compare State-Space models and Linear Recurrent Units against the more common Long Short-Term Memory networks.
-These have shown promising ability in sequence-to-sequence modeling tasks, showing a notable improvement in signal history encoding.
-Our comparative study uses these black-box neural modeling techniques with a variety of audio effects.
+In particular, we compare State-Space models and Linear Recurrent Units against the more common Long Short-Term Memory networks, with a variety of audio effects.
 We evaluate the performance and limitations of these models using multiple metrics, providing insights for future research and development.
-Our metrics aim to assess the models’ ability to accurately replicate energy envelopes and frequency contents, with a particular focus on transients in the audio signal.
-To incorporate control parameters into the models, we employ the Feature-wise Linear Modulation method.
-Long Short-Term Memory networks exhibit better accuracy in emulating distortions and equalizers, while the State-Space model, followed by Long Short-Term Memory networks when integrated in an encoder-decoder structure, and Linear Recurrent Unit outperforms others in emulating saturation and compression.
-When considering long time-variant characteristics, the State-Space model demonstrates the greatest capability to track history.
-The Long Short-Term Memory networks tend to introduce audio artifacts.
+Our metrics aim to assess the models’ ability to accurately replicate the signal’s energy and frequency contents, with a particular focus on transients.
+The Feature-wise Linear Modulation method is employed to incorporate effect parameters that influence the network’s response, enabling dynamic adaptability based on specified conditions.
+Experimental results suggest that Long Short-Term Memory networks offer an advantage in emulating distortions and equalizers, although performance differences are sometimes subtle yet statistically significant.
+On the other hand, encoder-decoder configurations of Long Short-Term Memory networks and State-Space models excel in modeling saturation and compression, effectively managing the dynamic aspects inherent in these effects.
+However, no models effectively emulate the low-pass filter, and Linear Recurrent Units show inconsistent performance across various audio effects.
 1 Introduction
 The emulation of analog musical devices has become pervasive in digital products due to their distinct sound, appealing to producers and musicians.
-Analog circuits use components like operational amplifiers and diodes, whose nonlinear behaviors contribute to creating a unique sound.
-Replicating these devices, called Virtual Analog (VA) modeling, is now a crucial area in digital audio signal processing.
+Analog circuits use components such as operational amplifiers and diodes whose non-linear behaviors contribute to creating a unique sound.
+Replicating these devices, called Virtual Analog (VA) modeling, is a crucial area in digital audio signal processing.
 There are two possible approaches for emulating analog audio effects: physics-based and data-driven.
@@
 Data-driven approaches identify a discrete model directly from discrete measurement data.
-Machine learning, as a data-driven method, has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks.
+Machine learning (ML), as a data-driven method, has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks.
 Hybrid solutions, incorporating knowledge of the modeling tasks, also exist and can offer advantages, but data-driven approaches can still be appealing in some contexts.
 While the former may not always be accurate enough, the latter may not be feasible for certain tasks.
-Data-driven models, also called Black-box, being applicable to different audio modeling tasks, have already been proposed  (Ramírez et al.,, 2020a), but these usually involve large networks that are not suitable for low-latency and real-time applications.
+Data-driven models, also called Black-box models, have been utilized for different audio modeling tasks  Ramírez et al., (2020a), but these usually involve large networks that are not suitable for low-latency and real-time applications.
 In addition, while machine learning modeling of nearly all types of analog effects has been investigated, they often do not provide a satisfactory response when continuously manipulating conditioning parameters, which is needed to offer users runtime control over the effect’s parameters.
-This article presents a comparative study that applies recurrent and state-based neural network architectures to model audio effects with different complexities.
-Neural networks utilizing internal states and a recurrent formulation are often preferred among the different types of artificial neural networks.
-The most known type is the Recurrent Neural Networks (RNNs).
-A Similar formulation is present in the State-Space Networks (SSMs).
-The internal states function as memory, making them particularly suitable in scenarios where time dependencies are crucial.
-Especially when the inputs are limited to one or a small segment of digital audio samples.
-For this reason, in this study, we consider Convolutional Neural Networks (CNNs) (LeCun et al.,, 1998) only when integrated with RNNs  (Simionato & Fasciani,, 2022).
-The aim is to work with small input vectors while also considering long-term temporal dependencies.
-Therefore, we focus on state-based structures.
-Analog audio effects are dynamic physical systems that exhibit time-dependent characteristics: their output depends on previous internal states.
-Therefore, dynamic physical systems possess internal states that describe their physical evolution, similar to stateful machine learning algorithms.
-We designed five architectures based on the popular Long Short-Term Memory (LSTM)  (Hochreiter & Schmidhuber,, 1997), the State-Space models (SSMs)  (Gu et al.,, 2022b; Gu & Dao,, 2023), and the Linear Recurrent Unit (LRU)  (Orvieto et al.,, 2023).
-All models consider conditioning on control parameters using the Feature-wise Linear Modulation (FiLM) method  (Perez et al.,, 2018) and employ the Gated Linear Unit (GLU)  (Dauphin et al.,, 2017).
-The architectures have a limited computational complexity, measured in the number of floating points operations per output audio sample, ensuring theoretical real-time computational capabilities on consumer-grade devices.
-The designed architectures also present limited intrinsic input-output audio samples latency, which we have limited to a maximum of $64$ samples ($1.33$ ms).
-Specifically, we refer to the intrinsic latency given by the architecture constraints, neglecting the latency given by the external DSP processing.
+Neural networks incorporating internal states and utilizing recurrent formulations have emerged as the most accurate option for modeling audio effects among various types of artificial neural networks, with Recurrent Neural Networks (RNNs) being the most recognized.
+A similar formulation exists in State-Space Models (SSMs).
+The internal states in these architectures serve as memory, rendering them particularly effective in handling scenarios with transients and time dependencies, especially when processing one or a few audio samples at a time to achieve low latency.
+Consequently, in this study, we include Convolutional Neural Networks (CNNs) LeCun et al., (1998)—which are inherently stateless—only when integrated with RNNs Simionato & Fasciani, (2022).
+This article presents a comprehensive comparative study that systematically evaluates state-based neural network architectures in modeling various types of audio effects, each posing unique modeling challenges, and in the context of achieving low-latency, computationally sustainable, and parametric solutions.
+Experimental results highlight the strengths and limitations of different state-based networks in their ability to learn diverse audio effects.
+Specifically, we explore the learning capabilities of these architectures, investigating the similarities and differences between them.
+By analyzing RNNs and SSMs, which use distinct recurrence methods to compute internal states, we provide insights into their variant modeling capabilities across different contexts.
+While architecture-specific optimizations, such as hyperparameter tuning or layer size adjustments, were not pursued, our study prioritized limiting computational complexity and minimizing input-output latency.
+We focused on developing models that are suitable for implementation on consumer-grade devices for live audio applications.
+We designed five architectures based on the popular Long Short-Term Memory (LSTM)  Hochreiter & Schmidhuber, (1997), the SSMs  Gu et al., (2022b); Gu & Dao, (2023), and the Linear Recurrent Unit (LRU)  Orvieto et al., (2023).
+All models consider conditioning on control parameters using the Feature-wise Linear Modulation (FiLM) method  Perez et al., (2018) and employ the Gated Linear Unit (GLU)  Dauphin et al., (2017).
+The architectures are engineered to operate with a computational load of less than $100$ MFLOPS (Million Floating Point Operations Per Second) when processing audio at $48$ kHz.
+This constraint facilitates the simultaneous operation of multiple software audio effects on consumer-grade personal computers and supports independent execution on bare-metal embedded systems.
+Additionally, the architectures are designed to exhibit minimal inherent input-output latency for audio samples.
+Specifically, we refer to the intrinsic latency given by the architecture constraints, neglecting the latency given by the external Digital Signal Processor (DSP) processing.
 This latency is due to the lag between the time index of the input and output samples.
@@
 A numerical scheme is formulated, and the discrete equations are solved iteratively to generate the digital audio output.
-Data-driven modeling is used when the inner circuitry of the audio effect is unknown.
-It treats the model as a mapping between input-output data: input-output measurements are used to approximate the system response.
-The procedure involves three main components: the dataset, candidate modeling techniques, and the validation process.
-The quality of the resulting model heavily depends on the selection of input signals, configuration of control parameters, and data collection.
-Data-driven approaches rely on measurements that are static representations of the system behavior, leading to greater difficulty in accurately emulating its dynamic aspects.
-Common examples of data-driven approaches include the block-oriented Hammerstein model, the Wiener model, and their combinations (Schoukens & Tiels,, 2017).
-In the Wiener model, a linear dynamic system is followed by a static nonlinearity, whereas in the Hammerstein model, the static nonlinearity precedes the linear dynamic system.
-These models offer a straightforward representation and are derived by estimating the model coefficients from real data using error-minimization techniques.
-A proposed method to classify whether a Wiener or a Hammerstein model should rather model the nonlinear system is in (Primavera et al.,, 2018).
-Recent examples are for guitar distortion pedal  (Eichas et al.,, 2015)  (Köper & Holters,, 2020)  (Darabundit et al.,, 2019), amplifier  (Jr & Bodmann,, 2016)  (Eichas et al.,, 2017), distortion circuit  (Eichas & Zölzer,, 2016).
-A more recent data-driven technique involves utilizing artificial neural networks (ANNs).
-ANNs consist of interconnected units called neurons, where each connection can transmit a signal.
-Each neuron processes these signals and can send them to other connected neurons.
-The transmitted signal is represented by a real number, and the neuron’s output is determined by applying a nonlinear function to the sum of its weighted inputs.
-Training ANNs involve optimizing the network’s parameters to minimize the error between predicted and actual target values in a dataset.
+Data-driven modeling is typically used when the internal circuitry of an audio effect is unknown, treating the model as a mapping of input-output data to approximate the system’s response.
+This process involves three main components: the dataset, modeling techniques, and validation methods.
+The model’s quality is heavily dependent on the selection of input signals, control parameter configurations, and data collection strategies.
+However, because data-driven approaches often rely on static measurements, accurately capturing dynamic behavior presents a notable challenge.
+Common examples of data-driven approaches include the block-oriented Hammerstein model, the Wiener model, and their combinations  Schoukens & Tiels, (2017).
+Recent examples of their application focus on guitar distortion pedal  Eichas et al., (2015)  Köper & Holters, (2020)  Darabundit et al., (2019), amplifier  Jr & Bodmann, (2016)  Eichas et al., (2017), distortion circuit  Eichas & Zölzer, (2016).
+More recent data-driven techniques leverage artificial neural networks (ANNs), comprising interconnected nonlinear units known as neurons.
+Each neuron propagates signals through its connections, processes these signals, and has the ability to transmit the processed signals to other connected neurons.
+Training ANNs involves optimizing the network’s parameters to minimize the error between predicted and actual target values in a dataset.
 Gradient-based techniques, like backpropagation, are commonly used to optimize the network’s weights.
@@
 In the last two decades, ANNs have been extensively used for modeling analog audio devices.
-In particular, traditional ML techniques, initially developed for different purposes, have been investigated to create models that emulate audio effects when taking and processing raw audio samples, such as traditional DSP techniques.
-In an early attempt using machine learning for audio effect modeling  (Mendoza,, 2005), a multilayer feedforward network was utilized to learn a high-pass filter and the Ibanez Tube-Screamer distortion effect.
-The high-pass filter training led to unsatisfactory results, while the distortion effect demonstrated the ability of the network to learn distortion effects.
-Other attempts used convolutional-based networks to model the preamplifier circuit from the Fender Bassman 56F-A vacuum-tube amplifier  (Damskägg et al.,, 2019a).
-The network consists of a stack of dilated causal convolution and fully connected layers.
-The dataset was simulated with different values of the gain, which is a variable resistor, and used to condition the model.
-The conditioning is realized in the convolutional layers and using the gated activation  (Oord et al.,, 2016).
-The work shows that CNNs accurately emulate the reference device better than a fully connected feedforward network and can respond to user control changes.
-Using the same architecture, the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals have been modeled as well  (Damskägg et al.,, 2019b).
-Here, the authors investigated more configurations, considering different dilation patterns and activation functions.
-In this work, the conditioning is not considered.
-Similarly to a fully connected feedforward network, in which networks’ ability to track the past information depends on the amount of input received, the convolutional network performance depends on the receptive field; a longer receptive field requires a greater number of layers.
-To overcome this problem, Temporal Convolutional networks (TCN) have been proposed for optical dynamic range compressors such as Teletronix LA-2A Leveling Amplifier (Steinmetz et al.,, 2022), which present longer time-variant characteristics than distortion units.
-TCNs’ strength consists of dilated convolutional networks with rapidly growing dilation factors ($2^{2}$ and $02^{1}$ 0).
-Increasing the dilation factors increases, in turn, the receptive field and the ability of the networks when dealing with long dependencies while, in turn, not increasing the number of layers.
-The authors compared TCN networks with presenting receptive fields of $100$, $300$, and $1000$ ms, respectively.
-The model with a receptive field of $300$ ms provided the best performances.
-TCNs provide accurate and efficient implementation with respect to computational complexity.
-On the other hand, convolutional architectures present an intrinsic latency equal to the length of the receptive field, which is detrimental for live audio applications.
-A model that combines different deep learning architectures was investigated in  (Ramírez et al.,, 2019).
-In this case, the model is divided into three parts: an adaptive front-end, latent space, and synthesis back-end.
-The front-end performs time-domain convolutions with the raw audio to map it into a latent space.
-The adaptive front-end consists of a convolutional encoder.
-The latent-space DNN modifies the latent representation of the input audio into a new latent representation, which is fed into the synthesis back-end.
-The synthesis back-end consists of an unpooling layer, a deep neural network with smooth adaptive activation functions (DNN-SAAF), and a single CNN layer.
-The SAAF  (Hou et al.,, 2017) is used as the nonlinearity for the last layer.
-Distortion, overdrive, and EQ effects have been modeled using this method.
-In  (Ramírez et al.,, 2020a), the previously described architecture is further explored and compared with different variations and with the CNN architecture based on  (Damskägg et al.,, 2019a).
-The variations mainly consist of incorporating Bi-LSTMs or CNNs into the latent space.
-In this work, the target devices were Universal Audio Vacuum-Tube Preamplifier 610-B, Universal Audio Transistor-Based Limiter Amplifier 1176LN, and 145 Leslie Speaker Cabinet.
-The highest score in the listening test was obtained by the variation incorporating Bi-LSTM.
-Another variation based on Bi-LSTM and SAAF as latent-space modules is used for plate and spring reverberators  (Ramírez et al.,, 2020b).
-These architectures represent a general-purpose deep learning framework for modeling audio effects, which has been shown to be capable of accurately modeling different types of effects.
-On the other hand, the models are large networks with high computational complexity, leading to difficulty when considering efficient implementation.
-In addition, using the Bi-LSTM layer makes the architecture non-causal, not allowing real-time applications.
-This architecture also lacks conditioning possibility and focuses on modeling a static representation of the devices (i.e., fixed parameters).
-The last group of machine-learning black-box modeling methods we review in this section are those based on RNNs.
-These have been applied to model a vacuum-tube amplifier  (Covert & Livingston,, 2013), where authors investigated different networks and input sizes in two scenarios: slightly distorted and overdriven signals.
-The networks’ input included a variable number of audio samples.
-The network is autoregressive: it generates the output sample at the current time step using current and past input samples, but also output samples generated at previous time steps.
-The work presents preliminary experiments, and although the models resulted in poor performance, the work demonstrates the feasibility of modeling vacuum-tube amplifiers with RNNs.
-RNNs, in particular LSTMs, have also been compared to CNNs when applied to the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals  (Wright et al.,, 2019).
-The architecture consists of one RNN layer and one linear fully connected layer as an output layer with one unit.
-LSTM and GRU are compared as RNN layers varying the number of units.
-The input signal is extended to include an additional value representing the user control’s parameter.
-The results show that RNNs achieve similar accuracy to the CNN model while requiring significantly less computational cost and latency since requiring only the current input sample to predict the current output.
-LSTM performed better than GRU networks, and the error decreased when the number of units was increased.
-Another RNN-based model utilized for the Klon Centaur guitar pedal circuit is presented in  (Chowdhury,, 2020), which is also compared with a WDF model of the same circuit.
-ML implementation resulted in better computational performance than the WDF model but sounded very similar, although the ML implementation slightly damped high frequencies.
-The architecture consists of a GRU layer and one linear fully connected layer with one unit as the output layer.
-Also, here, a single variable control parameter of the audio effect is added as an extra input to condition the inference of the networks.
-A delay-based effect is investigated in  (Mikkonen et al.,, 2023), where a tape-based delay is modeled using GRU networks.
-Here, the delay trajectory is analyzed and extracted using impulse train signals.
-This trajectory is used to demodulate the signal before training the models or used to directly guide a delay-line built in a differentiable way and based on  (Engel et al.,, 2020) work.
-RNNs rely on their internal memory mechanism to encode the history of the signal, but they can still suffer from remembering significant long-term dependencies.
-Indeed, LSTMs alone have not been successful for LFO delay-based effects, such as flanger and phaser.
-In  (Wright & Välimäki,, 2021), authors found it beneficial to use as extra input the LFO signal representing the frequency response of a time-varying system over time.
-The LFO signal is first extracted by analyzing the recording and then used to help the LSTM during the learning process.
-The extra input has shown to be beneficial to learning the LFO-based effect and can be freely controlled after model training, although in the study is considered only static parameter values and not LFO signal changing depth and width over time.
-A sequence-to-sequence-based model, based on LSTMs, was designed and proposed for the case of another optical compressor, such as the TubeTech CL 1B compressor  (Simionato & Fasciani,, 2022), to help LSTMs’ memory.
-The network is based on two LSTM layers: one acts as an encoder, and the other acts as a decoder.
-The first takes a segment of past input samples together with the controls’ parameters, and the decoder takes the current input sample (in case of single output) or samples (in case of multi-sample output).
-The encoder, when processing the input, computes the internal states.
-These states are used as initial states for the second LSTM.
-These states represent a summary of the encoder input, which helps the decoder to infer the output.
-The sequence-to-sequence-based model is compared to LSTMs and fully connected networks, resulting in significantly better accuracy.
-The authors initially considered two control parameters, threshold, and ratio, while attack and release parameters were fixed.
-In  (Simionato & Fasciani,, 2023), the latter parameters were included.
-The model’s encoder was replaced with a convolutional layer, and the model was compared to TCNs.
-Sequence-to-sequence-based models allow for more responsiveness and interactivity but at the expense of an increasing computational cost.
-Other works consider hybrid approaches.
-In  (Parker et al.,, 2019), fully connected neural networks exploit measurements within the circuits.
-Here, the network is embedded within a discrete-time state-space system and explored for diode-based guitar distortion circuits and the lowpass filter of the Korg MS-20 synthesizer.
-The method required measurements from the devices’ circuits, which are not always accessible.
-Continuing towards gray box approaches,  (Kuznetsov et al.,, 2020) is based on the Differentiable Digital Signal Processing (DDSP) framework and applies this concept to RC lowpass filter.
-Subsequently, the idea is applied to the parametric tone section of a guitar amplifier and a nonlinear guitar overdrive circuit in (Esqueda et al.,, 2021).
-The technique learns the parameters of a circuit model from measurements of real devices using a Wiener-Hammerstein model consisting of two IIR filters and a feedforward artificial neural network.
-The IIR filters are, in this way, trainable and derived from the analysis of the circuit.
-This approach is also used in (Nercessian et al.,, 2021).
-In this case, a cascade of biquad filters is trained, employing a Fourier transform, to model the BOSS MT-2 pedal distortion pedal.
-The model includes four tunable parameters.
-Differentiable Digital Filters have been applied for the case of a diode clipper circuit (Chowdhury & Clarke,, 2022), showing higher accuracy and similar computation time when compared to traditional physics-based models.
-In (Wilczek et al.,, 2022), the authors proposed to learn ordinary differential equations governing the first-order and the second-order diode clipper using fully connected layers and RNNs.
-The formulation similarly uses the discrete-time state-space as in (Parker et al.,, 2019).
-Neural networks have also been investigated in the case of wave digital filter formulations (Darabundit et al.,, 2022).
-A tube amplifier circuit is simulated using multiport nonlinearities inside a wave digital filter and fully connected layers to learn features from a Kirchhoff-domain dataset.
-Similarly, in  (Miklanek et al.,, 2023), authors combine recurrent neural networks (RNNs) with the Kirchhoff nodal analysis to create a grey-box guitar amplifier model.
-Lastly, an example of gray box modeling in the case of the LA-2A leveling amplifier is (Wright & Välimäki,, 2022), where it is modeled starting from the structure of a traditional digital compressor but using fully connected and RNN to predict the compressor’s parameters.
+In particular, traditional ML techniques have been investigated to create models that emulate audio effects processing raw audio samples, such as traditional DSP techniques.
+In an early attempt using ML for audio effect modeling  Mendoza, (2005), a multilayer feedforward network was utilized to learn a high-pass filter and the Ibanez Tube-Screamer distortion effect.
+The high-pass filter experiments led to unsatisfactory results as the model did not learn any aspect of the filtering process, while with the distortion effect, the network demonstrated the ability to replicate the signal alteration typical of such audio effects accurately.
+Other studies utilized CNNs to model the preamplifier circuit of the Fender Bassman 56F-A vacuum-tube amplifier  Damskägg et al., (2019a).
+The proposed network features stacked dilated causal convolution and fully connected layers, trained on a synthetic dataset with varying gain values for conditioning the network through a gated activation technique  Oord et al., (2016).
+Results indicate that CNNs outperform fully connected networks in emulating the device and they can also adapt to changes in the effect’s control parameters.
+The same architecture was later applied to model the Ibanez Tube Screamer, Boss DS-1, and Electro-Harmonix Big Muff Pi distortion pedals  Damskägg et al., (2019b), exploring various configurations without considering effects parameters for conditioning the network.
+Similarly to fully connected networks, CNNs’ ability to track the input’s past information depends on their receptive field, which requires more layers when the fields are longer.
+Temporal Convolutional Networks (TCNs) use dilated convolutions with increasing dilation factors ($2^{2}$ and $2^{10}$), enhancing the receptive field and long dependency handling without increasing layer count.
+TCNs have been applied to optical dynamic range compressors, such as the Teletronix LA-2A Leveling Amplifier  Steinmetz et al., (2022), where the utilization of a $300$ ms receptive field yielded the best performance.
+TCNs provide accurate and computationally efficient implementations, but network architectures with convolutional layers present intrinsic input-output latency equal to the receptive field length, which can hinder live audio applications.
+TCNs have also been used to design Open-Amp  Wright et al., (2025), a model emulating many guitar effects.
+The model is conditioned on a learnable look-up table of embeddings and learns one embedding vector per effect.
+In this case, the length of the receptive field is not reported.
+A model combining various deep learning architectures was explored in  Ramírez et al., (2019).
+The model includes an adaptive front-end, a latent space, and a synthesis back-end.
+The front-end uses a convolutional encoder to perform time-domain convolutions on raw audio, mapping it into a latent space.
+The latent-space transforms the input audio representation, which is then processed by the synthesis back-end, consisting of an unpooling layer, a deep neural network with smooth adaptive activation functions (SAAF)  Hou et al., (2017).
+This approach has been applied to model distortion, overdrive, and equalizer effects.
+In  Ramírez et al., (2020a), the previously described architecture was further investigated and compared with variations and the CNN-based architecture from  Damskägg et al., (2019a).
+Variations included incorporating Bi-LSTMs or CNNs into the latent space.
+The Bi-LSTM variation achieved the highest score in listening tests.
+Another Bi-LSTM and SAAF variation was used for plate and spring reverberators  Ramírez et al., (2020b).
+While these architectures effectively model various audio effects, they are large networks with high computational complexity, making efficient implementation challenging.
+The use of Bi-LSTMs also introduces non-causality, hindering real-time applications and interactivity with the effect’s parameters.
+In  Comunità et al., (2023), LSTMs are integrated to enhance intermediate data representation in TCN models, tested on fuzz and compressor effects, and the proposed architecture outperforms the baselines of LSTM and CNN.
+As in  Steinmetz et al., (2022), the models present long receptive fields, in this case, between $45$ ms and $2500$ ms.
+Finally, RNNs have been used to model vacuum-tube amplifiers  Covert & Livingston, (2013).
+Although preliminary experiments showed limited accuracy, they established RNN feasibility for this modeling task.
+RNNs, particularly LSTMs, were also compared to CNNs in modeling the Ibanez Tube Screamer, Boss DS-1, and Electro-Harmonix Big Muff Pi distortion pedals  Wright et al., (2019).
+The architecture included one RNN layer and a linear output layer, with LSTM and Gated Recurrent Unit (GRU) compared by varying the number of units.
+The input of the network includes both the audio signal and the effect parameters.
+The results indicated that RNNs achieved similar accuracy to CNNs, but with lower computational costs and latency, relying only on the current input sample for output prediction.
+LSTM outperformed GRU, with error reduction observed as unit count increased.
+In  Chowdhury, (2020), another RNN-based model for the Klon Centaur guitar pedal circuit is compared with a Wave Digital Filter (WDF) model.
+The machine learning approach presented a lower computational cost, though the emulated sound was very similar to the WDF model, with a slight damping of high frequencies.
+The architecture includes a GRU layer and a linear output layer with one unit, and a single control parameter is also fed as input to the network to condition the inference.
+In  Mikkonen et al., (2023), a tape-based delay effect is modeled using GRU networks, with the delay trajectory analyzed through impulse train signals.
+This trajectory demodulates the signal before model training or guides a differentiable delay line based on  Engel et al., (2020).
+In  Wright & Välimäki, (2021), the modeling of delay-based effects, such as flanger and phaser, is investigated using an extracted low-frequency oscillation (LFO) signal to aid LSTMs during training.
+This is provided as additional input to the network, and it is shown to improve the learning of these effects.
+A the same time, it allows for post-training control of the model; however, the experiments were focused on fixed effect parameters, without changes in the LFO depth and width over time.
+An encoder-decoder (ED) network was utilized to model the TubeTech CL 1B compressor  Simionato & Fasciani, (2022), featuring an encoder LSTM that processes past input samples and control parameters, and a decoder LSTM handling current and more recent past input samples.
+The encoder’s internal states summarize the input, aiding the decoder’s output inference.
+This model outperformed standard LSTMs and fully connected networks in accuracy, initially modeling only threshold and ratio control parameters, while the attack and release settings were fixed.
+In  Simionato & Fasciani, (2023), all four compressor control parameters were modeled, and the encoder was replaced with a convolutional layer for comparison with TCNs.
+While ED models offer increased responsiveness and interactivity, they also incur higher computational costs.
+Other studies explored hybrid modeling approaches.
+In  Parker et al., (2019), fully connected neural networks utilize circuit measurements, embedding the network within a discrete-time state-space system for diode-based guitar distortion circuits and the Korg MS-20 lowpass filter.
+This method relies on circuit measurements, which may not always be accessible.
+Similarly,  Peussa et al., (2021) proposes a circuit state-matching mechanism for GRU networks, training a variant of the GRU layer to match circuit states using a teaching-forcing method.
+The Differentiable Digital Signal Processing (DDSP) framework for modeling an RC lowpass filter is employed in  Kuznetsov et al., (2020), later also used for the parametric tone section of a guitar amplifier and a nonlinear overdrive circuit in  Esqueda et al., (2021).
+This method learns circuit parameters from real device measurements using a Wiener-Hammerstein model with two trainable Infinite Impulse Response (IIR) filters and a feedforward neural network.
+Similarly,  Nercessian et al., (2021) trains a cascade of biquad filters via the Fourier transform for the BOSS MT-2 distortion pedal, incorporating four tunable parameters.
+The DDSP framework was also successfully applied to a diode clipper circuit in  Chowdhury & Clarke, (2022), demonstrating higher accuracy with comparable computation times to traditional models.
+In  Wilczek et al., (2022), the authors learn ordinary differential equations governing diode clippers using fully connected layers and RNNs, aligned with the discrete-time state-space approach from  Parker et al., (2019).
+Investigations into WDF formulations are found in  Darabundit et al., (2022), simulating tube amplifier circuits with multiport nonlinearities and fully connected layers.
+Miklanek et al., (2023) combines RNNs with Kirchhoff nodal analysis to model guitar amplifiers.
+A hybrid approach for the LA-2A leveling amplifier is presented in  Wright & Välimäki, (2022), modeled using a traditional digital compressor structure alongside fully connected RNNs for parameter prediction.
+Lastly,  Huhtala et al., (2024) proposed a Koopman-Linearised Audio Neural Network structure, which comprises two gated linear unit multilayer perceptrons and differentiable biquad filters.
 3 Methods
 This study aims to identify which type of recurrent layer or architectural configuration accurately models different types of audio effects, if any.
-These insights can be used to understand the goodness and weakness of these models and determine which can better suit each audio effect.
-The states-based architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states rather than solely relying on the input received at each iteration.
-In addition, we restrict the experimentation to relatively small networks fed by a small number of input segments, which present low computational complexity and low input-output latency, allowing implementation on real-time consumer-grade digital audio systems.
-The designed models are applied to different audio effects and finally evaluated using different metrics.
-To be noticed, the models presented here, having limited parameters, are not expected to lead to perfectly accurate emulation but a good starting point.
-In the rest of this section, we first detail the four architectures used in this comparative study; then, we present the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets.
+These insights can be used to understand the strengths and weaknesses of these models and determine which can better suit each audio effect.
+The state-based architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states rather than solely relying on the input received at each iteration.
+In addition, we restrict the experimentation to relatively small networks that are fed by a limited number of input samples.
+This approach helps reduce computational complexity and minimizes input-output latency, enabling implementation on real-time consumer-grade digital audio systems.
+The designed models are applied to various audio effects and evaluated using a range of metrics.
+It is important to note that, given their limited number of trainable parameters, these models are not expected to achieve perfect emulation but rather serve as a solid foundation for further development.
+In the rest of this section, we first detail the five architectures used in this comparative study; then, we present the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets.
 Finally, explain the process we used to train and test the models.
 3.1 Architectures
-The following architectures, based on different types of recurrent layers, are compared in this study: LSTM (a), ED (b), LRU (c), and S4D (d).
-The four architectures are illustrated in Figure 1.
-ED is based on the concept of internal state sharing  (Simionato & Fasciani,, 2022).
+The following architectures, based on different types of recurrent layers, are compared in this study: LSTM, ED, LRU, S4D, and S6.
+The five architectures are illustrated in Figure 1.
+ED is based on the concept of internal state sharing  Simionato & Fasciani, (2022).
 These architectures follow the same design, consisting of a linear, fully connected layer, a recurrent layer, and a conditioning block.
@@
 This choice represents a trade-off: on the one hand, there is the amount of past input information used to generate the current output, where more is typically better; on the other hand, there is the resulting input-output latency, which increases with the number of input samples.
-Our goal is to keep this latency below $1.33$ ms (i.e., $64$ samples at $48$ kH.).
+Our goal is to keep this latency below $1.33$ ms (i.e., $64$ samples at $48$ kHz).
 The input of all architectures is an array, including the $64$ most recent input sample, used to generate the single current output sample.
-This implies that a new input sample enters the buffer with each inference cycle, while the oldest is discarded.
+This implies that a new input sample enters the buffer with each inference cycle while the oldest is discarded.
 Therefore, the set of the $64$ most recent input samples constitutes the network input, which, after initial compression by the FC layer, is used to update the states in the recurrent layers.
@@
 Consequently, the model must perform predictions at an audio rate, one for each output audio sample.
-The models map a vector of $64$ sample in one scalar output sample, give the internal states $\boldsymbol{h}$:
+The models map a vector of $64$ samples into one scalar output sample, giving the internal states $\boldsymbol{h}$:
 $y_{n}=g(\boldsymbol{x_{n}})=g(x_{n},x_{n-1},...,x_{n-62},x_{n-63};\boldsymbol{h})$ (1)
@@
 $\boldsymbol{u_{n}}=\boldsymbol{W_{l}}\boldsymbol{x_{n}}+\boldsymbol{b_{l}}$ (2)
-where $\boldsymbol{u_{n}}$ is the linear projected vector having a size equal to the number of hidden units of the fully connected layer.
+where $\boldsymbol{u_{n}}$ is the linearly projected vector having a dimensionality equal to the number of hidden units of the fully connected layer.
 $\boldsymbol{W_{l}}$ and is the weights’ matrices, and $\boldsymbol{b_{l}}$ the vector bias terms.
-Figure 1: The four compared architectures: LSTM, ED, LRU, S4D, and S6.
+Figure 1: The five compared architectures: LSTM, ED, LRU, S4D, and S6.
 All architectures present the same structure: a fully connected layer, a recurrent layer, and a conditioning layer.
 For the case of ED, the input vector is equally split into two vectors, $\boldsymbol{x_{e}}$ and $\boldsymbol{x_{d}}$.
-The first two architectures use LSTMs as the recurrent layer, designed to improve the memory capacity of the vanilla RNN by incorporating internal mechanisms known as gates, regulating the flow of information.
+The vector $\boldsymbol{p}$ denotes the effect parameters utilized to condition the networks.
+Following the linear projections, the models are characterized by various sets of equations, contingent upon the specific type of layer employed.
+LSTM layer
+The vector $\boldsymbol{u_{n}}$ is processed using a LSTM as the recurrent layer.
+LSTMs were designed to improve the memory capacity of the vanilla RNN by incorporating internal mechanisms known as gates, which regulate the flow of information.
 The LSTM architecture is governed by the following equations:
@@
 $\boldsymbol{h_{n}}$ $=\boldsymbol{o_{n}}\circ\phi(\boldsymbol{c_{n}})$ (3)
-where $\boldsymbol{h_{n}}$ is the hidden state vector, $\boldsymbol{f}$ is the forget gate, $\boldsymbol{i}$ the input gate, and $\boldsymbol{o}$ the output gate, and $\boldsymbol{c}$ is the cell state vector.
+where $\boldsymbol{h_{n}}$ is the hidden state vector, $\boldsymbol{f}$ is the forget gate, $\boldsymbol{i}$ the input gate, and $\boldsymbol{o}$ the output gate, $\boldsymbol{c}$ is the cell state vector, and $\circ$ the element-wise product.
 The forget gate determines what information from previous steps should be kept or discarded.
@@
 The forget vector and input vector are obtained using the sigmoid function, denoted by $\gamma$, which determines how much information should be kept.
-Values of 00 in this vector indicate information to be discarded, while a value of $1$ indicates information to be kept.
-Furthermore, the $\phi$ represents the hyperbolic tangent function, used to keep the value of the cell state between $-1$ and $1$.
+Values of 0 in this vector indicate information to be discarded, while a value of $1$ indicates information to be kept.
+Furthermore, the $\phi$ represents the hyperbolic tangent function, used to keep the value of the cell state between 1-1 and $1$.
 Finally, $\boldsymbol{W}$ and $\boldsymbol{U}$ are the weights’ matrices, and $\boldsymbol{b}$ the vector bias terms.
-After the LSTM layer, a linear FC layer reduces the dimensionality down to $4$, ensuring consistency with all other models before proceeding to the conditioning block.
-In the case of ED, the input split into two equal-size parts, $\boldsymbol{x_{d}}$ containing $[x_{n},...,x_{n-31}]$ and $\boldsymbol{x_{e}}$ containing containing $[x_{n-32},...,x_{n-63}]$.
-The first part feds two separate convolutional layers, which compute the states, $[\boldsymbol{h},\boldsymbol{c}]$.
+ED layer
+In this case, the input of the network $\boldsymbol{x}$ is split into two equal-size parts, $\boldsymbol{x_{d}}$ containing $x_{n},...,x_{n-31}]$ and $\boldsymbol{x_{e}}$ containing containing $x_{n-32},...,x_{n-63}]$.
+The first part feeds two separate convolutional layers, which compute the states, $\boldsymbol{h},\boldsymbol{c}]$.
 These states are combined with the internal states of the LSTM layer computed during the previous iteration.
-A sigmoid function is applied to the LSTM internal states and then element-wise multiplied to $[h,c]$.
-This operation determines how much information should be taken from the computed $[h,c]$.
+A sigmoid function is applied to the LSTM internal states and then element-wise multiplied by $\boldsymbol{h},\boldsymbol{c}]$.
+This operation determines how much information should be taken from the computed $\boldsymbol{h},\boldsymbol{c}]$.
 The result of this operation represents the new LSTM’s internal states.
 This process is detailed in Figure 2.
-Figure 2: The states sharing process in ED: the encoder input $\boldsymbol{x_{e}}$ is passed to two convolutional layers, which produce the states $[\boldsymbol{h},\boldsymbol{c}]$.
-These states are combined with $[\boldsymbol{h_{n-1}},\boldsymbol{c_{n-1}}]$, which are the LSTM states of the previous iteration.
-The sigmoid function is applied to $[\boldsymbol{h_{n-1}},\boldsymbol{c_{n-1}}]$ and the results are element-wise multiplied to $[\boldsymbol{h},\boldsymbol{c}]$.
-The outcome will be the new states for the LSTM layer $[\boldsymbol{h_{n}},\boldsymbol{c_{n}}]$.
-The second part of the input vector is linearly transformed and fed into the second LSTM, following the same procedure as the other models.
-The third architecture is based on LRU.
+Figure 2: The states sharing process in ED: the encoder input $\boldsymbol{x_{e}}$ is passed to two convolutional layers, which produce the states $\boldsymbol{h},\boldsymbol{c}]$.
+These states are combined with $\boldsymbol{h_{n-1}},\boldsymbol{c_{n-1}}]$, which are the LSTM states of the previous iteration.
+The sigmoid function is applied to $\boldsymbol{h_{n-1}},\boldsymbol{c_{n-1}}]$ and the results are element-wise multiplied to $\boldsymbol{h},\boldsymbol{c}]$.
+The outcome will be the new states for the LSTM layer $\boldsymbol{h_{n}},\boldsymbol{c_{n}}]$.
+LRU layer
 Compared to a vanilla RNN, the LRU uses a linear activation function to compute the hidden state instead of a nonlinear function.
@@
 Additionally, to further enhance stability, the LRU incorporates a complex-valued diagonal recurrent matrix initialization and an exponential parameterization.
-At the same time, a normalization scheme for the hidden states is used to boost the efficiency and accuracy (Orvieto et al.,, 2023).
-Lastly, the SSMs operate by considering a continuous-time representation of the state-space formulation that is discretized to obtain the following equations:
+At the same time, a normalization scheme for the hidden states is used to boost the efficiency and accuracy Orvieto et al., (2023).
+S4D layer
+The SSMs operate by considering a continuous-time representation of the state-space formulation that is discretized to obtain the following equations:
 $\boldsymbol{h_{n}}$ $=\boldsymbol{\bar{W}_{A}}\boldsymbol{h_{n-1}}+\boldsymbol{\bar{W}_{B}}\boldsymbol{u_{n}}$
 $\boldsymbol{o_{n}}$ $=\boldsymbol{\bar{W}_{C}}\boldsymbol{h_{n}}+\boldsymbol{\bar{W}_{D}}\boldsymbol{u_{n}}.$ (5)
-In our study, we employ the diagonal state space model variant S4D  (Gu et al.,, 2022a), and newer S6 variant (Gu & Dao,, 2023).
-The remarkable ability of S4D to capture long-range dependencies stems from their utilization of a specific state matrix $\boldsymbol{\bar{W_{A}}}$ known as the "HiPPO matrix" (Gu et al.,, 2020).
-This matrix is designed to encode all the past input history in $\boldsymbol{h_{n}}$, finding a map from $\boldsymbol{u_{n}}$ to a higher dimensional space $\boldsymbol{h_{n}}$ that represents the compression of the history.
+In our study, we employ the diagonal state-space model variant S4D  Gu et al., (2022a).
+The remarkable ability of S4D to capture long-range dependencies stems from its utilization of a specific state matrix $\boldsymbol{\bar{W_{A}}}$ known as the "HiPPO matrix" Gu et al., (2020).
+This matrix is designed to encode the entire history of past inputs within $\boldsymbol{h_{n}}$, effectively mapping $\boldsymbol{u_{n}}$ to a higher-dimensional space $\boldsymbol{h_{n}}$ that encapsulates the compressed representation of this historical input data.
 The matrix allows the model to be conceptualized as a convolutional model that decomposes an input signal onto an orthogonal system of smooth basis functions.
-In this way, the state $\boldsymbol{h_{n}}$ encodes the history of the signal.
+In this way, the state $\boldsymbol{h_{n}}$ encodes the history of the inputs.
 In the diagonal form, the S4D model is defined by parameterizing its state matrix as a diagonal matrix, boosting the model’s efficiency.
-By doing so, the basis kernels have closed-form formulas represented by normalized Legendre polynomials $L_{n}(t)$, resulting in the SSM decomposing the input signal $\boldsymbol{u_{n}}$ onto an infinitely long set of basis functions that are orthogonal with respect to an exponentially decaying measure.
-In contrast, the S6 layer introduces input-dependent matrices.
+By doing so, the basis kernels have closed-form formulas represented by normalized Legendre polynomials $L_{n}(t)$, resulting in the SSM decomposing the input $\boldsymbol{u_{n}}$ onto an infinitely long set of basis functions that are orthogonal with respect to an exponentially decaying measure.
+S6 layer
+The S6 layer Gu & Dao, (2023) introduces input-dependent matrices.
 Specifically, $\boldsymbol{\bar{W_{B}}}$ and $\boldsymbol{\bar{W_{C}}}$ become dependent on the input and computed using a linear FC layer, while $\boldsymbol{\bar{W_{A}}}$ and $\boldsymbol{\bar{W_{D}}}$’s parameters are independent of the input and learned during the training process.
-The recurrent layer’s output size will equal the number of hidden units of the layer.
-In the case of LRU and S4D, a fully connected layer with a hyperbolic tangent activation function also follows the recurrent layer to compensate for a lack of nonlinear activation functions in these layers:
-$\boldsymbol{\hat{o}_{n}}=tanh(\boldsymbol{W_{nl}}\boldsymbol{o_{n}}+\boldsymbol{b_{nl}})$ (6)
+After the recurrent layer—whether LSTM, ED, LRU, S4D or S6—a FC layer reduces the dimensionality to $4$, ensuring uniformity across all model variations before the conditioning block.
+For the LRU, S4D, and S6, this FC layer includes a hyperbolic tangent activation function, compensating for the absence of nonlinear activation functions within these layers.
+Conversely, the LSTM model uses a linear activation function in the FC layer, as detailed in the following equations:
+$\boldsymbol{\hat{o}_{n}}$ $=tanh(\boldsymbol{W_{nl}}\boldsymbol{o_{n}}+\boldsymbol{b_{nl}})\text{ if }\textbf{LRU / S4D / S6}$
+$\boldsymbol{\hat{o}_{n}}$ $=(\boldsymbol{W_{nl}}\boldsymbol{o_{n}}+\boldsymbol{b_{nl}})\text{ if }\textbf{LSTM}$ (6)
 The conditioning block is then applied just before the output layer, which consists of a linear fully connected layer with one unit, which is described in the following subsection.
-Placing the conditioning block after the recurrent layers rather than before has been found beneficial  (Fasciani et al.,, 2024a).
+Placing the conditioning block after the recurrent layers rather than before has been found beneficial  Fasciani et al., (2024a).
 This suggests that it is more beneficial for the networks to use the information given by the control parameters to project the output of the recurrent layer and determine the extent to which this information influences the final output rather than influencing the inference of the recurrent layer based on the control parameters.
@@
 Since the output layer has one unit, $\boldsymbol{W_{out}}$ is a vector of the same dimensionality as $\boldsymbol{o^{c}_{n}}$.
-The number of units is chosen to have approximately $800$ trainable parameters among the architectures and effects since an identical number is achievable due to major differences in the layers.
+The number of units is chosen to have approximately $800$ trainable parameters among the architectures and effects, since an identical number is achievable due to major differences in the layers.
 As a result, the selected LSTM and ED present $8$ units in the recurrent layer and $4$ in fully connected ones.
-LRU and S4D present $12$ units in the recurrent layer, and $6$ in fully connected ones.
-The number of trainable parameters also depends on the dimensional of the conditioning input $\boldsymbol{p}$, equal to the number of variable parameters in the datasets.
+LRU, S4D and S6 present $12$ units in the recurrent layer, and $6$ in fully connected ones.
+The number of trainable parameters also depends on the dimensionality of the conditioning input $\boldsymbol{p}$, equal to the number of variable parameters in the datasets.
 3.1.1 Conditioning Block
-The conditioning block consists of the FiLM method and the GLU layer, as in (Fasciani et al.,, 2024a), and is illustrated in Figure 3.
+The conditioning block consists of the FiLM method and the GLU layer, as in Fasciani et al., (2024a), and is illustrated in Figure 3.
 Figure 3: Conditioning block based on the FiLM method and the GLU layer.
@@
 FiLM is a technique that applies an affine transformation to a vector based on conditioning values.
-The parameters vector $\boldsymbol{p}$, if applicable, is the input to a linear fully connected layer.
+The parameters vector $\boldsymbol{p}$, if applicable, is the input to a linear FC layer.
 The output of this layer is then split into two vectors of equal size: $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$.
-If $\boldsymbol{o_{n}}$ is the vector to which the transformation is to be applied, the output of the linear fully connected layer will have a length double of $\boldsymbol{o_{n}}$.
-Consequently, FiLM performs am affine transform on $\boldsymbol{o_{n}}$ using $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$.
+If $\boldsymbol{o_{n}}$ is the vector to which the transformation is applied, the output of the linear FC layer is twice the length of $\boldsymbol{o_{n}}$.
+Consequently, FiLM performs an affine transform on $\boldsymbol{o_{n}}$ using $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$:
+$\boldsymbol{q}$ $=\boldsymbol{\theta}\boldsymbol{o_{n}}+\boldsymbol{\eta}.$ (8)
 Following the projection, it is common to apply a function, such as ReLU or sigmoid, to determine the amount of information that should be passed.
@@
 This layer is more stable than ReLU and exhibits faster learning compared to sigmoid.
-Similarly to the previous step, the GLU layer consists of a linear fully connected layer that takes the FiLM output vector as input and computes a vector with twice its length.
-The resulting output is then split equally into two vectors: $\boldsymbol{q_{1}}$ and $\boldsymbol{q_{2}}$.
-A function is applied to $\boldsymbol{q_{2}}$, and the resulting output is multiplied element-wise with $\boldsymbol{q_{1}}$.
-Hence, the conditioning block is described by the following:
-$\boldsymbol{o^{c}_{n}}=\boldsymbol{q_{1}}\odot softsign(\boldsymbol{q_{2}}),\text{ with }\boldsymbol{q_{1}}$ $=\boldsymbol{\dot{o}_{n}}\boldsymbol{K_{1}}+\boldsymbol{d_{1}}$
-$\boldsymbol{q_{2}}$ $=\boldsymbol{\dot{o}_{n}}\boldsymbol{K_{2}}+\boldsymbol{d_{2}}$
-$\boldsymbol{\dot{o}_{n}}$ $=(\boldsymbol{\theta}\boldsymbol{o_{n}}+\boldsymbol{\eta}).$ (8)
+Similarly to the previous step, the GLU layer consists of a linear FC layer that takes the FiLM output vector as input and computes a vector twice its length.
+The resulting output is then divided equally into two vectors: $\boldsymbol{q_{1}}$ and $\boldsymbol{q_{2}}$.
+A function is applied to $\boldsymbol{q_{2}}$, with the resulting output being multiplied element-wise with $\boldsymbol{q_{1}}$.
+This operation describes the conditioning block, represented by the following equation:
+$\boldsymbol{o^{c}_{n}}=\boldsymbol{q_{1}}\odot softsign(\boldsymbol{q_{2}})$ (9)
 The GLU layer determines the flow of information through the network, acting as a logical gate.
@@
 The softsign function controls the extent to which the control parameters should positively or negatively influence the final output.
-A value of 00 indicates a bypass, meaning no influence on the output.
+A value of 0 indicates a bypass, meaning no influence on the output.
 This allows for more flexibility.
@@
 This allows for a fair comparison between the trained models, as they all attempt to minimize the same loss function.
-Therefore, we selected the MSE for this comparative study as it helps to capture the dynamic aspects of the output signal more effectively  (Simionato & Fasciani,, 2023).
+Therefore, we selected the MSE for this comparative study as it helps to capture the dynamic aspects of the output signal more effectively  Simionato & Fasciani, (2023).
 In addition, our quantitative evaluation is based on various metrics to assess the models from different perspectives.
-In the following expressions, $y_{n}$ and $\hat{y}_{n}$ represent the target and predicted signal and the symbol $N$ represents the overall number of samples in the output signal.
-The first metric is the root-mean-square energy (RMSE) error.
-RMSE error provides insights into the energy deviation between the target and prediction.
+In the following expressions, $\boldsymbol{y}$ and $\hat{\boldsymbol{y}}$ represent the target and predicted audio signals, and the symbol $N$ represents the overall number of samples in the output signal.
+The first metric we utilize is the normalized root-mean-square energy (NRMSE) error, with normalization carried out using the RMSE of the target signal.
+NRMSE error provides insights into the energy deviation between the target and the prediction.
 The losses and metrics presented so far calculate the average difference between the target and prediction, which may hide large errors that occur only during signal transients.
@@
 In this case, we may obtain a relatively low error even with all transients poorly modeled.
-Therefore, we have included a metric related to the spectral flux, which considers the differences between pairs of spectrums computed on consecutive overlapping windows, as used in the library  [^1] https://github.com/magenta/ddsp developed after  (Engel et al.,, 2020):
-$\begin{split}M_{SF}=\Big{|}\Big{|}\Big{(}\Big{|}|STFT(y_{n})|-|STFT(y_{n-1})|\Big{|}\Big{)}\\ -\Big{(}\Big{|}|STFT(\hat{y}_{n})|-|STFT(\hat{y}_{n-1})|\Big{|}\Big{)}\Big{|}\Big{|}_{1}\end{split}$ (9)
+Therefore, we have included a metric related to the spectral flux, which considers the differences between pairs of spectra computed on consecutive overlapping windows, as used in the library  [^1] https://github.com/magenta/ddsp developed after  Engel et al., (2020):
+[…]
 where the Short-Time Fourier Transform (STFT) is computed with a window size equal to $2048$ and hop size equal to $512$ samples.
-This metric emphasizes the ability to predict portions of the output sound with rapid changes.
-To account for the frequency response, we use the multi-resolution STFT  (Engel et al.,, 2020):
-$\begin{split}M_{STFT}&=\frac{1}{N}\sum_{m}|||STFT_{m}(y_{n})|-|STFT_{m}(\hat{y}_{n})|||_{1}\\ &+\frac{1}{N}\sum_{m}||\log(|STFT_{m}(y_{n})|)-\log(|STFT_{m}(\hat{y}_{n})|)||_{1}\end{split}$ (10)
+The metric is normalized by the squared difference between the magnitudes of the spectra of two successive target frames, measuring the spectral change between them.
+This indicates the rate at which the power spectrum of a signal is evolving over time.
+To account for the frequency response, we use the multi-resolution STFT  Engel et al., (2020):
+[…]
 which compares linear and logarithmic spectral distances with varying frequency resolutions using the L1 norm.
+The linear component of $M_{STFT}$ is normalized by the magnitude of the STFT of the target.
 The error is calculated as the sum of the absolute values of the vectors.
@@
 We compute the metrics using $m=[256,512,1024]$ as resolutions.
-Similarly, we compare the MFCCs to compute a metric linked to human sound perception:
-$M_{MFCC}=\frac{1}{N}|||MFCC(y_{n})|-|MFCC(\hat{y}_{n})|||_{1}$ (11)
-where $1024$ is the FFT length, and we have $80$ Mel bands between $30$ Hz and Nyquist frequency.
 3.3 Audio Effects and Datasets
@@
 Table 2 summarizes all the datasets included in the study.
-Delay-based effects require a different approach than a direct black box  (Wright & Välimäki,, 2021),  (Mikkonen et al.,, 2023).
+Delay-based effects require a different approach than a direct black box  Wright & Välimäki, (2021),  Mikkonen et al., (2023).
 For this reason, they are not included in this study.
 The dataset we use is derived from the following hardware devices: Behringer OD300 overdrive pedal  [^2] https://www.behringer.com/product.html?modelCode=P0608, Behringer Neutron  [^3] https://www.behringer.com/product.html?modelCode=P0CM5’s overdrive module, Behringer Neutron’s filter module in low-pass mode, TC Electronic Bucket Brigade Analog Delay used as a saturator  [^4] https://www.tcelectronic.com/product.html?modelCode=P0EBV, CL 1B TubeTech  [^5] http://www.tube-tech.com/cl-1b-opto-compressor/and Teletronix LA-2A  [^6] https://www.uaudio.com/hardware/la-2a.html optical compressors.
-We have also collected data from two software VA plugins using  (Fasciani et al.,, 2024b): the Helper Saturator  [^7] https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ  [^8] https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
+We have also collected data from two software VA plugins using  Fasciani et al., (2024b): the Helper Saturator  [^7] https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ  [^8] https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
 Overdrive
@@
 We isolated the module and recorded the dataset at $10$ equally spaced values for the cutoff and $5$ for the resonance, selecting the low-pass filter type.
-The filter was disconnected from the ADSR envelope module.
+The filter was separated from the envelope module, which can be used to modulate it.
 Saturator
-Helper Saturator software emulates analog saturation, in particular tube and tape saturation.
+The Helper Saturator software emulates analog saturation, in particular tube and tape saturation.
 The software presents four parameters: lowpass and highpass cutoffs, saturation type switch (tape or tube), and saturation level.
@@
 The equalizer boosts and cuts the selected frequency bands, giving some saturation simultaneously.
-In this case, we recorded the dataset varying the boost and attenuator knobs at $5$ different equally-spaced values up to half of their range, setting $60$ Hz as the low-frequency switch value and 00 (sharp) as the bandwidth.
+In this case, we recorded the dataset varying the boost and attenuator knobs at $5$ different equally-spaced values up to half of their range, setting $60$ Hz as the low-frequency switch value and 0 (sharp) as the bandwidth.
 Compressor/Limiter
@@
 We consider three equally spaced values for the ratio, attack, and release time and four values for the threshold.
-On the other hand, LA-2A presents the peak reduction knob indicating the amount of compression ([00,$100$]) and a mode switch (limiter, compression).
+On the other hand, LA-2A presents the peak reduction knob indicating the amount of compression ([0,$100$]) and a mode switch (limiter, compression).
 We consider $10$ equally spaced values for the peak reduction.
@@
 Frequency Attenuator [0,10]
-LA-2A [^1] Recorded from hardware devices.  [^3] Publicly available. Optical Peak Reduction [0,100] $20$
+LA-2A [^1] Recorded from hardware devices. Optical Peak Reduction [0,100] 2020
 Compressor Switch Mode [Compressor,Limiter]
@@
 3.4 Experimenting and Learning
-The models are trained for $200$ epochs and using the Adam (Kingma & Ba,, 2014) optimizer with a gradient norm scaling of $1$  (Pascanu et al.,, 2013).
+The models are trained for $200$ epochs and use the Adam Kingma & Ba, (2014) optimizer with a gradient norm clipped at 11  Pascanu et al., (2013).
 The training was stopped earlier in case of no reduction of validation loss for $10$ epochs.
-We design a learning rate with an exponential decay according to the formula: $lr=LR*0.25^{e}$, where LR is the initial learning rate and it is set to $3\cdot 10^{-4}$, and $e$ is the number of the epochs.
+We design a learning rate with an exponential decay according to the formula: $lr=LR*0.25^{e}$, where LR is the initial learning rate and it is set to $3\cdot 10^{-4}$, and $e$ is the number of epochs.
 Test losses and evaluation metrics are computed using the model’s weights that minimize the validation loss throughout the training epochs.
 Finally, the input signal is split into segments of $2400$ samples (equivalent to $50$ ms) to be processed before updating the weights.
-Considering the network architecture described in Section 3.1, in which a buffer of the $64$ most recent samples is used as input, training layers with states necessitates sequential input-output examples.
-Consequently, our implementation adopts an input shape of ($B$, $L$, $F$), which is set to ($B$, $1$, $64$), where $B$ denotes the arbitrary batch size, $L$ the sequence length, and $F$ the input dimensionality, often identified as the number of input features.
-This configuration considers the $64$ input samples as the vector of input features at the current sampling instant.
+Considering the network architecture described in Section 3.1.
+Consequently, our implementation adopts an input shape of ($B$, $L$, $F$), which is set to ($B$, $2400$, $64$), where $B$ denotes the arbitrary batch size, $L$ the sequence length, and $F$ the input dimensionality, often identified as the number of input features.
+This configuration is used within the framework of Truncated Backpropagation Through Time (TBPTT).
 We adopted the same learning schedule and minibatch size for all models across all datasets.
-As explained in Section 3.3, we collected the datasets using the same approach and the same input file to minimize the difference among them.
+As explained in Section 3.3, we collected the datasets using the same approach and the same input file to minimize the differences among them.
 However, it is important to highlight that each dataset and modeling task may benefit from specific adjustments to achieve more accurate results.
 Different datasets and tasks may require slightly different learning rate schedules; an excessively slow or fast learning rate can lead to suboptimal solutions.
-The number of samples to process before updating the weights referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
-The composition of the training, validation, and test sets is identical across datasets.
-We used $80\%$ for the training set, $10\%$ for the validation set, and the remaining $10\%$ for the test set.
-For effects with variable control parameters, each $45$s recording in the dataset is associated with a specific combination of the effect parameters.
-The $80-10-10\%$ split was carried out at the individual recording level, ensuring that all control parameter combinations present an equal share in each subset and are equally used for training and evaluation of the model.
-The initial $10\%$ of all the recordings are used as the test set.
-The following $10\%$ portion of the recordings is used as a validation set, while the remaining parts were used as the training set.
-Minor manual adjustments are made to ensure that splitting points fall within segments of silence.
+The number of samples to process before updating the weights, referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
+To ensure robust evaluation of the models, we divided the collected data into $5$ distinct compositions of training, validation, and test sets.
+This strategy aimed to cross-validate the models across different data splits.
+In each composition, $80\%$ of the data is allocated for training, $10\%$ is reserved for validation, and the remaining $10\%$ constitutes the test set.
+Importantly, the validation and test sets in each composition include different data signals, ensuring that the model never uses the same audio signals in both validation and testing processes.
+This approach prevents any bias towards certain data signals and allows for a more accurate assessment of the model’s performance across varying conditions and inputs.
+For audio effects with variable control parameters, each $45$ s recording in the dataset corresponds to a specific combination of effect parameters.
+To maintain consistency across training, validation, and testing, we employ a $80-10-10\%$ split at the level of individual recordings.
+This ensures that all combinations of control parameters are proportionally represented in each subset, providing balanced usage for both training the model and evaluating its performance.
+Additionally, minor manual adjustments were made to the splitting points to ensure they occur during silent segments, preventing abrupt changes in audio dynamics that could affect model accuracy.
+The statistical significance of the results is analyzed using the Friedman test and the Wilcoxon signed-rank test  demvsar2006statistical; rainio2024evaluation.
+Both are non-parametric statistical tests employed to determine whether there are statistically significant differences between the medians of samples.
+These tests are well-suited for data that does not meet the assumptions required for parametric tests, such as normality.
+In this case, where we are evaluating $5$ models across $5$ different dataset compositions, the Friedman test is the most appropriate for assessing statistical significance.
+However, due to the limited number of dataset compositions, the reliability of the Friedman test might be reduced.
+To address this, we also conduct separate Wilcoxon signed-rank tests for each model pair.
+While this test is useful for assessing significant differences between two models, it’s important to note that it is not ideal for comparing multiple models.
+Specifically, conducting multiple Wilcoxon tests increases the risk of incorrectly finding significant differences due to the numerous comparisons involved.
 4 Results
-The evaluation is based on the metrics described in Section 3.2 and presented in Table 4.
-The MSE loss and the epoch-related epoch resulting in lower validation loss during training are reported in Table 3.
+The MSE loss, the metrics described in Section 3.2, and the mean of the epochs resulting in lower validation loss during training are reported in Table 3.
 The losses and the metrics consider all conditioning scenarios and provide an average error across all combinations of parameters.
-To further verify the model’s capacity to learn the conditioned effect, we provide a series of plots that compare the target signal with the predictions generated by the four distinct architectures.
+The average epoch at which the training met the early stopping condition is also reported in the table.
+LRU is generally the slowest to converge, taking significantly more epochs to reach the same range of errors as the other models.
+On the other hand, other variants are dataset-specific, with a tendency for the SSM-based model, such as S4D and S6, to converge faster and with a similar number of epochs.
+Similarly, both LSTM and ED require a comparable number of epochs, although they generally converge more slowly than models based on SSM.
+This slower convergence is attributed to their similar layer formulation.
+On the other hand, LRU utilizes specific initializations and normalization schemes, distinguishing it from traditional recurrent and SSM architectures.
+Table 3: Mean values of MSE loss, ESR, $M_{NRMSE}$, $M_{SF}$, and $M_{STFT}$ metrics computed across different test sets, with each set linked to a model trained on a distinct dataset split.
+The last column reports the average epoch at which the training met the early stopping condition.
+Dataset Model MSE ESR $M_{NRMSE}$ $M_{SF}$ $M_{STFT}$ Epoch
+OD300 LSTM $\boldsymbol{2.92\cdot 10^{-4}}$ $\boldsymbol{1.13\cdot 10^{-1}}$ $\boldsymbol{2.69\cdot 10^{-1}}$ $\boldsymbol{4.56\cdot 10^{-1}}$ $7.50\cdot 10^{-1}$ $157$
+Overdrive ED $3.59\cdot 10^{-4}$ $2.00\cdot 10^{-1}$ $2.90\cdot 10^{-1}$ $5.66\cdot 10^{-1}$ $7.78\cdot 10^{-1}$ $158$
+LRU $4.57\cdot 10^{-4}$ $1.77\cdot 10^{-1}$ $3.44\cdot 10^{-1}$ $5.35\cdot 10^{-1}$ $8.12\cdot 10^{-1}$ $170$
+S4D $3.14\cdot 10^{-4}$ $1.22\cdot 10^{-1}$ $2.88\cdot 10^{-1}$ $4.89\cdot 10^{-1}$ $\boldsymbol{7.40\cdot 10^{-1}}$ $189$
+S6 $4.45\cdot 10^{-4}$ $1.72\cdot 10^{-1}$ $3.57\cdot 10^{-1}$ $5.43\cdot 10^{-1}$ $8.82\cdot 10^{-1}$ $153$
+Neutron LSTM $\boldsymbol{5.64\cdot 10^{-4}}$ $\boldsymbol{6.21\cdot 10^{-2}}$ $\boldsymbol{2.51\cdot 10^{-1}}$ $\boldsymbol{3.25\cdot 10^{-1}}$ $5.45\cdot 10^{-1}$ $18$
+Overdrive ED $7.84\cdot 10^{-4}$ $8.64\cdot 10^{-2}$ $2.87\cdot 10^{-1}$ $3.53\cdot 10^{-1}$ $\boldsymbol{5.42\cdot 10^{-1}}$ $54$
+LRU $9.65\cdot 10^{-4}$ $1.10\cdot 10^{-1}$ $3.10\cdot 10^{-1}$ $4.09\cdot 10^{-1}$ $5.98\cdot 10^{-1}$ $106$
+S4D $1.12\cdot 10^{-3}$ $1.30\cdot 10^{-1}$ $4.20\cdot 10^{-1}$ $5.33\cdot 10^{-1}$ $5.98\cdot 10^{-1}$ $10$
+S6 $1.16\cdot 10^{-3}$ $1.32\cdot 10^{-1}$ $4.42\cdot 10^{-1}$ $5.08\cdot 10^{-1}$ $6.58\cdot 10^{-1}$ $10$
+Helper LSTM $3.47\cdot 10^{-4}$ $1.91\cdot 10^{-1}$ $3.81\cdot 10^{-1}$ $\boldsymbol{4.65\cdot 10^{-1}}$ $6.88\cdot 10^{-1}$ $180$
+Saturator ED $3.49\cdot 10^{-4}$ $1.66\cdot 10^{-1}$ $3.82\cdot 10^{-1}$ $4.82\cdot 10^{-1}$ $6.70\cdot 10^{-1}$ $94$
+LRU $6.11\cdot 10^{-4}$ $3.36\cdot 10^{-1}$ $4.96\cdot 10^{-1}$ $5.54\cdot 10^{-1}$ $8.90\cdot 10^{-1}$ $157$
+S4D $\boldsymbol{2.93\cdot 10^{-4}}$ $\boldsymbol{1.62\cdot 10^{-1}}$ $\boldsymbol{3.54\cdot 10^{-3}}$ $4.72\cdot 10^{-1}$ $6.59\cdot 10^{-1}$ $109$
+S6 $2.99\cdot 10^{-4}$ $1.63\cdot 10^{-1}$ $3.57\cdot 10^{-1}$ $4.85\cdot 10^{-1}$ $\boldsymbol{6.57\cdot 10^{-1}}$ $176$
+Pultec LSTM $\boldsymbol{4.26\cdot 10^{-5}}$ $\boldsymbol{1.14\cdot 10^{-2}}$ $\boldsymbol{9.63\cdot 10^{-3}}$ $\boldsymbol{7.11\cdot 10^{-2}}$ $\boldsymbol{1.85\cdot 10^{-1}}$ $66$
+Passive ED $1.43\cdot 10^{-4}$ $3.49\cdot 10^{-2}$ $56\cdot 10^{-1} 1.1$$18\cdot 10^{-1}$ $2.92\cdot 10^{-1}$ $130$
+Equalizer LRU $2.19\cdot 10^{-4}$ $5.60\cdot 10^{-2}$ $1.98\cdot 10^{-1}$ $2.17\cdot 10^{-1}$ $3.62\cdot 10^{-1}$ $173$
+S4D $1.62\cdot 10^{-4}$ $6.97\cdot 10^{-2}$ $8.73\cdot 10^{-2}$ $7.46\cdot 10^{-2}$ $2.28\cdot 10^{-1}$ 88
+S6 $4.79\cdot 10^{-5}$ $1.20\cdot 10^{-2}$ $1.00\cdot 10^{-1}$ $8.41\cdot 10^{-1}$ $1.96\cdot 10^{-1}$ $56$
+LA-2A LSTM $90\cdot 10^{-4} 1.1$$11\cdot 10^{-1}$ $11\cdot 10^{-1} 3.3$$34\cdot 10^{-1}$ $6.61\cdot 10^{-1}$ $57$
+Optical ED $4.22\cdot 10^{-5}$ $2.36\cdot 10^{-2}$ $1.60\cdot 10^{-1}$ $1.78\cdot 10^{-1}$ $3.47\cdot 10^{-1}$ $70$
+Comp LRU $5.08\cdot 10^{-5}$ $2.82\cdot 10^{-2}$ $1.67\cdot 10^{-1}$ $2.22\cdot 10^{-1}$ $3.97\cdot 10^{-1}$ $154$
+S4D $3.76\cdot 10^{-5}$ $2.07\cdot 10^{-2}$ $1.47\cdot 10^{-1}$ $1.62\cdot 10^{-1}$ $3.44\cdot 10^{-1}$ $54$
+S6 $\boldsymbol{3.60\cdot 10^{-5}}$ $\boldsymbol{1.96\cdot 10^{-2}}$ $\boldsymbol{1.43\cdot 10^{-1}}$ $\boldsymbol{1.53\cdot 10^{-1}}$ $\boldsymbol{3.31\cdot 10^{-1}}$ $22$
+CL 1B LSTM $3.37\cdot 10^{-4}$ $2.72\cdot 10^{-2}$ $4.88\cdot 10^{-1}$ $5.97\cdot 10^{-1}$ $9.16\cdot 10^{-1}$ $35$
+Optical ED $\boldsymbol{2.68\cdot 10^{-5}}$ $\boldsymbol{2.01\cdot 10^{-2}}$ $\boldsymbol{1.20\cdot 10^{-1}}$ $\boldsymbol{1.43\cdot 10^{-1}}$ $\boldsymbol{3.51\cdot 10^{-1}}$ $52$
+Comp LRU $6.21\cdot 10^{-5}$ $5.02\cdot 10^{-2}$ $1.96\cdot 10^{-1}$ $2.23\cdot 10^{-1}$ $4.59\cdot 10^{-1}$ $168$
+S4D $1.09\cdot 10^{-4}$ $9.06\cdot 10^{-2}$ $2.78\cdot 10^{-1}$ $2.95\cdot 10^{-1}$ $5.83\cdot 10^{-1}$ $28$
+S6 $6.64\cdot 10^{-5}$ $5.12\cdot 10^{-2}$ $1.83\cdot 10^{-1}$ $2.00\cdot 10^{-1}$ $4.66\cdot 10^{-1}$ $14$
+Neutron LSTM $3.18\cdot 10^{-3}$ $7.03\cdot 10^{-1}$ $8.28\cdot 10^{-1}$ $8.62\cdot 10^{-1}$ $1.19$ $76$
+Filter ED $2.91\cdot 10^{-3}$ $6.38\cdot 10^{-1}$ $\boldsymbol{6.82\cdot 10^{-1}}$ $\boldsymbol{7.57\cdot 10^{-1}}$ $1.13$ $54$
+LRU $2.62\cdot 10^{-3}$ $5.77\cdot 10^{-1}$ $7.18\cdot 10^{-1}$ $\cdot 10^{-1} 1.081.08$ $179$
+S4D $\boldsymbol{2.52\cdot 10^{-3}}$ $\boldsymbol{5.56\cdot 10^{-1}}$ $6.94\cdot 10^{-1}$ $7.90\cdot 10^{-1}$ $1.06$ $16$
+S6 $2.62\cdot 10^{-3}$ $5.77\cdot 10^{-1}$ $7.21\cdot 10^{-1}$ $8.11\cdot 10^{-1}$ $\boldsymbol{1.04}$ 77
+Figure 4: Comparison of the RMS energy between the target and predicted output.
+The RMS is computed using windows of $4096$ samples with $75\%$ overlap.
+When the model includes conditioning, the results refer to the highest conditioning values, representing the most challenging scenario.
+In the OD300 and Neutron’s overdrive dataset, the LSTM model demonstrates superior overall performance.
+However, it does not achieve the lowest $M_{STFT}$, although its values are close to the best.
+Additionally, S6 struggles more with emulating both effects, particularly with the Neutron’s overdrive dataset.
+When trained on the Saturator dataset, S4D, S6, and LRU deliver the strongest performance, with LSTM and ED slightly trailing.
+In contrast, for the Equalizer dataset, LSTM consistently performs better than the other models.
+The optical compressor features extended transients and lengthy time dependencies between the input signal and the alteration applied by the device, influenced by its adjustable attack and release parameters.
+This complexity is especially pronounced with the CL 1B compressor, presenting a challenging task for ANN modeling.
+Of the tested models, LSTM is the least accurate, with higher losses in both the LA-2A and CL 1B scenarios.
+In contrast, ED, S4D, and S6 demonstrate superior accuracy for these tasks, with S6 excelling in the LA-2A case and ED in the CL 1B case.
+In analyzing the metrics of models trained to emulate the low-pass filter, both S4D and ED show the best performance.
+However, all models reveal a considerable frequency mismatch with the target output.
+This suggests their inability to model the signal alteration process of this parametric effect, which selectively removes high-frequency components while leaving those below the cutoff frequency unchanged.
+Overdrive effects significantly distort the input signal, leading to the introduction of harmonic components.
+In these cases, LSTM-based models have been shown to be more effective.
+Conversely, the saturator determines mild distortion, similar to tape processing, by smoothing out the transients of the input signal.
+This situation may involve hysteresis and short delays associated with magnetic recording  Mikkonen et al., (2023), leading to longer temporal dependencies between input and output compared to distortion effects.
+This characteristic may explain the advantages observed with SSM-based layers.
+However, we lack information about the specific implementation of the software used, preventing us from confirming this hypothesis.
+Conversely, The Equalizer introduces tube-like saturation, which may explain the better performance of LSTMs in this context.
+SSM-based layers also demonstrated improved performance with optical compressors, where the effective management of transients and temporal dependencies is crucial.
+When ranging the models according to performance metrics, inconsistencies arise with $M_{SF}$ and $M_{STFT}$, as these metrics don’t always align with the trends observed in time-based metrics.
+Nonetheless, no distinct patterns are apparent.
+This could be due to the loss function’s emphasis on the time domain, where variations in values are typically subtle.
+To further verify the model’s capacity to learn the conditioned effect, we provide a series of plots that compare the target signal with the predictions generated by the five distinct architectures.
 For these plots, we chose the modeling scenario that we believe to be the most challenging based on the effect parameters used as conditioning factors for the networks.
@@
 Regarding the low-pass filter, the parameters are configured for minimum cutoff and maximum resonance.
-The plots show the RMSE in Figure  4, and the spectrograms in Figure  5.
-For datasets including conditioning parameters, the plots showing predictions versus the target are generated for a $2.5$ seconds input signal containing bass, guitar, and drum loop sound.
+The plots show the RMS energy in Figure  4, and the spectrograms in Figure  5.
+For datasets including conditioning parameters, the plots showing predictions versus the target are generated for a $10$ second input signal containing bass, guitar, and drum loop sound.
 These input signals are part of the respective test sets.
-To ensure a clearer comparative visualization of the spectral flux, we use only the first $1.25$ seconds.
-Table 3: MSE test loss computed on the test set. The test sets include all conditioning values, if applicable. The last column reports the epoch at which the training met the early stopping condition.
-Dataset Model MSE Epoch Dataset Model MSE Epoch
-OD300 LSTM $\boldsymbol{2.46\cdot 10^{-4}}$ $144$ Neutron LSTM $\boldsymbol{2.27\cdot 10^{-3}}$ $4$
-Overdrive ED $3.57\cdot 10^{-4}$ $176$ Overdrive ED $2.79\cdot 10^{-3}$ $3$
-LRU $2.98\cdot 10^{-4}$ $170$ LRU $2.43\cdot 10^{-3}$ $100$
-S4D $3.23\cdot 10^{-4}$ $165$ S4D $2.62\cdot 10^{-3}$ $8$
-S6 $3.39\cdot 10^{-4}$ $172$ S6 $2.41\cdot 10^{-3}$ $28$
-Helper LSTM $4.56\cdot 10^{-4}$ $165$ Pultec LSTM $\boldsymbol{7.41\cdot 10^{-4}}$ $19$
-Saturator ED $4.41\cdot 10^{-4}$ $138$ Passive ED $1.11\cdot 10^{-3}$ $101$
-LRU $\boldsymbol{4.14\cdot 10^{-4}}$ $176$ Equalizer LRU $1.03\cdot 10^{-3}$ $150$
-S4D $4.40\cdot 10^{-4}$ $149$ S4D $1.06\cdot 10^{-3}$ $11$
-S6 $4.53\cdot 10^{-4}$ $188$ S6 $9.68\cdot 10^{-4}$ $8$
-LA-2A LSTM $1.11\cdot 10^{-4}$ $59$ CL 1B LSTM $1.78\cdot 10^{-4}$ $74$
-Optical ED $9.54\cdot 10^{-5}$ $90$ Optical ED $\boldsymbol{8.96\cdot 10^{-5}}$ $38$
-Compressor LRU $1.14\cdot 10^{-4}$ $146$ Compressor LRU $6.76\cdot 10^{-4}$ $168$
-S4D $8.43\cdot 10^{-5}$ $62$ S4D $1.22\cdot 10^{-4}$ $14$
-S6 $\boldsymbol{8.25\cdot 10^{-5}}$ $42$ S6 $1.01\cdot 10^{-4}$ $19$
-Neutron LSTM $4.07\cdot 10^{-3}$ $46$
-Filter ED $3.36\cdot 10^{-3}$ $9$
-LRU $\boldsymbol{2.43\cdot 10^{-3}}$ $188$
-S4D $2.75\cdot 10^{-3}$ $12$
-S6 $3.16\cdot 10^{-3}$ $7$
-Table 4: ESR, $M_{RMSE}$, $M_{SF}$, and $M_{STFT}$ metrics computed on the test set. The test sets include all conditioning values, if applicable.
-Dataset Model ESR $M_{RMSE}$ $M_{SF}$ $M_{STFT}$ $M_{MFCC}$
-OD300 LSTM $\boldsymbol{1.06\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $1.60\cdot 10^{-2}$ $3.56\cdot 10^{-1}$ $6.08\cdot 10^{-1}$
-Overdrive ED $1.54\cdot 10^{-1}$ $1.30\cdot 10^{-2}$ $1.19\cdot 10^{-2}$ $3.52\cdot 10^{-1}$ $5.80\cdot 10^{-1}$
-LRU $2.35\cdot 10^{-1}$ $1.65\cdot 10^{-2}$ $1.77\cdot 10^{-2}$ $4.55\cdot 10^{-1}$ $1.05$
-S4D $1.40\cdot 10^{-1}$ $\boldsymbol{1.20\cdot 10^{-2}}$ $\boldsymbol{8.50\cdot 10^{-3}}$ $3.11\cdot 10^{-1}$ $\boldsymbol{5.38\cdot 10^{-1}}$
-S6 $1.46\cdot 10^{-1}$ $1.25\cdot 10^{-2}$ $1.04\cdot 10^{-2}$ $\boldsymbol{2.78\cdot 10^{-1}}$ $5.61\cdot 10^{-1}$
-Neutron LSTM $\boldsymbol{2.50\cdot 10^{-1}}$ $\boldsymbol{1.86\cdot 10^{-2}}$ $5.74\cdot 10^{-3}$ $4.80\cdot 10^{-1}$ $3.49\cdot 10^{-1}$
-Overdrive ED $3.07\cdot 10^{-1}$ $2.36\cdot 10^{-2}$ $6.33\cdot 10^{-3}$ $5.64\cdot 10^{-1}$ $3.80\cdot 10^{-1}$
-LRU $2.70\cdot 10^{-1}$ $1.94\cdot 10^{-2}$ $6.64\cdot 10^{-3}$ $\boldsymbol{3.40\cdot 10^{-1}}$ $3.39\cdot 10^{-1}$
-S4D $2.88\cdot 10^{-1}$ $2.21\cdot 10^{-2}$ $5.71\cdot 10^{-3}$ $5.20\cdot 10^{-1}$ $3.59\cdot 10^{-1}$
-S6 $2.65\cdot 10^{-1}$ $1.89\cdot 10^{-2}$ $\boldsymbol{4.97\cdot 10^{-3}}$ $3.86\cdot 10^{-1}$ $\boldsymbol{3.05\cdot 10^{-1}}$
-Helper LSTM $2.67\cdot 10^{-1}$ $1.05\cdot 10^{-2}$ $2.78\cdot 10^{-2}$ $5.97\cdot 10^{-1}$ $4.72\cdot 10^{-1}$
-Saturator ED $2.58\cdot 10^{-1}$ $1.06\cdot 10^{-2}$ $1.29\cdot 10^{-2}$ $5.33\cdot 10^{-1}$ $4.14\cdot 10^{-1}$
-LRU $3.46\cdot 10^{-1}$ $\boldsymbol{9.72\cdot 10^{-3}}$ $1.42\cdot 10^{-2}$ $5.44\cdot 10^{-1}$ $\boldsymbol{2.46\cdot 10^{-1}}$
-S4D $\boldsymbol{2.57\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $\boldsymbol{1.22\cdot 10^{-2}}$ $5.41\cdot 10^{-1}$ $3.67\cdot 10^{-1}$
-S6 $2.65\cdot 10^{-1}$ $1.01\cdot 10^{-2}$ $1.31\cdot 10^{-2}$ $\boldsymbol{4.28\cdot 10^{-1}}$ $3.88\cdot 10^{-1}$
-Pultec LSTM $\boldsymbol{2.15\cdot 10^{-1}}$ $1.29\cdot 10^{-2}$ $8.85\cdot 10^{-3}$ $9.33\cdot 10^{-1}$ $2.06\cdot 10^{-1}$
-Passive ED $3.22\cdot 10^{-1}$ $1.24\cdot 10^{-2}$ $2.49\cdot 10^{-3}$ $4.94\cdot 10^{-1}$ $1.91\cdot 10^{-1}$
-Equalizer LRU $3.00\cdot 10^{-1}$ $\boldsymbol{1.16\cdot 10^{-2}}$ $1.99\cdot 10^{-3}$ $\boldsymbol{1.73\cdot 10^{-1}}$ $9.81\cdot 10^{-2}$
-S4D $3.08\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $\boldsymbol{1.79\cdot 10^{-3}}$ $4.09\cdot 10^{-1}$ $\boldsymbol{9.44\cdot 10^{-2}}$
-S6 $2.80\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $3.89\cdot 10^{-3}$ $2.87\cdot 10^{-1}$ $1.26\cdot 10^{-1}$
-LA-2A LSTM $1.03\cdot 10^{-1}$ $4.51\cdot 10^{-3}$ $1.12\cdot 10^{-2}$ $4.37\cdot 10^{-1}$ $3.10\cdot 10^{-1}$
-Optical ED $8.91\cdot 10^{-2}$ $4.10\cdot 10^{-3}$ $7.28\cdot 10^{-3}$ $2.30\cdot 10^{-1}$ $2.30\cdot 10^{-1}$
-Compressor LRU $1.06\cdot 10^{-1}$ $4.67\cdot 10^{-3}$ $7.84\cdot 10^{-3}$ $2.70\cdot 10^{-1}$ $2.93\cdot 10^{-1}$
-S4D $7.87\cdot 10^{-2}$ $\boldsymbol{3.51\cdot 10^{-3}}$ $1.00\cdot 10^{-2}$ $2.52\cdot 10^{-1}$ $2.53\cdot 10^{-1}$
-S6 $\boldsymbol{7.71\cdot 10^{-2}}$ $3.63\cdot 10^{-3}$ $\boldsymbol{6.26\cdot 10^{-3}}$ $\boldsymbol{1.84\cdot 10^{-1}}$ $\boldsymbol{2.12\cdot 10^{-1}}$
-CL 1B LSTM $2.92\cdot 10^{-1}$ $9.33\cdot 10^{-3}$ $5.21\cdot 10^{-3}$ $3.56\cdot 10^{-1}$ $2.26\cdot 10^{-1}$
-Optical ED $\boldsymbol{1.46\cdot 10^{-1}}$ $\boldsymbol{4.05\cdot 10^{-3}}$ $\boldsymbol{4.18\cdot 10^{-3}}$ $4.17\cdot 10^{-1}$ $\boldsymbol{2.09\cdot 10^{-1}}$
-Compressor LRU $1.88\cdot 10^{-2}$ $6.38\cdot 10^{-3}$ $5.49\cdot 10^{-3}$ $4.38\cdot 10^{-1}$ $4.52\cdot 10^{-1}$
-S4D $2.00\cdot 10^{-1}$ $5.02\cdot 10^{-3}$ $6.12\cdot 10^{-3}$ $\boldsymbol{2.79\cdot 10^{-1}}$ $2.93\cdot 10^{-1}$
-S6 $1.67\cdot 10^{-1}$ $4.41\cdot 10^{-3}$ $6.23\cdot 10^{-3}$ $4.99\cdot 10^{-1}$ $2.36\cdot 10^{-1}$
-Neutron LSTM $1.06$ $3.15\cdot 10^{-2}$ $3.03\cdot 10^{-2}$ $9.33\cdot 10^{-1}$ $6.57\cdot 10^{-1}$
-Filter ED $8.77\cdot 10^{-1}$ $2.18\cdot 10^{-2}$ $2.21\cdot 10^{-2}$ $9.38\cdot 10^{-1}$ $6.32\cdot 10^{-1}$
-LRU $\boldsymbol{6.58\cdot 10^{-1}}$ $\boldsymbol{1.66\cdot 10^{-2}}$ $\boldsymbol{1.95\cdot 10^{-2}}$ $\boldsymbol{7.51\cdot 10^{-1}}$ $\boldsymbol{5.98\cdot 10^{-1}}$
-S4D $7.18\cdot 10^{-1}$ $1.97\cdot 10^{-2}$ $2.00\cdot 10^{-2}$ $9.01\cdot 10^{-1}$ $6.28\cdot 10^{-1}$
-S6 $8.23\cdot 10^{-1}$ $2.17\cdot 10^{-2}$ $2.18\cdot 10^{-2}$ $8.94\cdot 10^{-1}$ $6.27\cdot 10^{-1}$
-Figure 4: Comparison of the RMSE between the target and predicted output.
-The RMSE is computed using windows of $4096$ samples with $75\%$ overlap.
-When the model includes conditioning, the results refer to the highest conditioning values, representing the most challenging scenario.
-In the OD300 and Neutron’s overdrive dataset, the LSTM model presents the lowest MSE and ESR, indicating better general amplitude matching between target and prediction.
-On the other hand, the S4D and S6 present better accuracy when considering the transients, the RMS envelope, and frequency-related content.
-This can be observed from the general lower error in terms of RMSE, spectral flux, STFT, and MFCC.
 In the case of maximum distortion, we can notice LRU prediction having significantly lower energy than other models.
 Notably, the spectrograms shown in Figure 5 illustrate that S4D and S6 models better match the energy in the harmonics at higher frequency regions.
-LRU fails to generate the harmonics in the OD300 case while ED presents the opposite behavior in Neutron’s overdrive.
+LRU fails to generate the harmonics in the OD300 case, while ED presents the opposite behavior in Neutron’s overdrive.
 Models trained on the Neutron OD dataset exhibit larger errors, suggesting that this dataset presents a more challenging modeling case for the investigated architectures.
@@
 The results refer to the highest conditioning values, representing the most challenging scenario.
-For models trained on the Saturator dataset, the LRU model presents lower MSE and $M_{RMSE}$ values, but when considering other metrics, no models appear clearly outperforming.
-When inspecting Figures  4, LRU is again left behind, while other models present closer performances.
+When inspecting models emulating the Saturator, LRU is again left behind, while other models present closer performances.
 In contrast, LRU achieves closer matching in the spectrograms, visible in Figure 5.
-Other models tend to generate more harmonic, especially LSTM that introduces artifacts, which can be clearly heard when listening to the audio output and negatively affect its performance.
-When training models with the Equalizer dataset, LSTM performs better in the time domain, having lower MSE and ESR errors while performing poorly in the frequency domain.
-Figures  4 show close performance among the models, while looking at the spectrograms S6 present better matching.
-The optical compressor features a time-variant response with pronounced temporal dependencies, largely influenced by its variable attack and release parameters.
-This is especially noticeable in the case of the CL 1B compressor, which represents a more challenging task for neural network modeling.
-Among the tested models, the LSTM and LRU are the least accurate in the time domain metrics, exhibiting higher loss in both the LA-2A and CL 1B cases.
-In contrast, in the frequency domain, the gap is less prominent.
-The S6 and ED models demonstrate superior accuracy for these two tasks.
-Specifically, S6 outperforms the ED in the LA-2A case, while the ED excels in the CL 1B case.
-S4D presents similar performance than S6.
+Other models tend to produce more harmonics, particularly the LSTM, which introduces noticeable distortions that are clearly audible in the audio output and adversely impact its performance.
+Models emulating the Equalizer demonstrate similar performance across different designs.
+However, when examining the spectrograms, S6 shows superior matching to the target audio characteristics.
 In the LA-2A case, we observe an overall good match for RMSE in Figure 4.
@@
 Upon examining Figure  5, spectrograms related to the S6 model have the best match with the target output, while the other models present similar mismatches when looking at the higher frequencies.
-More difficult is the CL 1B case, where LRU presents the best envelope matching in Figure 4, and Figure  5.
-In the case of the compressor, the capability to encode past information of the signal is critical, LSTM is confirmed to have a worse temporal tracking ability than the other models, which present design improvements.
-These improvements are particularly evident with the CL 1B when setting long attack and release time through the conditioning control parameters.
+More difficult is the CL 1B case, where LRU presents the best energy matching in Figure 4, and Figure  5.
+In the case of the compressor, the capability to encode past information of the signal is critical, and LSTM is confirmed to have a worse temporal tracking ability than the other models, which present design improvements.
+These improvements are particularly evident with the CL 1B when setting long attack and release times through the conditioning control parameters.
 These reach up to $300$ ms and $5$ seconds, respectively.
-Modeling the low-pass filter is inherently a more complex task than the equalizer, given its more pronounced alterations to the frequency content of the processed signal.
-Nevertheless, informal listening evaluations of the predicted outputs indicate that all models struggle with accurately emulating the filter behavior; instead of eliminating high-frequency contents, they attenuate the whole signal, as apparent in Figure 5.
-However, when considering the quantitative metrics, the LRU model outperforms the others.
-The LSTM model, in particular, tends to incorrectly predict excessive energy on the RMS envelope, as visible in Figure 4, and to introduce artifact, as shown in Figure 5.
+In the case of the low-pass filter, input-output signal alterations are the most pronounced and diverse among the effects considered, challenging the network’s ability to learn this mapping.
+The difficulty mainly stems from the filter’s range, which varies from a straightforward pass-through at high cutoff frequencies to a highly selective filter at low cutoff frequencies.
+Nevertheless, informal listening of the predicted outputs indicates that all models struggle with accurately emulating the filter behavior; instead of eliminating high-frequency contents, they attenuate the whole signal, as apparent in Figure 5.
+The LSTM model, in particular, tends to incorrectly predict excessive energy on the RMS energy, as visible in Figure 4, and to introduce distortions, as shown in Figure 5.
 Generally, performance metrics are relatively close across models across most effects examined in this study.
-Although the S4D model slightly outperforms other models based on the metrics alone, it appears to be the most suitable architecture analog effect modeling, at least considering those within the scope of this study.
-The S6 model often performs very closely, while LRU presents the bigger variance across the different metrics and datasets.
-The ED shows its quality when modeling the compressors.
-The ED, LRU, and particularly the S4D and S6 models are more effective at encoding past signal information within their internal states.
-This is suggested by the case of the cases of optical compressors, where they significantly outperform other models.
-Conversely, the LSTM models appear to perform well only when emulating effects with shorter and less critical time dependencies, such as overdrive, equalizer, and filters.
-Moreover, LSTM models tend to introduce artifacts in some cases.
+When looking at objective metrics, no models appear clearly outperforming the others, but all share similar learning capability.
+MSE, ESR, and $M_{NRMSE}$ errors present the same trends, while performing better in the time domain does not lead to the same in the frequency domain.
+The S4D and S6 models often yield similar performance, while LRU exhibits the greatest variance across different metrics and datasets.
+When modeling compressors, S4D, S6, and ED demonstrate high quality.
+The ED, LRU, and especially S4D and S6 models excel at encoding historical signal information within their internal states, significantly outperforming LSTM models in compressor scenarios.
+Conversely, the LSTM models appear to perform well only when emulating effects with shorter and less critical transients and time dependencies, such as the overdrives.
+Furthermore, in certain instances, LSTM models tend to introduce significant distortion.
+Table 4 presents the $p$-values calculated using the Friedman test, which determines whether there are statistically significant differences between the medians of samples.
+All values are below $0.05$, indicating statistical significance in the results.
+The Wilcoxon results are omitted for brevity, as they involve computing $p$-values for each pair of models trained on the same dataset.
+However, we observe $p$-values exceeding $0.05$ solely when comparisons involve LRU or when models are trained with the low-pass filter dataset.
+This supports the previously mentioned difficulty in learning low-pass filters and the high variance in the LRU model’s performance.
+Table 4: $p$-values calculated from the Friedman test for each metric and dataset presented in Table 3, indicating the statistical significance of the rankings.
+Dataset MSE ESR $M_{NRMSE}$ $M_{SF}$ $M_{STFT}$
+OD300 $0.001$ $0.001$ $0.0008$ $0.011$ $0.0008$
+Neutron Overdrive $0.001$ $0.001$ $0.0004$ $0.0006$ $0.0010$
+Helper Saturator $0.002$ $0.002$ $0.002$ $0.001$ $0.0004$
+Pultec Passive EQ $0.002$ $0.002$ $0.001$ $0.0007$ $0.013$
+LA-2A $0.002$ $0.002$ $0.012$ $0.0006$ $0.002$
+CL 1B $0.002$ $0.002$ $0.002$ $0.001$ $0.0004$
+Neutron Filter $0.002$ $0.002$ $0.004$ $0.018$ $0.002$
 5 Conclusion
-The computational challenges of simulating the nonlinearities of analog effect, described by complex differential equations, are a significant obstacle for real-time applications.
-The complexity of these nonlinearities makes them difficult to model accurately, but they also contribute to the unique sound characteristics of the devices, making them appealing for both academic and industrial purposes.
-Modeling these nonlinear aspects is a difficult task that requires simplifications and assumptions to ensure stability and efficiency in the numerical schemes.
-Artificial neural networks (ANN) have emerged as a popular data-driven approach for this modeling process.
-While they offer a convenient and automated approach, they may struggle with the computational complexity and flexibility required for handling control parameters.
-In some cases, hybrid solutions that incorporate physics knowledge or measurements from internal components may provide advantages.
-However, there are situations where black-box models are necessary and feasible.
-Recurrent neural networks (RNNs) utilize internal states as memory mechanisms, allowing them to encode the signal history.
-This feature enables more interactive solutions as their memory capacity is not solely dependent on the input amount.
-In this paper, we present a comparative analysis that examines the performance of the widely used Long Short-Term Memory (LSTM) models against recently proposed recurrent models, namely Linear Recurrent Unit (LRU) and State-Space models (SSM), which have demonstrated remarkable results in sequence modeling.
-To carry out this study, we design five models incorporating parameter conditioning and providing low-latency response.
-The control parameters are integrated into the models using the Feature-wise Linear Modulation (FiLM) method.
-Our selection of effects includes overdrive, saturation, low pass filter, equalization, and compressors.
-We evaluate these models using various metrics, which provide valuable insights for future research and development.
-Specifically, our metrics consider the root-mean-squared energy envelopes, the transients, and the frequency content.
-LSTM-based networks excel in replicating distortion and equalizer effects, showing their ability to encode the audio dependencies enough for these specific effects effectively.
-Conversely, LSTMs tend to introduce more artifacts than other models considered in this study.
-In contrast, LSTM, when designed in an encoder-decoder fashion, LRUs and, in particular, SSMs, outperform other models when emulating saturation and compression.
-In the case of compression, they demonstrate their advantage in accurately capturing the signal’s history.
-Generally, SSMs have also proved to be an appropriate candidate technique in VA, although the gap with the other models is not always so pronounced.
-Lastly, LRUs are less consistent in their quality across the different metrics.
-In this study, we did not carry out architecture—and effect-specific optimizations, which can definitely boost the accuracy, such as tuning hyperparameters and changing the size of the various layers.
-We limited computational complexity and input-output latency and focused the study on models that can be theoretically implemented in consumer-grade devices for live audio applications.
+Modeling analog effects is challenging, requiring simplifications to maintain numerical stability and efficiency.
+ANNs have gained popularity for their convenient, automated approach, but they can struggle with computational complexity and flexibility for control parameters.
+While hybrid solutions incorporating physics can offer advantages, black-box models are sometimes necessary.
+Recurrent neural networks, utilizing internal states for memory, provide interactive solutions by encoding signal history, enhancing their modeling capability beyond input size limitations.
+In this paper, we conducted a comparative analysis to assess the performance of LSTM models against recently proposed recurrent models, specifically LRU and SSMs, which have shown impressive results in sequence modeling.
+Five models were designed with parameter conditioning and low-latency response using the Feature-wise Linear Modulation method, focusing on effects like overdrive, saturation, low-pass filter, equalization, and compressors.
+Our evaluation employed metrics considering root-mean-squared energy, transients, and frequency content for valuable insights into future research.
+Experimental results indicate that LSTM-based networks perform well in emulating distortion and equalizer effects.
+However, they tend to introduce noticeable distortions compared to other models in this comparative study.
+Conversely, LSTM structured in an encoder-decoder configuration, and SSMs, perform better in emulating tape saturation and compression effects, where encoding historical signal information may significantly impact the accuracy.
+However, the performance gap is not always pronounced.
+Regarding the low-pass filter, no architecture successfully learns its audio processing characteristics.
+In general, MSE, ESR, and $M_{NRMSE}$ errors present the same trends, while performing better in the time domain does not lead to the same in the frequency domain.
+Lastly, LRU shows inconsistent performance across different datasets.